### v1

# MS-IMAP - A Multi-Scale Graph Embedding Approach for Interpretable Manifold Learning

MS-IMAP - A Multi-Scale Graph Embedding Approach for Interpretable Manifold Learning
Shay Deutsch  [^1] Corresponding author.  [^4] Emerging Capabilities Research Group, Discover Financial Services Inc., Riverwoods, IL.
shaydeu@gmail.com
   Lionel Yelibi  [^4] Emerging Capabilities Research Group, Discover Financial Services Inc., Riverwoods, IL.
ylblio001@myuct.ac.za
   Alex Tong Lin  [^4] Emerging Capabilities Research Group, Discover Financial Services Inc., Riverwoods, IL.
atlin271@gmail.com
   Arjun Ravi Kannan  [^4] Emerging Capabilities Research Group, Discover Financial Services Inc., Riverwoods, IL.
arjun.kannan@gmail.com
1 Abstract
Deriving meaningful representations from complex, high-dimensional data in unsupervised settings is crucial across diverse machine learning applications. This paper introduces a framework for multi-scale graph network embedding based on spectral graph wavelets that employs a contrastive learning approach. A significant feature of the proposed embedding is its capacity to establish a correspondence between the embedding space and the input feature space which aids in deriving feature importance of the original features. We theoretically justify our approach and demonstrate that, in Paley-Wiener spaces on combinatorial graphs, the spectral graph wavelets operator offers greater flexibility and better control over smoothness properties compared to the Laplacian operator. We validate the effectiveness of our proposed graph embedding on a variety of public datasets through a range of downstream tasks, including clustering and unsupervised feature importance.
2 Introduction
Graph Embeddings and Manifold Learning [24, 1, 27, 28, 3] are pivotal in analyzing complex data structures prevalent in diverse machine learning applications. The emergent representations from these techniques offer insights into the underlying data structure in conjunction with downstream tasks like clustering and visualization and are particularly valuable when labels are unavailable or unreliable. However, limitations of these techniques include reliance on low frequencies of the graph Laplacian and $k$-Nearest Neighbors graph connectivity, which provide limited information, including in recent contrastive learning methods such as Uniform Manifold Approximation and Projection (UMAP). To address these limitations, this paper proposes three contributions. (1) A framework employing multi-scale graph representation using contrastive learning which enhances the expressiveness of an embedding by optimizing low and high-frequency information, capturing intricate data details. (2) A characterization of the theoretical properties of the representation power of the spectral graph wavelets (SGW) operator by considering functions sampled from the Paley-Wiener spaces [22] on combinatorial graphs and a demonstration of the flexibility and better control over smoothness properties of the SGW operator in comparison to the Laplacian operator. (3) A novel sampling technique that leverages the graph structure by sampling edges or nodes from a distribution reflecting network characteristics.
A key distinction between our approach and many manifold learning methods such as UMAP is that we leverage multi-scale graph representation (using SGW) to increase the dimensionality of the data thus enhancing data representation learning. On the other hand, other manifold learning techniques predominantly employ nonlinear dimensionality reduction to reduce the dimensions of the output embedding space. We use stochastic gradient descent (SGD) to optimize the embedding space and employ a novel contrastive learning approach using a 3D-tensor to capture both low and high frequencies.
Another limitation of non-linear manifold learning methods is they inherently lack direct connections to input features, thereby limiting their ability to offer interpretable results which is crucial in various applications (e.g. in finance where customer behavior needs to be summarized with a few important features to design products or strategies benefiting them and in biology where easily interpretable features may lead to life saving discoveries). Although not the main focus of this article, our proposed approach establishes a mapping between the original features and the constructed optimized embeddings and enables the measurement of the importance of the original features with respect to the embedding space (see Figure.1 for an illustration of the mapping generated by the proposed embedding). To the best of our knowledge, current manifold learning techniques produce embeddings lacking explicit interpretation. In contrast, our approach demonstrates the ability to generate interpretable embedding representations, while also being competitive with the state-of-the-art graph embeddings across applications such as finance, vision, and biology.
3 Related Work
Extensive efforts have been dedicated to exploring non-linear dimensionality reduction aimed at achieving a low-dimensional embedding that preserves the underlying manifold structure. Early manifold learning techniques include Locally Linear Embedding (LLE) [24], Laplacian Eigenmaps [1], Isomap [27], Diffusion Maps [2], and t-distributed Stochastic Neighbor Embedding (t-SNE) [28].
Balancing local and global structure in the resulting embeddings is challenging in the presence of noise. To address the challenge, researchers have introduced de-noising methods [17, 10, 6, 9, 5, 11, 7, 8]. These techniques seek to enhance algorithms robustness, thereby broadening the applicability of these techniques in real-world scenarios.
UMAP [21] has proven effective in providing vivid visualizations and cluster separation, yet its reliance on Laplacian Eigenmaps initialization and negative sampling optimization has limitations. Laplacian Eigenmaps initialization focuses on low-frequency signal overlooking higher-frequency patterns. It uniformly samples edges without considering topological importance, overemphasizing graph connectivity and introducing further distortion.
Recent work [4] further highlights that negative sampling has a significant impact on the effective loss and introduces distortions with t-SNE and UMAP which motivates the exploration of alternate sampling schemes. Other relevant studies include [30] which investigates competing negative sampling strategies, [25] on generalization error bounds , and [20] on topological and sampling bias due to high degree nodes.
The lack of explicit mapping linking the original high dimensional dataset to its low dimensional embedding has prevented their use in areas where interpretability is critical. Establishing this link clearly would lead to a better understanding of the original feature space and improve feature selection, understanding of the resulting visualizations, and robustness to noise [18, 13, 31].
SGW serve as an efficient tool for localization in both the spectral and vertex domains, playing a crucial role in our method (see Section 5.1 for details). SGW are constructed by applying a set of bandpass filters to the eigenvalues of the graph Laplacian, enabling multi-resolution representation of graph signals. An alternative multi-scale framework for analyzing functions on graphs is Diffusion Wavelets [23]. The Geometric Graph Scattering transform and the Graph Scattering Transform [14] utilize Diffusion Wavelets to extract graph features, which are then employed in graph classification tasks. In our work, we leverage the localization properties of SGW to develop a novel feature representation scheme, achieving superior performance in our target applications.
4 Preliminaries
Consider a set of points $\mathbf{x}=\left\{\mathbf{x}_{i}\right\},\,i=1,...N,\mathbf{x}_{i}\in\mathbb{R}^{D}$, which are sampled from an unknown manifold $M\subset\mathbb{R}^{D}$. An undirected, weighted graph $G=(V,\mathbf{W})$, is constructed over $\mathbf{x}$, where $V$ corresponds to the nodes and $\mathbf{W}$ to the set of edges on the graph. The adjacency matrix $\mathbf{W}=(w_{ij})$ consists of the weights $w_{ij}$ between node $i$ and node $j$. The weights can be chosen using different techniques, such as the Gaussian kernel function or adaptive graph construction. If the weights are chosen using the Gaussian kernel function then
$(\mathbf{W})_{ij}=\left\{\begin{matrix}\exp\left(\frac{-||\mathbf{x}_{i}-\mathbf{x}_{j}||_{2}^{2}}{2\sigma_{D}^{2}}\right)&\mbox{if}\>\mathbf{x}_{j}\in\text{kNN}(\mathbf{x}_{i})\\ 0&\mbox{otherwise}\end{matrix}\right.$ (1)
where is (||||)(||\,||)( | | | | ) denotes the $L2$ distance between the points $\mathbf{x}_{i},\mathbf{x}_{j}$, kNN$(\mathbf{x}_{i})$ denotes the $k$ nearest neighbors of $\mathbf{x}_{i}$, and $2\sigma_{D}^{2}$ are parameters used to construct the graph.
An alternative is to use adaptive graph construction as proposed in [21]. The global smoothness of the graph signal function ${f}\in\mathbb{R}^{N}$ (a function over the vertices of the graph $G$) is defined using the graph Laplacian quadratic form as follows:$||\bigtriangledown{f}||^{2}=\sum_{V(i,j)}{w_{ij}(f(i)-f(j)})^{2}={f}^{T}\mathbf{L}{f}$, where $\mathbf{L}$ denotes the combinatorial graph Laplacian, defined as $\mathbf{L}=\mathbf{D}-\mathbf{W}$, with $\mathbf{D}$ the diagonal degree matrix with entries $d_{ii}=d(i)$. The degree $d(i)$ of vertex $i$ is defined as the sum of weights of edges that are connected to $i$. The normalized Laplacian is defined as ${\mathbf{L}}_{N}=\mathbf{D}^{-\frac{1}{2}}\mathbf{L}\mathbf{D}^{-\frac{1}{2}}=\mathbf{I}-\mathbf{D}^{-\frac{1}{2}}\mathbf{W}\mathbf{D}^{-\frac{1}{2}}$ and its real eigenvalues are in the interval $[0,2]$. The eigenvalues and eigenvectors of $\mathbf{L}$ are $\lambda_{1},\ldots,\lambda_{N}$ and $\mathbf{\phi}_{1},\ldots,\phi_{N}$, respectively. The graph Fourier transform (GFT) $\hat{f}$ of the graph signal $f$ is defined as the expansion of $f$ in terms of the eigenvectors $\phi$ of the graph Laplacian, so that for frequency $\lambda_{l}$ we have:
$\hat{f}(\lambda_{l})=\sum_{i}{f(i)\phi_{l}^{*}(i)}.$ (2)
We will denote the matrix representations of the eigenvectors and eigenvalues of the graph Lapacian $\mathbf{L}$ as $\Phi,\lambda$, respectively. Our approach assumes that all graph signals $f_{r}\in\mathbb{R}^{N}$ used to create the embedding correspond to the input coordinates, $\mathbf{x}=\left(f_{1},f_{2},..,f_{r},..f_{D}\right)$. The output embedding space provides a multi-scale representation approximating the intrinsic manifold coordinates.
4.1 Multi-scale representations using SGW
In the last two decades, several multi-scale representations using irregular graphs have been suggested, including SGW [15] and Diffusion Wavelets [23]. In this work we focus on the multi-scale graph transform utilizing SGW. SGW provides a natural way of trading off between spectral and spatial resolution, since the SGW coefficients are localized in both domains. These wavelets are constructed using a kernel function $g(\mathbf{L})$ operator, which acts on a function by modulating each of its Fourier modes [15], that helps in capturing a trade off between vertex (spatial) and spectral localization. Spatial localization in the graph domain is implicitly controlled by a single wavelet scale parameter defined in the spectral domain, such that the more the vertex is localized in the vertex domain, the spectral band is wider. The scale parameter enables the model to adjust the effective neighborhood sizes to the properties of the data.
To represent a signal ${f}\in\mathbb{R}^{N}$ in multiple scales $S=[s_{1},s_{2},..s_{K}]$, the SGW transform is constructed as follows: Assume that $g(\lambda)$ is a filter in the spectral domain. Let $\delta_{i}\in\mathbb{R}^{N}$ be the delta function centred at the vertex $i\in G:\delta_{i}(j)=1\,\,\mbox{if}\,\,i=j$, and $\delta_{i}(j)=0$ otherwise. Given a bandpass filter $g(\lambda)$ and a wavelet centered on node $i$ at scale $s$, the wavelets $\psi(s,i),i=1,..N$ are calculated by applying them to the delta function on a single vertex $i$, given by
$\psi(s,i)=\Phi g(s\lambda)\Phi^{T}\delta_{i}$ (3)
The value of $\psi(s,i)$ with respect to a vertex $m$ can be written as $\psi(s,i)(m)=\sum_{l=1}^{N}g(s\lambda_{l})\phi_{l}^{*}(i)\phi_{l}(m)$. Given a graph signal ${f}\in\mathbb{R}^{N}$, the SGW coefficient at node $i$ and scale $s$ can be expressed as follows:
$\psi_{f}(s,i)=\sum_{l=1}^{N}g(s\lambda_{l})\hat{f}(\lambda_{l})\phi_{l}(i)$ (4)
Fast computation using Chebyshev polynomials: Directly computing the SGW coefficients requires calculating the entire eigensystem of the Laplacian, which is computationally intensive - $O(N^{3})$ for $N$ points. Instead, Hammond et al. [15] suggested computing the SGW using a fast algorithm based on approximating the scaled generating kernels through low-order polynomials of $\mathbf{L}$ applied to the input data (details in Appendix).
5 Our Proposed framework: Multi-scale IMAP
Figure 1: The encoding step of our framework is illustrated as a mapping between each coordinate of the input features and a corresponding dimension in the embedding space. This mapping is facilitated through our proposed approach, which aligns the input dimensions of the features with the latent embedding space. By establishing this correspondence, we acquire the capability to interpret each dimension in the embedding space. Such interpretation allows for various analyses, including providing feature importance relative to the latent space.
In this section, we introduce Multi-scale IMAP, a framework for interpretable embedding via manifold learning utilizing a multi-scale graph representation. This approach enables us to maintain global regularity and preserve local structure without sacrificing interpretability. Our method imposes a differentiable structure on the mapping $h:\mathbf{x}\rightarrow\psi_{\mathbf{x}}$, supported on a discrete graph $G=(V,\mathbf{W})$, where $\psi_{\mathbf{x}}$ represents the encoded multi-scale graph transform of $\mathbf{x}$. Multi-scale IMAP consists of three main components:
Step 1: Features representation encoding using multi-scale graph representation.
In this step, our approach constructs a multi-scale graph representation by incorporating feature signals through the SGW Transform across multiple scales and graph frequencies. We introduce two methods for encoding multi-scale transform: in the first, based on simple concatenation, all of the features transformed for all scales are represented by concatenating all the SGW coefficients for each point in a vector form. In the second based on 3D tensor optimization, we construct a 3D tensor with dimensions ${K}\times{N}\times{D}$. We leverage the encoded 3D tensor structure to align the optimized embedding with all scales simultaneously to enforce differentiable structure of the transformed features.
Step 2: Optimization design using multi-scale network structure.
We integrate the multi-scale representation within the SGD optimization framework, simultaneously leveraging both low and high-frequency information. This integration leads to fine-grained manifold regularization and improved robustness in downstream tasks.
Step 3: Network features sampling strategy.
Our approach introduces a sampling scheme that selects network features based on their estimated importance in the graph network. Employed within the optimization using SGD, this adaptive strategy dynamically prioritizes and focuses on informative parts of the graph during the embedding process. This adaptability enhances the optimization process, leading to increased efficiency in generating embeddings.
5.1 Feature Representation to encode multi-scale structure in the optimized embeddings
We propose two alternative methods to encode multi-scale representations used for subsequent optimization. Note that for both method 1 and method 2, each dimension in the embedding space is constructed using a single feature in the original feature set, which is an essential characteristic that can be leveraged for interpreting graph embeddings.
[…]
$\mathbf{c}(\psi_{{f}_{i}}(s_{k},:),\psi_{{f}_{j}}(s_{k},:))=\psi_{{f}_{i}}(s_{k},:)\left|\right|\psi_{{f}_{j}}(s_{k},:)$
denoting the concatenation of the vectors corresponding to the multi-scale representation $\psi_{{f}_{i}}(s_{k},:)$ and $\psi_{{f}_{j}}(s_{k},:)$. For method (i) (also noted as method 1) where all features and all scales are concatenated together, the resulting matrix $\psi_{\mathbf{x}}$ can be represented as
$\psi_{\mathbf{x}}=\psi_{{f}_{1}}(s_{1},:)\|\psi_{{f}_{2}}(s_{1},:)\|....\|\psi_{{f}_{D-1}}(s_{K},:)\|\psi_{{f}_{D}}(s_{K},:)$
where $\psi_{\mathbf{x}}\in\mathbb{R}^{KD}\times\mathbb{R}^{N}$.
(ii) 3D-Tensor Encoding - Method 2 The second encoding method concatenates the multi-scale representation for all features at a fixed scale, generating a matrix representation $\psi_{\mathbf{x}}(s_{j},:,:)\in\mathbb{R}^{D}\times\mathbb{R}^{N}$ for each scale $s_{j}$. For method 2, after concatenating all $f_{i}$ at a fixed scale $s_{j}$, we have
$\psi_{\mathbf{x}}(s_{j},:,:)=\psi_{{f}_{1}}(s_{j},:)\|\psi_{{f}_{2}}(s_{j},:)\|...\|\psi_{{f}_{D}}(s_{j},:)$
The optimization can be performed separately for each scale $s_{j}$, or jointly for all scales $s_{j}$ using the 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$ We detail the proposed feature representation encoding methods in the pseudo code algorithm 1.
  Input: Set of points $\left\{\mathbf{x}_{i}\right\}_{i=1}^{N}$.
  Output: Initial Node Embeddings ${\psi}_{\mathbf{x}}$
  Step 1: Construct $G=(V,\mathbf{W})$ from $\left\{\mathbf{x}_{i}\right\}_{i=1}^{N}$.
  Step 2: Construct the Laplacian $\mathbf{L}$.
  Step 3: Compute $\lambda_{\mbox{max}}(\mathbf{L})$.
1  Step 4: Initial Embedding construction:
2For $r=1,..,D$ associated with the feature signals $\left\{f_{r}\right\}$ do:
3Compute $\psi_{{f}_{r}}(s_{j},:)$ at scales $s_{j},j=1,...K$ using Chebyshev approximation.
4If Encoding Method 1: Concatenate $\psi_{{f}_{r}}(s_{j},:)$, $j=1,...K$, $r=1,...,D$ in a matrix form $\psi_{\mathbf{x}}\in\mathbb{R}^{KD}\times\mathbb{R}^{N}$.
If Encoding Method 2: Concatenate $\psi_{{f}_{r}}(s_{j},:)$ for $r=1,...,D$ in a matrix form $\psi_{\mathbf{x}}(s_{j},:,:)$ for each fixed scale $s_{j}$, representing all scales using 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$.
Algorithm 1 Encoding multi-scale structure in the optimized embeddings
5.2 Optimization
We propose to use optimization based on SGD that begins with the initial embedding within the spectral graph domain. While employing positive and negative sampling, our aim is to optimize the embedding space directly in the SGW domain, revealing the intrinsic structure of manifold data while retaining high-frequency information associated with the graph Laplacian. This optimization within the SGW domain provides more effective regularization, allowing for the direct removal of noise from each spectral band. We outline the steps of the proposed approach, presenting several alternatives for optimization through the incorporation of multi-scale representations (refer to method 1 and method 2 in the algorithm description below).
1. Optimize Embedding Method 1:
Given the encoded multi-scale representation $\psi_{\mathbf{x}}\in\mathbb{R}^{KD}\times\mathbb{R}^{N}$ perform optimization in the SGW domain, using the following fuzzy cross entropy loss function:
$\mathcal{L}(\tilde{\psi}_{\mathbf{x}}|\mathbf{W})=\sum_{i,j}\left(w_{ij}\mbox{log}\frac{w_{ij}}{v_{ij}^{\psi_{\mathbf{x}}}}+(1-w_{ij})\mbox{log}\frac{1-w_{ij}}{1-v_{ij}^{\psi_{\mathbf{x}}}}\right)$ (5)
where $v_{ij}^{\psi_{\mathbf{x}}}=\frac{1}{1+||\psi_{\mathbf{x}_{i}}-\psi_{\mathbf{x}_{j}}||^{2}}$.
Dropping terms that do not depends on $\psi_{\mathbf{x}_{i}}$, the gradient of the loss is approximated by:
$\frac{\partial\mathcal{L}(\tilde{\psi}_{\mathbf{x}}|\mathbf{W})}{\partial\psi_{\mathbf{x}_{i}}}\sim\sum_{j}w_{ij}v_{ij}^{\psi_{\mathbf{x}}}(\psi_{\mathbf{x}_{i}}-\psi_{\mathbf{x}_{j}})-\sum_{j}\frac{1}{||\psi_{\mathbf{x}_{i}}-\psi_{\mathbf{x}_{j}}||^{2}}v_{ij}^{\psi_{\mathbf{x}}}(\psi_{\mathbf{x}_{i}}-\psi_{\mathbf{x}_{j}})$ (6)
Our optimization process involves using SGD with positive and negative sampling, similar to recent graph embeddings method such as UMAP and LargeVis [26]. Positive edge samples are associated with attraction (first term on the right hand side), while negative samples (second term on the right hand side) refer to a pair of nodes that are not connected in the graph, which create repulsion among dissimilar points.
Output: Regularized embedding space $\tilde{\psi}_{\mathbf{x}}$.
2. 3D-Tensor Based Optimization: Embedding Method 2:
In method 2, the encoded manifold representation given is a 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$. A SGD update rule at iteration $t$ applied to the 3D tensor is:
$(\tilde{\psi}_{\mathbf{x}}^{(t+1)})_{i,j,k}=(\tilde{\psi}_{\mathbf{x}}^{(t)})_{i,j,k}-\alpha\frac{\partial\mathcal{L}}{(\partial\psi_{\mathbf{x}})_{i,j,k}}$ (7)
where $\alpha$ is the learning parameter. In our case we employ the cross entropy loss function:
$\mathcal{L}(\tilde{\psi}_{\mathbf{x}}|\mathbf{W})=\sum_{i,j,k}\left(w_{ij}\mbox{log}\frac{w_{ij}}{v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}}+(1-w_{ij})\mbox{log}\frac{1-w_{ij}}{1-v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}}\right)$ (8)
We apply the optimization based on SGD with respect to each scale $s_{k}$:
$\mathcal{L}(\tilde{\psi}_{\mathbf{x}}(s_{k},:,:)|\mathbf{W})=\sum_{i,j}\left(w_{ij}\mbox{log}\frac{w_{ij}}{v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}}+(1-w_{ij})\mbox{log}\frac{1-w_{ij}}{1-v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}}\right)$ (9)
which has the gradient of the loss is approximated by:
$\frac{\partial\mathcal{L}(\tilde{\psi}_{\mathbf{x}}(s_{k},:,:)|\mathbf{W})}{\partial\psi_{\mathbf{x}_{i}}(s_{k},:,:)}=\sum_{j}w_{ij}v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}(\psi_{\mathbf{x}_{i}}(s_{k},:,:)-\psi_{\mathbf{x}_{j}}(s_{k},:,:))$
$-\sum_{j}\frac{1}{||\psi_{\mathbf{x}_{i}}(s_{k},:,:)-\psi_{\mathbf{x}_{j}}(s_{k},:,:)||^{2}}v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}(\psi_{\mathbf{x}_{i}}(s_{k},:,:)-\psi_{\mathbf{x}_{j}}(s_{k},:,:))$
The optimization process uses SGD with positive and negative sampling for all scales $s_{k}$ given a pair of positive or negative pair of nodes $i,j$.
We compute the final embedding by summing up all the optimized embeddings with respect to each scale $s_{k}$
$\tilde{\psi}_{\mathbf{x}}=\sum_{k}\tilde{\psi}_{\mathbf{x}}(s_{k},:,:)$ (10)
In the 3D tensor based optimization method 2, the dimensionality of the final embedding $\tilde{\psi}_{\mathbf{x}}\in\mathbb{R}^{D}\times\mathbb{R}^{N}$, maintaining the same dimensionality as the original input, and thus there exists a one-to-one correspondence between the coordinates of the input features $\mathbf{x}$ and $\tilde{\psi}_{\mathbf{x}}$. This correspondence ensures that each dimension of the feature space is directly mapped to a single dimension in the embedding space $\tilde{\psi}_{\mathbf{x}}$.
Sampling approaches from network features: We propose a strategic sampling of edges, which further refines the optimization by assessing the significance of edges with respect to the topological structure of the graph. Specifically, we construct probability distribution over $V$ (or $\mathbf{W}$), providing the significance of each node/edge, and sample from this distribution. We also propose a novel method to implement the sampling strategy, rather than selecting edges randomly as in [21, 26]. Refer to Appendix for more details.
6 Theoretical results: Sampling set for Smooth Manifolds with functions defined over Paley-Wiener Spaces
In this section, we characterize the theoretical properties of the representation power of the SGW operator by considering functions sampled from the Paley-Wiener spaces [22] on combinatorial graphs. Pesenson introduced the Paley-Wiener spaces and demonstrated that Paley-Wiener functions of low type are uniquely determined by their values on certain subsets, known as uniqueness sets $U$. We show that the SGW operator can represent functions $f$ within the Paley-Wiener space more efficiently than the graph Laplacian operator $\mathcal{L}$. This efficiency is demonstrated by showing that the SGW operator is more effective in representing functions with larger bandwidth $\omega$ in the Paley-Wiener spaces (i.e., with higher frequencies) using subsets of nodes from the uniqueness sets $U$.
To characterize the representation properties of functions defined over $PW_{\omega}(G)$, we employ the notion of the $\Lambda$-set, introduced by Pesenson which is central to our investigation. Formally, the Paley-Wiener space of $\omega$-bandlimited signals is defined as $PW_{\omega}(G)=\left\{f\mid\hat{f}(\lambda)=0\,\,\forall\,\,\lambda>\omega\right\}$. The space $L_{2}(G)$ is defined as the Hilbert space of all complex-valued functions, and $L_{2}(S)$ is defined as the space of all functions from $L_{2}(G)$ with support in $S$: $L_{2}(S)=\left\{\varphi\in L_{2}(G)\mid\varphi(v)=0,v\in V(G)\setminus S\right\}$.
The $\Lambda$-set is defined as follows: a set of vertices $S\subset V$ is a $\Lambda(S)$-set if any $\varphi\in L_{2}(S)$ satisfies the inequality $||\varphi||\leq\Lambda||\mathcal{L}\varphi||$ for some constant $\Lambda(S)>0$. The infimum of all $\Lambda>0$ for which $S$ is a $\Lambda$-set is denoted by $\Lambda$.
The ability of the SGW operator to efficiently represent functions $f\in PW_{\omega}(G)$ can be summarized in the following theorem, which highlights the role of the $\Lambda_{\psi}$-set with respect to the operator $\psi$. We show that any signal $f\in PW_{\omega}(G)$, where $\lambda_{1}\leq\omega<\Omega_{G}$ for some $\Omega_{G}<\lambda_{N}$, can be uniquely represented by its samples on the uniqueness set $U$ using the SGW operator. Under certain conditions related to the SGW operator, its associated $\Lambda_{\psi}$-set is smaller than the $\Lambda$-set associated with the Laplacian operator.
Theorem 1
Let $G=(V,\mathbf{W})$ be a connected graph with $n$ vertices and $f\in PW_{\omega}(G)$ for $\lambda_{1}\leq\omega<\lambda_{\text{max}}$. The SGW operator $\psi$ can be constructed such that the set $S$ is a $\Lambda_{\psi}$-set and the set $U=V\setminus S$ is a uniqueness set for any space $PW_{\omega}(G)$ with $\omega<1/\Lambda_{\psi}$ and $\Lambda_{\psi}<\Lambda$ for any $\varphi\in L_{2}(S)$, where $\Lambda$ is the $\Lambda$-set of the Laplacian operator.
The proof is provided in the Appendix.
7 Experimental Results
We demonstrates the advantage of our approach over existing approaches using both synthetic and real datasets. We evaluate our method by testing its clustering performance using the output embedding and comparing it to several representative methods, including UMAP, t-SNE, Diffusion Maps, ISOMAP, and HeatGeo [19]. For HeatGeo, we hyperparameter tune the performance on each dataset and report the best one in Section 11. For each dataset, we briefly describe their properties and leave fuller details in Section 9.
Evaluation metrics: We test performance using the Adjusted Rand Index (ARI) and Adjusted Mutual Information (AMI). In all experiments we used $k$-means to cluster the data in the embedding space.
7.1 Synthetic dataset
We assess the robustness of our method using the two moons dataset, which is a 2D manifold depicting two interleaving half-circles. We sampled $N=600$ points, and set the Gaussian noise’s standard deviation to $0.12$. While spectral based methods such as UMAP are effective under relatively “modest” noise levels, their performance deteriorates significantly in the presence of larger amounts of noise (See Figure 3). As shown in Table 1, our approach shows robustness to noise and correctly clusters most points in two moons manifold, despite the large amount of noise, and outperforms competing methods. We do further experiments on more synthetic datasets in Section 10.
Method / Accuracy ARI AMI
UMAP 0.54 0.51
t-SNE 0.42 0.35
ISOMAP 0.36 0.3
Diffusion Maps 0.25 0.19
HeatGeo 0.54 0.52
MS-IMAP Method 1 0.75 0.73
MS-IMAP Method 2 0.89 0.87
Table 1: Comparison of clustering performance on the Two Moons datasets.
7.2 Real Datasets
We study the performance of MS-IMAP compared to other methods for real datasets. We chose a mix of financial, biological, and image datasets: the Census dataset [12] is a financial dataset containing information about individuals extracted from the 1994 US Census; the Zilionis dataset [32] is a biological dataset containing single-cell sequencing data from different types of cells, and the Animals with Attributes (AWA) [29] dataset is an image dataset containing images of animals. More information about each dataset can be found in Section 9.
As shown in Table 2, MS-IMAP Method 2 obtains the best performance across all datasets, with a tie between UMAP and MS-IMAP Method 2 in the AWA dataset. In the Census dataset, MS-IMAP Method 2 achieves a $47\%$ increase in ARI over HeatGeo, the next best method excluding our own. In the Zilionis dataset, MS-IMAP 2 achieves a $9\%$ increase in ARI over t-SNE (we found issues running the official HeatGeo code [19] on Zilionis, so were unable to obtain results for this dataset). And in AWA, we have the ARI of MS-IMAP Method 2 is slightly better than UMAP.
Dataset Census Zilionis AWA
Method / Accuracy ARI AMI ARI AMI ARI AMI
UMAP 0.05 0.08 0.45 0.71 0.73 0.80
t-SNE 0.01 0.02 0.64 0.34 0.68 0.76
HeatGeo 0.15 0.10 N/A N/A 0.65 0.74
MS-IMAP Method 1 0.15 0.14 0.69 0.76 0.72 0.80
MS-IMAP Method 2 0.22 0.15 0.70 0.76 0.74 0.81
Table 2: Clustering results comparison using ARI and AMI the Census, Zilionis, and AWA datasets. Note HeatGeo had code execution issues on the Zilionis dataset.
8 Explicit Feature Correspondence with Laplacian Scoring
The correspondence established between the features and the embedding space becomes valuable in obtaining feature importance. One simple measure for feature importance is the Laplacian score (LS) [16], which evaluates each feature based on its correlation with the graph Laplacian eigensystem. Specifically, the importance of a feature $f_{r}\in\mathbb{R}^{N}$ is determined from the importance of its corresponding coordinate $l$ in the embedding space, $\tilde{\psi}_{\mathbf{x}_{l}}$. We calculate the Laplacian score [16] with respect to the embedding features $\tilde{\psi}_{\mathbf{x}}$ using its graph Laplacian $\mathbf{L}$ and its associated degree matrix $\mathbf{D}$. This is achieved by first removing the mean and then computing the Laplacian score using:
$L_{s}(\tilde{\psi}_{\mathbf{x}})_{l}=\frac{(\tilde{\psi}_{\mathbf{x}})_{l}^{T}\mathbf{L}(\tilde{\psi}_{\mathbf{x}})_{l}}{(\tilde{\psi}_{\mathbf{x}})_{l}^{T}\mathbf{D}(\tilde{\psi}_{\mathbf{x}})_{l}}$ (11)
Smaller scores imply that the feature $(\tilde{\psi}_{\mathbf{x}})_{l}$ has a larger projection on the subspace of the eigenvectors associated with the smallest eigenvalues, indicating higher importance with respect to the graph structure. Each coordinate $l$ in the embedding space $\tilde{\psi}_{\mathbf{x}}$, namely $(\tilde{\psi}_{\mathbf{x}})_{l}$, was constructed using a single coordinate of the original feature $f_{r}$ space. Hence, the feature importance of $(\tilde{\psi}_{\mathbf{x}})_{l}$ can be explicitly interpreted as the importance of the associated original feature.
8.1 Application to Feature importance
Figure 2: AWA Dataset Laplacian Score plotline showing in order from high to low, the importance of the original features with respect to the embeddings.
We demonstrate the effectiveness of our approach using a computer vision example where providing interpretation, including feature importance, is of significant interest. Here we show in Figure (2) the estimated Laplacian score of the 85 semantic attributes for the AWA datasets sorted from high to low. Because the Laplacian score is a function of explained variance, one can argue that this score is a measure of feature information and we observe that the number of features capturing the highest scores is small (on the order of 10). This result highlights that most of the information in the data-set which correlates with the embedding is captured by a small group of variables which suggests their relatively higher importance. The Laplacian score ranking allows us to select a small number of variables for further analysis in order to better understand the output of statistical studies performed using the embedded data.
8.2 Runtime and Computational Complexity
The execution time of our method with Python code implementation with 32 cores Intel Xeon 8259CL running at 2.50Ghz and 256GB of RAM on the Cancer QC data-set of 48,969 samples and 306 features took 12.44 mins using Method 1. The computational complexity of Multi-Scale IMAP is of $O(NDlog(N))$. Additional details are provided in the appendix.
8.3 Conclusions and Limitations
Identifying the key drivers of high-dimensional datasets is a necessary ingredient to take advantage of unlabelled data in many practical applications in fields such as finance and biology. In this article, we introduce a novel contrastive learning framework for manifold learning via graph embeddings, capitalizing on both low and high-frequency information. In particular, we use SGW to construct a multi-scale graph representation of the underlying input feature space. We then use an SGD-based optimization scheme together with innovative strategies such as 3D Tensor encoding to derive the embeddings. We study the theoretical properties of the spectral graph wavelet (SGW) representation by considering functions in Paley-Wiener spaces on combinatorial graphs and prove that the SGW operator provides a more effective representation using the concept of the $\Lambda$set. Finally, we show that the embeddings are interpretable using a simple derivation of feature importance of the embedding and original feature spaces. The construction of our methodology for generating the embeddings implicitly gives us a way to tie the original and transformed feature spaces together which is lacking in current non-linear manifold learning techniques to the best of our knowledge.
Limitations: While our work has demonstrated several significant properties with strong performance on challenging datasets, it is not without limitations. Firstly, we rely on the assumption that the features, and thus the similarity measure, adequately capture information to reveal geodesic distances between manifold points. Secondly, utilizing the current optimization framework with SGD may present difficulties when extending to out-of-sample problems. Furthermore, although our embedding feature correspondence proves valuable in assessing feature importance through a global method, there remains potential for enhancement through the incorporation of local methods.
Appendix
9 Dataset Details
Two Moons. The two moons dataset depicts two interleaving half-circles. We sampled $N=600$ points and used a Gaussian noise level having standard deviation $0.12$. An example is show in Figure 3. More specifically, to produce these points we use Sci-kit Learn’s sklearn.datasets.make_moons function with $n\_samples=600$ and $noise=0.12$.
In Figure 3, we show example clusterings for each of the methods mentioned in Section 7.
Census. From the UCI Machine Learning Repository [12], this dataset contains 14 features that are a mix of categorical, numerical, and binary. Such features include age, marital status, sex, etc. The goal is to predict whether a sample makes less than or equal to \$$\$50,000$, or strictly more. We use $32,561$ samples in our dataset.
Lung Cancer. The Zilionis dataset is widely used, and consists of single-cell RNA sequencing data. It has $306$ features, and $48,969$ samples. The data has 20 classes corresponding to cell type. More can be found at [32]
Animals with Attributes. The Animals with Attributes (AWA) dataset, contains 5,000 data points corresponding to 10 unseen classes, where the testing image features are obtained from the pre-trained ResNet architecture, with $D=2,048$ dimensions, and the semantic features are provided with $D=85$ dimensions. More information can be found in Section 4.1 of [29].
Figure 3: Clustering results using two moons
10 Experimental Results on More Datasets
Here we show further experiments on more datasets.
Dense cluster inside a sparse circle: In this example, we sampled $500$ points from a dense cluster (using a uniform distribution) situated in the interior of a sparse cluster (using 100 points), maintaining a ratio of 5:1. UMAP relies on a well-initialized kNN graph, however the disparate densities in the two manifolds lead to erroneous edge weights, resulting in errors in segmentation. Despite utilizing the same initial graph construction as UMAP, our method, employing a multi-scale representation, is more successful in segmenting the two manifolds accurately in the presence of differing densities. HeatGeo, a competitive state-of-the-art method, does the best here, and is able to separate the dense-sparse circles from each other.
Data Dense-sparse clusters
Method/Accuracy ARI AMI
UMAP 0.75 0.72
t-SNE 0.30 0.33
ISOMAP 0.74 0.65
Diffusion Maps 0.81 0.76
HeatGeo 1.0 1.0
Ms-IMAP Method 1 0.82 0.8
Ms-IMAP Tensor-encoding Method 2 0.77 0.73
Table 3: Comparison of clustering performance on Two Moons and Dense-sparse clusters datasets.
11 HeatGeo Hyperparameter Tuning Details
We tune HeatGeo on each dataset in Section 7, by tuning over the parameter space:
Hyperparameter Set of values
knn 5, 10, 15
lap_type normalized, combinatorial
harnack_regul 0, 0.5, 1
Table 4: Space of parameters in which we tuned HeatGeo.
12 Ablation Studies of Hyperparameters for MS-IMAP
Here we study what effect the hyperparameters – the number of nearest neighbors, and the number of filters – have on the clustering performance of MS-IMAP with method 2. In table 5, we demonstrate the effect of varying the number of neighbors, between 10, 15, and 20 neighbors. We see that the results are mostly the same, thus showing MS-IMAP is robust on real datasets.
Dataset Census Zilionis AwA
Number of Neighbors / Accuracy ARI AMI ARI AMI ARI AMI
10 0.22 0.15 0.71 0.78 0.73 0.80
15 0.22 0.15 0.70 0.76 0.74 0.81
20 0.22 0.15 0.70 0.77 0.71 0.79
Table 5: Ablation study on MS-IMAP with method 2, varying the number of neighbors. The number of filters is kept at 5.
We also study the effect of varying the number filters. In table 6, we also see similar results when using different filters, showing the stability of MS-IMAP.
Dataset Census Zilionis AwA
Number of Filters / Accuracy ARI AMI ARI AMI ARI AMI
5 0.22 0.15 0.70 0.76 0.74 0.81
6 0.22 0.15 0.70 0.76 0.73 0.81
7 0.22 0.15 0.72 0.78 0.74 0.81
Table 6: Ablation study on MS-IMAP with method 2, varying the number of filters. The number of neighbors is kept at 15.
13 Ablation Studies of Hyperparameters for t-SNE, Isomaps, and Diffusion Maps
We study the effect of varying the number of neighbors for the methods: t-SNE, Isomaps, and Diffusion Maps. For t-SNE this would be the perplexity hyperparameter, for Isomaps this would be the number of neighbors, and for diffusion maps, this would be the parameter that affects the width of the Gaussian kernel, i.e. $\exp(\cdot/\alpha)$.
In Table 7, we see choosing a smaller perplexity of 15 does worse than the perplexity of 30 and 60.
Dataset Census Zilionis AwA
Perplexity / Accuracy ARI AMI ARI AMI ARI AMI
15 0.03 0.04 0.37 0.67 0.69 0.76
30 0.15 0.15 0.38 0.68 0.73 0.80
60 0.17 0.17 0.38 0.69 0.73 0.80
Table 7: Ablation study on t-SNE, varying the perplexity.
In Table 8, we see a similar pattern to t-SNE, where choosing too low causes reduction in performance. But performance stabilizes around choosing the number of neighbors as 5-10+.
Dataset Census Zilionis AwA
Neighbors / Accuracy ARI AMI ARI AMI ARI AMI
2 -0.05 0.02 0.41 0.55 0.43 0.58
5 0.18 0.09 0.44 0.56 0.52 0.60
10 0.18 0.09 0.44 0.57 0.51 0.62
Table 8: Ablation study on Isomap, varying the number of neighbors.
In Table 9, we find the ablation study for Diffusion Maps. Diffusion Maps perform the worst, but as with t-SNE and Isomap, benefits from considering more neighbors.
Dataset Census Zilionis AwA
alpha / Accuracy ARI AMI ARI AMI ARI AMI
1 0 0 0 0.01 0.23 0.44
2 0 0 0 0.01 0.22 0.43
3 0.06 0.16 0 0.02 0.22 0.42
Table 9: Ablation study on Diffusion Map, varying alpha which affects the width of the kernel.
14 Fast computation using Chebyshev polynomials
We provide additional details regarding the fast computation of SGW coefficients [15]. Directly computing the SGW coefficients above requires calculating the entire eigensystem of the Laplacian, which is computationally intensive - $O(N^{3})$ for $N$ points. Instead, Hammond et al. [15] suggested computing the SGW using a fast algorithm based on approximating the scaled generating kernels through low-order polynomials. The wavelet coefficients at each scale are then computed as a polynomial of $\mathbf{L}$ applied to the input data, using approximating polynomials given by truncated Chebyshev polynomials.
The Chebyshev polynomials $T_{k}(y)$ are computed using the recursive relations: $T_{k}(y)=2yT_{k-1}(y)-T_{k-2}(y)$ for $k\geq 2$, where $T_{0}=1$ and $T_{1}=y$.
The SGW coefficients are then approximated using wavelet and scaling function coefficients as follows:
$\psi_{f}(s_{j},i)\sim\left(\frac{1}{2}c_{j,0}f+\sum_{k=1}^{K}c_{j,k}\bar{T}_{j,k}(\mathbf{L})f\right)_{i}$ (12)
where $c_{j,k},j>0$ are the Chebyshev coefficients and $\bar{T}_{j,k}$ are the shifted Chebyshev polynomials $\bar{T}_{k}(x)=T_{k}\left(\frac{x-a}{a}\right)$ for $x\in[0,\lambda_{\mbox{max}}]$, where $x=a(y+1)$, a = $\lambda_{\mbox{max}}$. The scaling function coefficients, which are corresponding to a low-pass filter operation, are approximated in a similar way using Chebyshev polynomials. Note that the scaling kernel function is a low pass filter $h$ satisfying $h(0)>0$ and $h(x)\rightarrow 0$ when $x\rightarrow\infty$. If the graph is sparse, we obtain a fast computation of the matrix-vector multiplication $\bar{T}_{j,k}(\mathbf{L})f$, where the computational complexity scales linearly with the number of points, resulting in a complexity of $O(N)$ for an input signal $f\in\mathbb{R}^{N}$. The SGWs efficiently map an input graph signal (a vector of dimension $N$) to $NK$ scaling and wavelet coefficients.
15 Runtime and Computational Complexity
We have performed experimental runtime studies on empirical datasets. Experiments were performed on Virtual Server containers with 32 cores Intel Xeon 8259CL running at 2.50Ghz and 256GB of RAM. On the AWA dataset of $5,685$ samples and $2,048$ features, Method 1 and the 3D based Tensor Method 2 both run in 1.35 mins. Finally on the Cancer QC data-set of 48,969 samples and 306 features, Method 1 took 12.44 mins. The computational complexity of Multi-Scale UMAP is of $O(NDlog(N))$ for construction of the multi-scale representations which includes the $k$ nearest neighbor graph using k-d tree, the SGW transform which is $O(N)$ for each dimension of the manifold for sparse graphs. The optimization stage has a complexity which scale with the number of edges in the graph, which has a complexity of $O(kDN)$.
16 Theoretical results: Sampling set for Smooth Manifolds with functions defined over Paley-Wiener Spaces
In this work, we characterize the theoretical properties of the representation power of the SGW operator by considering functions sampled from the Paley-Wiener spaces [22] on combinatorial graphs. The Paley-Wiener spaces were introduced on combinatorial graphs in [22] and a corresponding sampling theory was developed which resembles the classical one. Pesenson proved in [22] that Paley-Wiener functions of low type are uniquely determined by their values on certain subgraphs (which are composed from set of nodes known as the uniqueness sets) and can be reconstructed from such sets in a stable way. We demonstrate that the SGW operator can represent functions $f$ that reside in the Paley-Wiener space on combinatorial graphs more efficiently than the graph Laplacian operator $\mathcal{L}$. The effectiveness of the SGW operator representation in this case can be understood in several ways. In one way, by the ability of the SGW operator to accurately represent functions with larger bandwidth, i.e., $f\in PW_{\omega^{\prime}}(G)$ where $\omega<\omega^{\prime}$.
In order to prove these results, we will first need the following definitions:
Definition 1
The Paley-Wiener space of $\omega$bandlimited signals is defined as follows:
$PW_{\omega}(G)=\left\{f|\hat{f}(\lambda)=0\,\,\forall\,\,\lambda>\omega)\right\}$ (13)
We summarize the main notions and definitions. We consider simple undirected unweighted and connected graph $G=(V,\mathbf{W})$, where $V$ is its set of $N$ vertices and $\mathbf{W}$ is its set of edges. The degree of $v$ is number of vertices adjacent to a vertex $v$ is and is denoted by $d(v)$. We assume that degrees of all vertices are bounded by the maximum degree denoted as
$d(G)=\mbox{max}_{v\in V}d(v)$ (14)
The following definition [22] explains the uniqueness set.
Definition 2
A set of vertices $U\subset V$ is a uniqueness set for a space $PW_{\omega}(G)$ if we have two functions from $PW_{\omega}(G)$ that coincide on $U$ then they coincide on $V$.
Next the following definition of $L_{2}(S)$ is provided as the set of finite energy signals whose support is contained in $S$. The space $L_{2}(G)$ is the Hilbert space of all complex-valued functions $f:V\rightarrow\mathbb{C}$ with the following inner product $\left\langle f,g\right\rangle=\sum_{v\in G}f(v)\overline{g(v)}$ and the norm
$||f||=\left(\sum_{v\in V}{|f(v)|^{2}}\right)^{1/2}$ (15)
The Laplace (normalized) operator $\cal L$ is defined by the formula [22]:
$\mathcal{L}f(v)=\frac{1}{\sqrt{d(v)}}\sum_{u\sim v}\left(\frac{f(v)}{\sqrt{d(v)}}-\frac{f(u)}{\sqrt{d(u)}}\right),f\in L_{2}(G)$ (16)
Definition 3
For a subset $S\subset V$, denote $L_{2}(S)$ as the space of all functions from $L_{2}(G)$ with support in $S$:
$L_{2}(S)=\left\{\varphi\in L_{2}(G),\varphi(v)=0,v\in V(G)\setminus S\right\}$
Definition 4
[22] We say that a set of vertices $S\subset V$ is a $\Lambda$set if for any $\varphi\in L_{2}(S)$ it admits a Poincare inequality with a constant $\Lambda(S)>0$
$||\varphi||\leq\Lambda||\mathbf{\cal L}\varphi||,\,\varphi\in L_{2}(S)$ (17)
The infimum of all $\Lambda(S)>0$ for which $S$ is a $\Lambda$set will be called the Poincare constant of the set $S$ and denoted by $\Lambda$.
Definition above provides a tool to determine when bandlimited signals in Paley-Wiener spaces $PW_{\omega}(G)$ can be uniquely represented from their samples on a given set. The role of $\Lambda$-sets was explained and proved in the following Theorem by Pesenson [22], that shows that if $S\subset V$ then any signal $f\in PW_{\omega}(G)$ can be uniquely represented by its samples in the complement set $U=V(G)\setminus S$:
Theorem 2
[22] If $S\subset V$ is a $\Lambda$- set, then the set $U=V(G)\setminus S$ is a uniqueness set for any space $PW_{\omega}(G)$ with $\omega<\frac{1}{\Lambda}$.
Remark: Note that non-trivial uniqueness sets can not exist for functions from any Paley-Wiener subspace $PW_{\omega}(G)$ with any $\lambda_{0}\leq\omega\leq\lambda_{N}$, but can they can exist for some range $\lambda_{0}\leq\omega<\Omega$, as was shown in [22].
We state one of our main results, in which we employ the SGW operator $\psi$ to characterize the uniqueness set using the $\Lambda_{\psi}$-set, therefore extending the $\Lambda$-set concerning the graph Laplacian operator $\cal L$.
Theorem 3
Let $G=(V,\mathbf{W})$ be a connected graph with $N$ vertices. Assume that there exist a set of vertices $S\subset V$ for which the conditions (1)-(2) in Lemma 1 hold true. Let $\psi$ be the SGW operator using a polynomial $p(\mathbf{\cal L})$ with the coefficients $\left\{a_{k}\right\}_{k=0}^{K}$ such that $\psi_{f}=\sum_{k=0}^{K}a_{k}\mathbf{\cal L}^{k}f$. Then, for any $\varphi\in L_{2}({S})$, we have that the following inequality holds:
$||\varphi||\leq\Lambda_{\psi}||\psi_{{\varphi}}||$ (18)
and thus the set $S$ is a $\Lambda_{\psi}$-set for the operator $\psi$ with $\Lambda_{\psi}={\frac{1}{\sqrt{\sum_{k=0}^{K}\frac{a_{k}^{2}}{\Lambda^{2k}}}}}$.
We recall the following results from [22]. Note that [22] established the construction of a $\Lambda$-set by imposing specific assumptions on the sets $S$ and $U$. Our result Theorem 3 holds similar assumptions as in [22].
Lemma 1
[22] Given a connected graph $G=(V,\mathbf{W})$ graph with $N$ vertices for which the following conditions hold true: (1) For every $s\in S$ there exists $u\in U$ that is a neighbor of $s$, i.e., $w(u,s)>0$.
(2) for every $s\in S$ there exists at least one neighbor node $u\in U$ whose adjacency set intersects $S$ only over $s$.
Then there exist a set of vertices $S\subset V$ which is a $\Lambda$-set, with $\Lambda=d(G)$.
In the next theorem, we expand the characterization of the uniqueness set using the $\Lambda$-set concerning the graph Laplacian operator $\mathcal{L}$ to include cases where we employ the SGW operator $\psi$. We thereby characterize the uniqueness set using the $\Lambda_{\psi}$-set for the SGW operator.
We now turn to prove Theorem 3, which was stated earlier.
Proof of Theorem 3:
Assuming $f\in PW_{\omega}(G)$, $f$ can be efficiently represented using a polynomial $p(\mathcal{L})$ instead of a kernel function $g$, where the wavelet at node $i$ and a fixed scale $s$ is calculated using the polynomial $p(\mathbf{\cal L})$ of the Laplacian:
$\psi(i)=\sum_{k=0}^{K}a_{k}\mathbf{\cal L}^{k}\delta_{i}$ (19)
and we can write the SGW coefficients with respect to the function $\varphi$ as:
$\psi_{{\varphi}}=\sum_{l=0}^{N-1}\sum_{k=0}^{K}a_{k}\mathbf{\cal L}^{k}\varphi$
Following the assumptions of Theorem 1 implies that there exists a subset $U^{*}\subset U$ such that for every $s\in S$ there exists at least one point $u^{*}\in U^{*}$ whose adjacency set intersects $S$ only over $s$ and from this property we have
$\mathbf{\cal L}\varphi(u^{*})=-\frac{\varphi(s)}{\sqrt{d(s)d(u^{*})}},u^{*}\in U^{*},s\in S$ (20)
We will use this property to show that the set $S$ is a $\Lambda_{\psi}$-set with respect to the operator $\psi$. Using 20 and taking taking powers of the Laplacian operator considering $u^{*}\in U^{*},s\in S$ we have that:
$\mathbf{\cal L}^{k}\varphi(u^{*})=(-1)^{k}\frac{\varphi(s)}{{(d(s)d(u^{*}))^{k/2}}},u^{*}\in U^{*},s\in S$ (21)
and
$|\mathbf{\cal L}^{k}\varphi(u^{*})|\geq\frac{|\varphi(s)|}{\Lambda^{k}},u^{*}\in U^{*},s\in S$ (22)
Thus we have that $\forall\varphi\in L_{2}(S)$
$||\psi_{\varphi}||=\left(\sum_{i,j}\sum_{k=0}^{K}|a_{k}({\cal L}^{k}\varphi)_{ij}|^{2}\right)^{1/2}\geq\left(\sum_{s\in S}\sum_{k=0}^{K}|a_{k}({\cal L}^{k}\varphi(u^{*}))|^{2}\right)^{1/2}$ (23)
using similar argument as in Lemma 1 we have that
$\left(\sum_{k=0}^{K}\sum_{s\in S}|a_{k}{\cal L}^{k}\varphi(u^{*})|^{2}\right)^{1/2}=\left(\sum_{k=0}^{K}|a_{k}|^{2}\sum_{s\in S}|{\cal L}^{k}\varphi(u^{*})|^{2}\right)^{1/2}\geq||\varphi||\left(\sum_{k=0}^{K}|a_{k}|^{2}\frac{1}{\Lambda^{2k}}\right)^{1/2}$ (24)
which proves the claim of the Theorem. $\square$
Remark 1: An important property which can be observed from Theorem 3 is the following: given the SGW $\psi_{{f}}$ as a spectral representation operator, we can choose coefficients $\left\{a_{k}\right\}_{k=0}^{K}$ such that we obtain a $\Lambda_{\psi}$ set associated with the operator $\psi$, which is smaller than the $\Lambda$ set associated with the Laplacian operator $\cal L$. This implies that the operator $\psi$ provides more flexibility and better control over smoothness properties in comparison to the Laplacian operator.
Remark 2: Since the $\Lambda_{\psi}$ set can be chosen to be smaller than $\Lambda$ set (for a proper choice of the coefficients $\left\{a_{k}\right\}_{k=0}^{K}$ using the operator $\psi$) then the SGW operator $\psi$ provides a more efficient representation for $f\in PW_{\omega^{\prime}}(G)$ with $\omega<\omega^{\prime}$ using the same subsets of nodes from the uniqueness set $U$ in comparison to the Laplacian operator $\cal L$.
Remark 3: Note that the characterization of the the uniqueness set does not rely on a reconstruction method of the graph signal values of $f(S)$ from their known values on $U$.
The next Theorem demonstrates the role of the $\Lambda_{\psi}$set with respect to the operator $\psi$, where we show that any signal $f\in PW_{\omega}(G)$ can be uniquely represented by its samples on the uniqueness set $U$. This results resembles the role of $\Lambda$-sets with respect to the graph Laplacian operator $\mathbf{\cal L}$, yet with a different bound then Lemma 2.
Theorem 4
Let $G=(V,\mathbf{W})$ be a connected graph with $N$ vertices and $f\in PW_{\omega}(G)$ for $\lambda_{1}<\omega<\lambda_{\mbox{max}}$. Given the SGW operator $\psi$, and a set $S$ which is a $\Lambda_{\psi}$ set. Then the set $U=V\setminus S$ is a uniqueness set for any space $PW_{\omega}(G)$ with $\omega<1/\Lambda_{\psi}$.
Proof: Given $f,\tilde{f}\in PW_{\omega}(G)$, then $f-\tilde{f}\in PW_{\omega}(G)$. Assume that $f\neq\tilde{f}$. If $f,\tilde{f}$ coincide on $U=V\setminus S$, then $f-\tilde{f}\in L_{2}(S)$ and therefore
$||f-\tilde{f}||\leq\Lambda_{\psi}||\psi_{f-\tilde{f}}||$ (25)
Since $\psi_{{f}}\in\mathbb{R}^{N}$, we have that by properties of a vector space in $\mathbb{R}^{N}$, using the Cauchy–Schwarz inequality and assuming $|a_{k}|\leq 1\,\forall k$, we have:
$||\psi_{f-\tilde{f}}||\leq\omega||f-\tilde{f}||$ (26)
Combining the inequalities above and using the inequality $\Lambda_{\psi}\omega<1$ we have that
$||f-\tilde{f}||\leq\Lambda_{\psi}||\psi_{f-\tilde{f}}||\leq\Lambda_{\psi}\omega||f-\tilde{f}||<||f-\tilde{f}||$ (27)
which is a contradiction to the assumption that $f\neq\tilde{f}$. Thus, the set $U=V\setminus S$ is a uniqueness set for any space $PW_{\omega}(G)$ with $\omega<1/\Lambda_{\psi}$. $\square$
Remark 1: Note that $\Lambda\omega<1$ implies that $\Lambda_{\psi}\omega<1$ given $\Lambda_{\psi}<\Lambda$, then we can increase the size number of nodes in the uniqueness set $U$ for $PW_{\omega}(G)$ (for a proper choice of the coefficients $\left\{a_{k}\right\}_{k=0}^{K}$ using the operator $\psi$). In other words, we may increase the size of $S$ (thus reducing the size of $U$) and still obtain a uniqueness set with a smaller size for the graph signals in $PW_{\omega}(G)$.
Remark 2: We note that the results of Theorem 4 concerning the uniqueness set are independent from the stability properties of the representation. In order to achieve stability which is important for reconstruction, it is required to construct a wavelet operator using multiple scales $t_{j},j=1,...K$, as proposed in [15], which will yield a collection of $nK$ wavelets. More specifically, we can express the function $\varphi$ using multiple scales $t_{j}$ (here we replace the previous notation of scale $s_{j}$ with $t_{j},j=1,..K$, not to confuse with nodes $s\in S$). Then for a fixed scale $t_{j}$ we have that the SGW is given by $||\psi_{\varphi}(t_{j})||=\left(\sum_{i=1}^{N}\sum_{k=0}^{K}|a_{t_{j},k}({\cal L}^{k}\varphi(i))|^{2}\right)^{1/2}$ and
$||\psi_{\varphi}||_{=}\left(\sum_{j}\sum_{i=1}^{N}\sum_{k=0}^{K}|a_{t_{j},k}({\cal L}^{k}\varphi(i))|^{2}\right)^{1/2}$ (28)
In a similar way to the arguments provided in Theorem 3 we can choose coefficients $a_{t_{j},k}$ associated with the Laplacian polynomial such that the inequality $||\varphi||\leq\Lambda_{\psi}||\psi_{{\varphi}}||$ is satisfied.
17 Sampling approach from network features
Our approach revolves around the strategic sampling of edges, emphasizing their importance in the overall network structure. We introduce innovative method for this purpose, such as one leveraging edge betweenness centrality (EBC). In these proposed methods, our approach to edge sampling is guided by the assessed significance of edges with respect to the topological structure of the graph. This methodology extends random sampling techniques commonly employed in contemporary graph embedding, which typically rely on information derived solely from sparse graph connectivity, encompassing nodes within a 1-hop distance on the graph. To elaborate on our sampling approach, we extend the definition of the network $G=(V,\mathbf{W})$ to a triplet $G=(V,\mathbf{W},\gamma)$ such that $\gamma:V\rightarrow(0,1]$ is some probability distribution over $V$ (or $\mathbf{W}$) providing the significance of each node/edge. We choose to estimate $\gamma$ directly from the graph network using kernel density estimation (KDE).
The kernel density estimator of the distribution ${\gamma}({{V})}$, is given by
$\hat{\gamma}(x)=\frac{1}{Nh}\sum_{i\in V}{\frac{\mathbf{K}(\bar{v}(i)-x)}{h}}$ (29)
Where $\bar{v}$ accumulate statistics measuring the extent of diffusion spread using SGW among nodes, $\mathbf{K}$ is symmetric kernel function (chosen to be a smooth Gaussian symmetric function), $h>0$ is the smoothing bandwidth that controls the amount of smoothing.
In a similar way, one can also construct a probability measure $\gamma:V\times V\rightarrow(0,1]$ describing the significant of an edge. In this case, we choose to calculate the distribution from the edge betweenness centrality measurements (EBC), $\left\{w_{i,j}^{BE}|i,j\in V,i\sim j\right\}$ (see Eq. 31) where we use kernel density estimation (KDE) to estimate the density distribution. Let $\gamma:V\times V\rightarrow\mathbb{R}^{+}$ denote the probability distribution over edges in the graph $G=(V,\mathbf{W})$, using the distribution of EBC, (EBC is defined in Eq. 30). We denote the estimated distribution calculated directly from the EBC measurements as $\gamma({\mathbf{W^{BE}})}$. Formally, the kernel density estimator of distribution of EBC, ${\gamma}({\mathbf{W^{BE}})}$ is given by
$\hat{\gamma}(x)=\frac{1}{N(e)h}\sum_{i\sim j}{\frac{\mathbf{K}(w_{i,j}^{BE}-x)}{h}}$ (30)
Where $N(e)$ is a fixed number of positive edges to be sampled.
The advantage of using KDE is that it is a non-parametric density estimator, which does not require assumption that the underlying density function is from a parametric family. Note that KDE works by estimating the density at each point (node or edge) as a weighted sum of the densities of neighboring points, with the weights given by a kernel function $\mathbf{K}$. In our experiments, we used the Gaussian kernel for simplicity, but one can use other kernel choices.
Figure 4: Sampling from Betweenness Centrality distribution
17.1 Sample from Network Features using Edge Betweenness Centrality
Motivation: We employ Edge Betweenness Centrality (EBC) based sampling to create a strategy that samples edges from the EBC distribution. This enables us to optimize embeddings by predominantly using edges with low EBC, which are concentrated in dense clusters. In contrast, high-EBC edges typically occur in transitions between clusters and form a sparser set that is less frequently sampled. During optimization with SGD and negative sampling, these high-EBC edges play a crucial role in revealing the global structure of the network and connections between different clusters.
Consider the illustration in Fig. 4, where edges marked with black arrows correspond to edges with high edge betweenness centrality values, indicating their importance in connecting different dense clusters. Sampling edges with low EBC assists in identifying dominant local structures, while infrequently sampling high betweenness centrality edges is critical for understanding large clusters and the network’s global structure.
  Input: $G=(V,\mathbf{W})$, kernel density function $\mathbf{K}$, bandwidth parameter $h$, $N_{e}$ for the number of edge samples.
  Step 1: Construct the centrality graph $G=(V,\mathbf{W}^{BE})$ using Eq. (31).
  Step 2: Utilize $\mathbf{K}$, $\bar{v}$, and $\mathbf{W}^{BE}$ in Kernel Density Estimation (Eq. 30) to estimate $\hat{\gamma}$.
1  Step 3: Initialize a list of positive edges $C_{e}=\left\{\right\}$. Set $i=1$.
  Step 4: While $i\leq N_{e}$ do:
  Step 5: Sample an edge $w_{ij}$ using $\hat{\gamma}$ , $w_{ij}\sim\hat{\gamma}(\mathbf{W}^{BE})$.
  Step 6: Add $w_{ij}$ to $C_{e}$, and and set $i=i+1$.
  Step 7: Repeat steps 4-6. STOP when $i=N_{e}$
  Output: Set of edges $C_{e}$ for positive samples in Algorithm 1.
Algorithm 2 Edge Betweenness Centrality Sampling
The Edge Betweenness Centrality:
Edge Betweenness Centrality (EBC) quantifies edge importance based on the fraction of shortest paths that pass through it. For a graph with vertices and edge weights, the EBC of an edge $(i,j)$ is given by:
$w_{i,j}^{BE}=\sum_{k\neq t}\frac{\sigma_{kt}(w_{ij})}{\sigma_{kt}}$ (31)
Here, $\sigma_{kt}$ is the number of shortest paths from node $k$ to node $t$, and $\sigma_{kt}(w_{ij})$ is the number of those paths that include the edge $w_{ij}$. The centrality graph $G=(V,\mathbf{W}^{BE})$ is constructed using these EBC values, modifying the edges of the original graph $G$.
Sampling Strategy of EBC using KDE estimator:
Let $\gamma:V\times V\rightarrow\mathbb{R}^{+}$ denote the probability distribution over edges in the graph $G=(V,\mathbf{W})$, using the distribution of EBC. We denote the estimated distribution calculated directly from the EBC measurements as $\gamma({\mathbf{W^{BE}})}$.
Given the set of edge betweenness centrality $\left\{w_{i,j}^{BE}|i,j\in V,i\sim j\right\}$ as an input to the Kernel Density Estimation (KDE) estimator, the density function is estimated automatically. With a fixed number of positive edges $N(e)$ to be sampled, we sample $N(e)$ edges from the estimated distribution. Although we used the Gaussian kernel in our experiments, other kernel choices are viable. In EBC based-sampling, we sample from the distribution of edge betweenness centrality measurements. This approach prioritizes the sampling of edges based on their importance in the graph structure. By treating edges as bottlenecks, we sample high betweenness centrality edges less frequently, thereby revealing clusters. Edge betweenness centrality serves as a measure of an edge’s bottleneck potential, guiding our SGD sampling and prioritizing edges based on their relevance and significance in capturing underlying relationships and connectivity patterns in the graph. Combined with the proposed encoding and initial embedding of multi-scale graph structure, the approach provides a comprehensive representation of the data.

### v4

@@
 MS-IMAP - A Multi-Scale Graph Embedding Approach for Interpretable Manifold Learning
-MS-IMAP - A Multi-Scale Graph Embedding Approach for Interpretable Manifold Learning
-Shay Deutsch  [^1] Corresponding author.  [^4] Emerging Capabilities Research Group, Discover Financial Services Inc., Riverwoods, IL.
-shaydeu@gmail.com
-   Lionel Yelibi  [^4] Emerging Capabilities Research Group, Discover Financial Services Inc., Riverwoods, IL.
-ylblio001@myuct.ac.za
-   Alex Tong Lin  [^4] Emerging Capabilities Research Group, Discover Financial Services Inc., Riverwoods, IL.
-atlin271@gmail.com
-   Arjun Ravi Kannan  [^4] Emerging Capabilities Research Group, Discover Financial Services Inc., Riverwoods, IL.
-arjun.kannan@gmail.com
-1 Abstract
+Abstract
 Deriving meaningful representations from complex, high-dimensional data in unsupervised settings is crucial across diverse machine learning applications.
 This paper introduces a framework for multi-scale graph network embedding based on spectral graph wavelets that employs a contrastive learning approach.
-A significant feature of the proposed embedding is its capacity to establish a correspondence between the embedding space and the input feature space which aids in deriving feature importance of the original features.
-We theoretically justify our approach and demonstrate that, in Paley-Wiener spaces on combinatorial graphs, the spectral graph wavelets operator offers greater flexibility and better control over smoothness properties compared to the Laplacian operator.
-We validate the effectiveness of our proposed graph embedding on a variety of public datasets through a range of downstream tasks, including clustering and unsupervised feature importance.
+We theoretically show that in Paley-Wiener spaces on combinatorial graphs, the spectral graph wavelets operator provides greater flexibility and control over smoothness compared to the Laplacian operator, motivating our approach.
+An additional key advantage of the proposed embedding is its ability to establish a correspondence between the embedding and input feature spaces, enabling the derivation of feature importance.
+We validate the effectiveness of our graph embedding framework on multiple public datasets across various downstream tasks, including clustering and unsupervised feature importance.
 1 Introduction
@@
 However, limitations of these techniques include reliance on low frequencies of the graph Laplacian and $k$-Nearest Neighbors graph connectivity, which provide limited information, including in recent contrastive learning methods such as Uniform Manifold Approximation and Projection (UMAP).
-To address these limitations, this paper proposes three contributions.
-(1) A framework employing multi-scale graph representation using contrastive learning which enhances the expressiveness of an embedding by optimizing low and high-frequency information, capturing intricate data details.
+To address these limitations, this paper proposes three main contributions.
+(1) A framework employing multi-scale graph representation using 3D based tensor and optimization using contrastive learning which enhances the expressiveness of an embedding by optimizing low and high-frequency information, capturing intricate data details.
+(2) An application of manifold interpretation to feature importance allows each dimension of our proposed embedding to be uniquely traced back and associated with the original features.
 (3) A characterization of the theoretical properties of the representation power of the spectral graph wavelets (SGW) operator by considering functions sampled from the Paley-Wiener spaces [27] on combinatorial graphs and a demonstration of the flexibility and better control over smoothness properties of the SGW operator in comparison to the Laplacian operator.
-(3) A novel sampling technique that leverages the graph structure by sampling edges or nodes from a distribution reflecting network characteristics.
 A key distinction between our approach and many manifold learning methods such as UMAP is that we leverage multi-scale graph representation (using SGW) to increase the dimensionality of the data thus enhancing data representation learning.
-On the other hand, other manifold learning techniques predominantly employ nonlinear dimensionality reduction to reduce the dimensions of the output embedding space.
 We use stochastic gradient descent (SGD) to optimize the embedding space and employ a novel contrastive learning approach using a 3D-tensor to capture both low and high frequencies.
-Another limitation of non-linear manifold learning methods is they inherently lack direct connections to input features, thereby limiting their ability to offer interpretable results which is crucial in various applications (e.g.
-in finance where customer behavior needs to be summarized with a few important features to design products or strategies benefiting them and in biology where easily interpretable features may lead to life saving discoveries).
-Although not the main focus of this article, our proposed approach establishes a mapping between the original features and the constructed optimized embeddings and enables the measurement of the importance of the original features with respect to the embedding space (see Figure.1 for an illustration of the mapping generated by the proposed embedding).
+Generally, manifold learning methods perform nonlinear dimensionality reduction by embedding the data into lower-dimensional spaces, although some approaches may embed the data into higher dimensions [16, 31].
+A common limitation of these and other graph embedding methods is their lack of direct connections to the input features, which restricts their ability to produce interpretable results.
+This can be crucial in fields like finance, where behavioral analysis must be summarized using key features to guide decisions, or in biology, where interpretable features are essential for understanding complex systems.
+Our proposed approach establishes a mapping between the original features and the constructed optimized embeddings and enables the measurement of the importance of the original features with respect to the embedding space.
 To the best of our knowledge, current manifold learning techniques produce embeddings lacking explicit interpretation.
@@
 SGW are constructed by applying a set of bandpass filters to the eigenvalues of the graph Laplacian, enabling multi-resolution representation of graph signals.
-An alternative multi-scale framework for analyzing functions on graphs is Diffusion Wavelets [23].
-The Geometric Graph Scattering transform and the Graph Scattering Transform [14] utilize Diffusion Wavelets to extract graph features, which are then employed in graph classification tasks.
-In our work, we leverage the localization properties of SGW to develop a novel feature representation scheme, achieving superior performance in our target applications.
-4 Preliminaries
+Also related to our approach are graph embedding and multi-scale framework using Diffusion Wavelets [28] that embed the data in higher dimensions such as the Geometric Scattering for Graph Data Analysis [17] and the Graph Scattering Transform [16] that extract graph features, which are then employed in graph classification tasks.
+2.1 Preliminaries
 Consider a set of points $\mathbf{x}=\left\{\mathbf{x}_{i}\right\},\,i=1,...N,\mathbf{x}_{i}\in\mathbb{R}^{D}$, which are sampled from an unknown manifold $M\subset\mathbb{R}^{D}$.
@@
 The adjacency matrix $\mathbf{W}=(w_{ij})$ consists of the weights $w_{ij}$ between node $i$ and node $j$.
-The weights can be chosen using different techniques, such as the Gaussian kernel function or adaptive graph construction.
-If the weights are chosen using the Gaussian kernel function then
-$(\mathbf{W})_{ij}=\left\{\begin{matrix}\exp\left(\frac{-||\mathbf{x}_{i}-\mathbf{x}_{j}||_{2}^{2}}{2\sigma_{D}^{2}}\right)&\mbox{if}\>\mathbf{x}_{j}\in\text{kNN}(\mathbf{x}_{i})\\ 0&\mbox{otherwise}\end{matrix}\right.$ (1)
-where is (||||)(||\,||)( | | | | ) denotes the $L2$ distance between the points $\mathbf{x}_{i},\mathbf{x}_{j}$, kNN$(\mathbf{x}_{i})$ denotes the $k$ nearest neighbors of $\mathbf{x}_{i}$, and $2\sigma_{D}^{2}$ are parameters used to construct the graph.
-An alternative is to use adaptive graph construction as proposed in [21].
-The global smoothness of the graph signal function ${f}\in\mathbb{R}^{N}$ (a function over the vertices of the graph $G$) is defined using the graph Laplacian quadratic form as follows:$||\bigtriangledown{f}||^{2}=\sum_{V(i,j)}{w_{ij}(f(i)-f(j)})^{2}={f}^{T}\mathbf{L}{f}$, where $\mathbf{L}$ denotes the combinatorial graph Laplacian, defined as $\mathbf{L}=\mathbf{D}-\mathbf{W}$, with $\mathbf{D}$ the diagonal degree matrix with entries $d_{ii}=d(i)$.
+The weights can be chosen using different techniques, such as the Gaussian kernel function between the points $\mathbf{x}_{i},\mathbf{x}_{j}$ using the $k$ nearest neighbors of $\mathbf{x}_{i}$, kNN$(\mathbf{x}_{i})$ (which denotes the $k$ nearest neighbors) or adaptive graph construction as proposed in [25].
+$\mathbf{L}$ denotes the combinatorial graph Laplacian, defined as $\mathbf{L}=\mathbf{D}-\mathbf{W}$, with $\mathbf{D}$ the diagonal degree matrix with entries $d_{ii}=d(i)$.
 The degree $d(i)$ of vertex $i$ is defined as the sum of weights of edges that are connected to $i$.
@@
 The eigenvalues and eigenvectors of $\mathbf{L}$ are $\lambda_{1},\ldots,\lambda_{N}$ and $\mathbf{\phi}_{1},\ldots,\phi_{N}$, respectively.
-The graph Fourier transform (GFT) $\hat{f}$ of the graph signal $f$ is defined as the expansion of $f$ in terms of the eigenvectors $\phi$ of the graph Laplacian, so that for frequency $\lambda_{l}$ we have:
-$\hat{f}(\lambda_{l})=\sum_{i}{f(i)\phi_{l}^{*}(i)}.$ (2)
+In this work we use the normalized Laplacian $\mathbf{L}_{N}$, which from now on we refer to as $\mathbf{L}$ for simplicity.
 We will denote the matrix representations of the eigenvectors and eigenvalues of the graph Lapacian $\mathbf{L}$ as $\Phi,\lambda$, respectively.
@@
 Spatial localization in the graph domain is implicitly controlled by a single wavelet scale parameter defined in the spectral domain, such that the more the vertex is localized in the vertex domain, the spectral band is wider.
-The scale parameter enables the model to adjust the effective neighborhood sizes to the properties of the data.
 To represent a signal ${f}\in\mathbb{R}^{N}$ in multiple scales $S=[s_{1},s_{2},..s_{K}]$, the SGW transform is constructed as follows: Assume that $g(\lambda)$ is a filter in the spectral domain.
 Let $\delta_{i}\in\mathbb{R}^{N}$ be the delta function centred at the vertex $i\in G:\delta_{i}(j)=1\,\,\mbox{if}\,\,i=j$, and $\delta_{i}(j)=0$ otherwise.
-Given a bandpass filter $g(\lambda)$ and a wavelet centered on node $i$ at scale $s$, the wavelets $\psi(s,i),i=1,..N$ are calculated by applying them to the delta function on a single vertex $i$, given by
-$\psi(s,i)=\Phi g(s\lambda)\Phi^{T}\delta_{i}$ (3)
+Given a bandpass filter $g(\lambda)$ and a wavelet centered on node $i$ at scale $s$, the wavelets $\psi(s,i),i=1,..N$ are calculated by applying them to the delta function on a single vertex $i$, given by $\psi(s,i)=\Phi g(s\lambda)\Phi^{T}\delta_{i}$.
 The value of $\psi(s,i)$ with respect to a vertex $m$ can be written as $\psi(s,i)(m)=\sum_{l=1}^{N}g(s\lambda_{l})\phi_{l}^{*}(i)\phi_{l}(m)$.
@@
 $\psi_{f}(s,i)=\sum_{l=1}^{N}g(s\lambda_{l})\hat{f}(\lambda_{l})\phi_{l}(i)$ (1)
-Fast computation using Chebyshev polynomials: Directly computing the SGW coefficients requires calculating the entire eigensystem of the Laplacian, which is computationally intensive - $O(N^{3})$ for $N$ points.
+Fast computation using Chebyshev polynomials: Directly computing the SGW coefficients is computationally intensive - $O(N^{3})$ for $N$ points.
 Instead, Hammond et al.
-[15] suggested computing the SGW using a fast algorithm based on approximating the scaled generating kernels through low-order polynomials of $\mathbf{L}$ applied to the input data (details in Appendix).
+[18] suggested computing the SGW using a fast algorithm based on approximating the scaled generating kernels through low-order polynomials of $\mathbf{L}$ applied to the input data (details in supplementary material).
 3 Our Proposed framework: Multi-Scale IMAP
-Figure 1: The encoding step of our framework is illustrated as a mapping between each coordinate of the input features and a corresponding dimension in the embedding space.
-This mapping is facilitated through our proposed approach, which aligns the input dimensions of the features with the latent embedding space.
-By establishing this correspondence, we acquire the capability to interpret each dimension in the embedding space.
-Such interpretation allows for various analyses, including providing feature importance relative to the latent space.
 In this section, we introduce Multi-scale IMAP, a framework for interpretable embedding via manifold learning utilizing a multi-scale graph representation.
@@
 Our method imposes a differentiable structure on the mapping $h:\mathbf{x}\rightarrow\psi_{\mathbf{x}}$, supported on a discrete graph $G=(V,\mathbf{W})$, where $\psi_{\mathbf{x}}$ represents the encoded multi-scale graph transform of $\mathbf{x}$.
-Multi-scale IMAP consists of three main components:
+Multi-scale IMAP consists of two main components:
 Step 1: Features representation encoding using multi-scale graph representation.
 In this step, our approach constructs a multi-scale graph representation by incorporating feature signals through the SGW Transform across multiple scales and graph frequencies.
-We introduce two methods for encoding multi-scale transform: in the first, based on simple concatenation, all of the features transformed for all scales are represented by concatenating all the SGW coefficients for each point in a vector form.
-In the second based on 3D tensor optimization, we construct a 3D tensor with dimensions ${K}\times{N}\times{D}$.
+We suggest a 3D tensor-based approach (detailed in Algorithm 1), where we construct a 3D tensor with dimensions ${K}\times{N}\times{D}$ to provide an effective representation for encoding multi-scale transform.
 We leverage the encoded 3D tensor structure to align the optimized embedding with all scales simultaneously to enforce differentiable structure of the transformed features.
+Note that an alternative would be to use simple concatenation, all of the features transformed for all scales are represented by concatenating all the SGW coefficients for each point in a vector form.
+For completeness we also describe this alternative in the Appendix.
+The 3D tensor-based approach (detailed in Algorithm 1) provides the most effective representation for the following reasons:
+(i) Joint and simultaneous optimization: using the 3D tensor-based method enables the joint and simultaneous optimization of all scales for each sample by leveraging the SGW coefficients of each spectral band within the cross-entropy loss framework while adhering to the fixed graph structure (Equation (3.2)).
+This approach effectively balances coarse and fine-grain resolutions, with summation performed only after the contrastive learning optimization process is complete.
+By smoothing both low- and high-frequency information jointly and simultaneously with respect to the graph structure, this method reduces the reliance on threshold-based techniques for removing high-frequency spectral bands, which may be less efficient.
+Jointly optimizing across scales ensures consistency in the optimization process for all scales with respect to a fixed pair of nodes.
+Although there may be correlations between overlapping SGWs corresponding to different spectral bands, each scale is associated with a distinct bandwidth of frequencies, effectively capturing graph structures at varying levels of neighborhood information.
+By integrating multiple scales, SGWs capture neighborhood information continuously across varying granularities, enabling a representation that strikes a balance between local and global structural details.
+(ii) Enhanced Interpretability: the proposed representation establishes a one-to-one correspondence between input features and the final embedding features, making feature interpretation more straightforward and intuitive.
+This is demonstrated in Section 6, where we show that the embeddings produced enable effective feature importance rankings.
+By ensuring this direct correspondence, our approach facilitates meaningful interpretation of the embedding space.
+This combination of optimization efficiency and enhanced interpretability highlights the clear advantages of the 3D tensor-based representation.
 Step 2: Optimization design using multi-scale network structure.
@@
 This integration leads to fine-grained manifold regularization and improved robustness in downstream tasks.
-Step 3: Network features sampling strategy.
-Our approach introduces a sampling scheme that selects network features based on their estimated importance in the graph network.
-Employed within the optimization using SGD, this adaptive strategy dynamically prioritizes and focuses on informative parts of the graph during the embedding process.
-This adaptability enhances the optimization process, leading to increased efficiency in generating embeddings.
-5.1 Feature Representation to encode multi-scale structure in the optimized embeddings
-We propose two alternative methods to encode multi-scale representations used for subsequent optimization.
-Note that for both method 1 and method 2, each dimension in the embedding space is constructed using a single feature in the original feature set, which is an essential characteristic that can be leveraged for interpreting graph embeddings.
+3.1 Step 1: Feature Representation to encode multi-scale structure
+To encode multi-scale representations used for subsequent optimization we focus on a 3D tensor based method.
+Note that each dimension in the embedding space is constructed using a single feature in the original feature set, which is an essential characteristic that can be leveraged for interpreting graph embeddings.
+3D-Tensor Encoding: The 3D-Tensor Encoding method concatenates the multi-scale representation for all features at a fixed scale, generating a matrix representation $\psi_{\mathbf{x}}(s_{j},:,:)\in\mathbb{R}^{D}\times\mathbb{R}^{N}$ for each scale $s_{j}$.
+After concatenating all $f_{i}$ at a fixed scale $s_{j}$, we have
+$\psi_{\mathbf{x}}(s_{j},:,:)=\psi_{{f}_{1}}(s_{j},:)\|\psi_{{f}_{2}}(s_{j},:)\|...\|\psi_{{f}_{D}}(s_{j},:)$
 […]
-$\mathbf{c}(\psi_{{f}_{i}}(s_{k},:),\psi_{{f}_{j}}(s_{k},:))=\psi_{{f}_{i}}(s_{k},:)\left|\right|\psi_{{f}_{j}}(s_{k},:)$
-denoting the concatenation of the vectors corresponding to the multi-scale representation $\psi_{{f}_{i}}(s_{k},:)$ and $\psi_{{f}_{j}}(s_{k},:)$.
-For method (i) (also noted as method 1) where all features and all scales are concatenated together, the resulting matrix $\psi_{\mathbf{x}}$ can be represented as
-$\psi_{\mathbf{x}}=\psi_{{f}_{1}}(s_{1},:)\|\psi_{{f}_{2}}(s_{1},:)\|....\|\psi_{{f}_{D-1}}(s_{K},:)\|\psi_{{f}_{D}}(s_{K},:)$
-where $\psi_{\mathbf{x}}\in\mathbb{R}^{KD}\times\mathbb{R}^{N}$.
-(ii) 3D-Tensor Encoding - Method 2 The second encoding method concatenates the multi-scale representation for all features at a fixed scale, generating a matrix representation $\psi_{\mathbf{x}}(s_{j},:,:)\in\mathbb{R}^{D}\times\mathbb{R}^{N}$ for each scale $s_{j}$.
-For method 2, after concatenating all $f_{i}$ at a fixed scale $s_{j}$, we have
-$\psi_{\mathbf{x}}(s_{j},:,:)=\psi_{{f}_{1}}(s_{j},:)\|\psi_{{f}_{2}}(s_{j},:)\|...\|\psi_{{f}_{D}}(s_{j},:)$
+$\mathbf{c}(\psi_{{f}_{i}}(s_{k},:,:),\psi_{{f}_{j}}(s_{k},:,:))=\psi_{{f}_{i}}(s_{k},:,:)\left|\right|\psi_{{f}_{j}}(s_{k},:,:)$
 The optimization can be performed separately for each scale $s_{j}$, or jointly for all scales $s_{j}$ using the 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$ We detail the proposed feature representation encoding methods in the pseudo code algorithm 1.
@@
 3Compute $\psi_{{f}_{r}}(s_{j},:)$ at scales $s_{j},j=1,...K$ using Chebyshev approximation.
-4If Encoding Method 1: Concatenate $\psi_{{f}_{r}}(s_{j},:)$, $j=1,...K$, $r=1,...,D$ in a matrix form $\psi_{\mathbf{x}}\in\mathbb{R}^{KD}\times\mathbb{R}^{N}$.
-If Encoding Method 2: Concatenate $\psi_{{f}_{r}}(s_{j},:)$ for $r=1,...,D$ in a matrix form $\psi_{\mathbf{x}}(s_{j},:,:)$ for each fixed scale $s_{j}$, representing all scales using 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$.
+  Step 5: Concatenate $\psi_{{f}_{r}}(s_{j},:)$ for $r=1,...,D$ in a matrix form $\psi_{\mathbf{x}}(s_{j},:,:)$ for each fixed scale $s_{j}$, representing all scales using 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$.
 Algorithm 1 Encoding multi-scale structure in the optimized embeddings
-5.2 Optimization
+3.2 Step 2: Optimization design using multi-scale network structure.
 We propose to use optimization based on SGD that begins with the initial embedding within the spectral graph domain.
@@
 This optimization within the SGW domain provides more effective regularization, allowing for the direct removal of noise from each spectral band.
-We outline the steps of the proposed approach, presenting several alternatives for optimization through the incorporation of multi-scale representations (refer to method 1 and method 2 in the algorithm description below).
-1. Optimize Embedding Method 1:
-Given the encoded multi-scale representation $\psi_{\mathbf{x}}\in\mathbb{R}^{KD}\times\mathbb{R}^{N}$ perform optimization in the SGW domain, using the following fuzzy cross entropy loss function:
-$\mathcal{L}(\tilde{\psi}_{\mathbf{x}}|\mathbf{W})=\sum_{i,j}\left(w_{ij}\mbox{log}\frac{w_{ij}}{v_{ij}^{\psi_{\mathbf{x}}}}+(1-w_{ij})\mbox{log}\frac{1-w_{ij}}{1-v_{ij}^{\psi_{\mathbf{x}}}}\right)$ (5)
-where $v_{ij}^{\psi_{\mathbf{x}}}=\frac{1}{1+||\psi_{\mathbf{x}_{i}}-\psi_{\mathbf{x}_{j}}||^{2}}$.
-Dropping terms that do not depends on $\psi_{\mathbf{x}_{i}}$, the gradient of the loss is approximated by:
-$\frac{\partial\mathcal{L}(\tilde{\psi}_{\mathbf{x}}|\mathbf{W})}{\partial\psi_{\mathbf{x}_{i}}}\sim\sum_{j}w_{ij}v_{ij}^{\psi_{\mathbf{x}}}(\psi_{\mathbf{x}_{i}}-\psi_{\mathbf{x}_{j}})-\sum_{j}\frac{1}{||\psi_{\mathbf{x}_{i}}-\psi_{\mathbf{x}_{j}}||^{2}}v_{ij}^{\psi_{\mathbf{x}}}(\psi_{\mathbf{x}_{i}}-\psi_{\mathbf{x}_{j}})$ (6)
-Our optimization process involves using SGD with positive and negative sampling, similar to recent graph embeddings method such as UMAP and LargeVis [26].
-Positive edge samples are associated with attraction (first term on the right hand side), while negative samples (second term on the right hand side) refer to a pair of nodes that are not connected in the graph, which create repulsion among dissimilar points.
-Output: Regularized embedding space $\tilde{\psi}_{\mathbf{x}}$.
-2. 3D-Tensor Based Optimization: Embedding Method 2:
-In method 2, the encoded manifold representation given is a 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$.
-A SGD update rule at iteration $t$ applied to the 3D tensor is:
-$(\tilde{\psi}_{\mathbf{x}}^{(t+1)})_{i,j,k}=(\tilde{\psi}_{\mathbf{x}}^{(t)})_{i,j,k}-\alpha\frac{\partial\mathcal{L}}{(\partial\psi_{\mathbf{x}})_{i,j,k}}$ (7)
-where $\alpha$ is the learning parameter. In our case we employ the cross entropy loss function:
-$\mathcal{L}(\tilde{\psi}_{\mathbf{x}}|\mathbf{W})=\sum_{i,j,k}\left(w_{ij}\mbox{log}\frac{w_{ij}}{v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}}+(1-w_{ij})\mbox{log}\frac{1-w_{ij}}{1-v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}}\right)$ (8)
-We apply the optimization based on SGD with respect to each scale $s_{k}$:
-$\mathcal{L}(\tilde{\psi}_{\mathbf{x}}(s_{k},:,:)|\mathbf{W})=\sum_{i,j}\left(w_{ij}\mbox{log}\frac{w_{ij}}{v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}}+(1-w_{ij})\mbox{log}\frac{1-w_{ij}}{1-v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}}\right)$ (9)
-which has the gradient of the loss is approximated by:
-$\frac{\partial\mathcal{L}(\tilde{\psi}_{\mathbf{x}}(s_{k},:,:)|\mathbf{W})}{\partial\psi_{\mathbf{x}_{i}}(s_{k},:,:)}=\sum_{j}w_{ij}v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}(\psi_{\mathbf{x}_{i}}(s_{k},:,:)-\psi_{\mathbf{x}_{j}}(s_{k},:,:))$
-$-\sum_{j}\frac{1}{||\psi_{\mathbf{x}_{i}}(s_{k},:,:)-\psi_{\mathbf{x}_{j}}(s_{k},:,:)||^{2}}v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}(\psi_{\mathbf{x}_{i}}(s_{k},:,:)-\psi_{\mathbf{x}_{j}}(s_{k},:,:))$
+We outline the steps of the proposed approach.
+3D-Tensor Based Optimization for graph Embedding
+The encoded manifold representation is a 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$. An SGD update rule at iteration $t$, applied to the 3D tensor, is:
+$(\tilde{\psi}_{\mathbf{x}}^{(t+1)})_{i,j,k}=(\tilde{\psi}_{\mathbf{x}}^{(t)})_{i,j,k}-\alpha\frac{\partial\mathcal{L}}{\partial(\psi_{\mathbf{x}})_{i,j,k}}$ (2)
+where $\alpha$ is the learning parameter. In this case, we employ the cross-entropy loss function:
+$\mathcal{L}(\tilde{\psi}_{\mathbf{x}}|\mathbf{W})=\sum_{i,j,k}$ $\left(w_{ij}\log\frac{w_{ij}}{v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}}\right.$
+$\left.+(1-w_{ij})\log\frac{1-w_{ij}}{1-v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}}\right)$ (3)
+where $v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}=\frac{1}{1+||\psi_{\mathbf{x}_{i}}(s_{k},:,:)-\psi_{\mathbf{x}_{j}}(s_{k},:,:)||^{2}}$.
+We apply the optimization based on SGD with respect to each scale $s_{k}$, where the gradient of the loss is approximated by:
+$\frac{\partial\mathcal{L}\left(\tilde{\psi}_{\mathbf{x}}(s_{k},:,:)\mid\mathbf{W}\right)}{\partial\psi_{\mathbf{x}_{i}}(s_{k},:,:)}$ $=\sum_{j}w_{ij}v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}\,r_{ij}(s_{k})$ (4)
+$\quad-\sum_{j}\frac{1}{\|r_{ij}(s_{k})\|^{2}}v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}\,r_{ij}(s_{k})$ (5)
+where $r_{ij}(s_{k})=\psi_{\mathbf{x}_{i}}(s_{k},:,:)-\psi_{\mathbf{x}_{j}}(s_{k},:,:)$.
 The optimization process uses SGD with positive and negative sampling for all scales $s_{k}$ given a pair of positive or negative pair of nodes $i,j$.
@@
 $\tilde{\psi}_{\mathbf{x}}=\sum_{k}\tilde{\psi}_{\mathbf{x}}(s_{k},:,:)$ (6)
-In the 3D tensor based optimization method 2, the dimensionality of the final embedding $\tilde{\psi}_{\mathbf{x}}\in\mathbb{R}^{D}\times\mathbb{R}^{N}$, maintaining the same dimensionality as the original input, and thus there exists a one-to-one correspondence between the coordinates of the input features $\mathbf{x}$ and $\tilde{\psi}_{\mathbf{x}}$.
+Note that the dimensionality of the final embedding
+$\tilde{\psi}_{\mathbf{x}}\in\mathbb{R}^{D}\times\mathbb{R}^{N}$, maintaining the same dimensionality as the original input, and thus there exists a one-to-one correspondence between the coordinates of the input features $\mathbf{x}$ and $\tilde{\psi}_{\mathbf{x}}$.
 This correspondence ensures that each dimension of the feature space is directly mapped to a single dimension in the embedding space $\tilde{\psi}_{\mathbf{x}}$.
-Sampling approaches from network features: We propose a strategic sampling of edges, which further refines the optimization by assessing the significance of edges with respect to the topological structure of the graph.
-Specifically, we construct probability distribution over $V$ (or $\mathbf{W}$), providing the significance of each node/edge, and sample from this distribution.
-We also propose a novel method to implement the sampling strategy, rather than selecting edges randomly as in [21, 26].
-Refer to Appendix for more details.
+Our optimization process involves using SGD with positive and negative sampling, similar to recent graph embeddings methods [33, 25].
+Positive edge samples are associated with attraction (first term on the right hand side), while negative samples (second term on the right hand side) refer to a pair of nodes that are not connected in the graph, which create repulsion among dissimilar points.
+The optimization can be performed jointly for all scales using the 3D tensor.
+Specifically, our implementation employs an optimization process that operates jointly across all scales, as described in the SGD update rule for iteration $t$ in Eq.
+(2).
+At each iteration $t$, associated with either a positive sample (an edge) or a negative sample (a pair of unconnected nodes), the update rule from Eq.(2).
+is applied.
+The derivative of the loss function, as detailed in Eq.
+(4), is computed for all scales $(s_{1},...s_{K})$, with the sign in Eq.(2) depending on whether the sample is positive or negative.
+Once the optimization concludes, the final embedding is derived from the 3D tensor by summing over the scales, as specified in Eq.(6).
+Encoding the representation in a 3D tensor offers two key advantages.
+First, is the computational efficiency: the update rule at each iteration, associated with a pair of nodes, is performed jointly and simultaneously for all scales.
+This avoids the need for independent optimization per scale, which would significantly increase computational complexity.
+The second is optimization alignment: jointly optimizing across scales ensures alignment in the optimization process for all scales with respect to a fixed pair of nodes (positive or negative sample).
+This approach is more effective than optimizing scales independently, as it maintains consistency and improves overall performance.
 4 Theoretical results: Sampling set for Smooth Manifolds with functions defined over Paley-Wiener Spaces
@@
 Pesenson introduced the Paley-Wiener spaces and demonstrated that Paley-Wiener functions of low type are uniquely determined by their values on certain subsets, known as uniqueness sets $U$.
-We show that the SGW operator can represent functions $f$ within the Paley-Wiener space more efficiently than the graph Laplacian operator $\mathcal{L}$.
+We show that the SGW operator can represent functions $f$ within the Paley-Wiener space more efficiently than the graph Laplacian operator $\mathbf{L}$.
 This efficiency is demonstrated by showing that the SGW operator is more effective in representing functions with larger bandwidth $\omega$ in the Paley-Wiener spaces (i.e., with higher frequencies) using subsets of nodes from the uniqueness sets $U$.
@@
 The space $L_{2}(G)$ is defined as the Hilbert space of all complex-valued functions, and $L_{2}(S)$ is defined as the space of all functions from $L_{2}(G)$ with support in $S$: $L_{2}(S)=\left\{\varphi\in L_{2}(G)\mid\varphi(v)=0,v\in V(G)\setminus S\right\}$.
-The $\Lambda$-set is defined as follows: a set of vertices $S\subset V$ is a $\Lambda(S)$-set if any $\varphi\in L_{2}(S)$ satisfies the inequality $||\varphi||\leq\Lambda||\mathcal{L}\varphi||$ for some constant $\Lambda(S)>0$.
+The $\Lambda$-set is defined as follows: a set of vertices $S\subset V$ is a $\Lambda(S)$-set if any $\varphi\in L_{2}(S)$ satisfies the inequality $||\varphi||\leq\Lambda||\mathbf{L}\varphi||$ for some constant $\Lambda(S)>0$.
 The infimum of all $\Lambda>0$ for which $S$ is a $\Lambda$-set is denoted by $\Lambda$.
@@
 The proof is provided in the Appendix.
+Figure 1: Illustration of the proposed method.
+(Left) Initial encoding using SGW in (b) and regularized embedding in (c).
+(Right) Feature embedding space representation, which has a one-to-one correspondence with the input features, can be used along with clustering for interpretability models, such as feature importance.
 5 Experimental Results
-We demonstrates the advantage of our approach over existing approaches using both synthetic and real datasets.
-We evaluate our method by testing its clustering performance using the output embedding and comparing it to several representative methods, including UMAP, t-SNE, Diffusion Maps, ISOMAP, and HeatGeo [19].
-For HeatGeo, we hyperparameter tune the performance on each dataset and report the best one in Section 11.
-For each dataset, we briefly describe their properties and leave fuller details in Section 9.
+We demonstrate the advantage of our approach over existing approaches using both synthetic and real datasets.
+We evaluate our method by testing its clustering performance using the output embedding and comparing it to several representative methods, including UMAP, t-SNE, Diffusion Maps, ISOMAP, PHATE [26], PaCMAP [36], TriMap [1] and HeatGeo [22].
+For UMAP, t-SNE, PHATE, PaCMAP, TriMap and HeatGeo, we hyperparameter tune the performance on each dataset, running the method five times for each hyperparameter configuration and taking the average.
+Then we report the best average score.
+Additional details are provided in the supplementary material.
 Evaluation metrics: We test performance using the Adjusted Rand Index (ARI) and Adjusted Mutual Information (AMI). In all experiments we used $k$-means to cluster the data in the embedding space.
-7.1 Synthetic dataset
-We assess the robustness of our method using the two moons dataset, which is a 2D manifold depicting two interleaving half-circles.
-We sampled $N=600$ points, and set the Gaussian noise’s standard deviation to $0.12$.
-While spectral based methods such as UMAP are effective under relatively “modest” noise levels, their performance deteriorates significantly in the presence of larger amounts of noise (See Figure 3).
-As shown in Table 1, our approach shows robustness to noise and correctly clusters most points in two moons manifold, despite the large amount of noise, and outperforms competing methods.
-We do further experiments on more synthetic datasets in Section 10.
-Method / Accuracy ARI AMI
-UMAP 0.54 0.51
-t-SNE 0.42 0.35
-ISOMAP 0.36 0.3
-Diffusion Maps 0.25 0.19
-HeatGeo 0.54 0.52
-MS-IMAP Method 1 0.75 0.73
-MS-IMAP Method 2 0.89 0.87
-Table 1: Comparison of clustering performance on the Two Moons datasets.
 5.1 Real Datasets
 We study the performance of MS-IMAP compared to other methods for real datasets.
-We chose a mix of financial, biological, and image datasets: the Census dataset [12] is a financial dataset containing information about individuals extracted from the 1994 US Census; the Zilionis dataset [32] is a biological dataset containing single-cell sequencing data from different types of cells, and the Animals with Attributes (AWA) [29] dataset is an image dataset containing images of animals.
-More information about each dataset can be found in Section 9.
-As shown in Table 2, MS-IMAP Method 2 obtains the best performance across all datasets, with a tie between UMAP and MS-IMAP Method 2 in the AWA dataset.
-In the Census dataset, MS-IMAP Method 2 achieves a $47\%$ increase in ARI over HeatGeo, the next best method excluding our own.
-In the Zilionis dataset, MS-IMAP 2 achieves a $9\%$ increase in ARI over t-SNE (we found issues running the official HeatGeo code [19] on Zilionis, so were unable to obtain results for this dataset).
-And in AWA, we have the ARI of MS-IMAP Method 2 is slightly better than UMAP.
-Dataset Census Zilionis AWA
-Method / Accuracy ARI AMI ARI AMI ARI AMI
-UMAP 0.05 0.08 0.45 0.71 0.73 0.80
-t-SNE 0.01 0.02 0.64 0.34 0.68 0.76
-HeatGeo 0.15 0.10 N/A N/A 0.65 0.74
-MS-IMAP Method 1 0.15 0.14 0.69 0.76 0.72 0.80
-MS-IMAP Method 2 0.22 0.15 0.70 0.76 0.74 0.81
-Table 2: Clustering results comparison using ARI and AMI the Census, Zilionis, and AWA datasets. Note HeatGeo had code execution issues on the Zilionis dataset.
-8 Explicit Feature Correspondence with Laplacian Scoring
-The correspondence established between the features and the embedding space becomes valuable in obtaining feature importance.
-One simple measure for feature importance is the Laplacian score (LS) [16], which evaluates each feature based on its correlation with the graph Laplacian eigensystem.
-Specifically, the importance of a feature $f_{r}\in\mathbb{R}^{N}$ is determined from the importance of its corresponding coordinate $l$ in the embedding space, $\tilde{\psi}_{\mathbf{x}_{l}}$.
-We calculate the Laplacian score [16] with respect to the embedding features $\tilde{\psi}_{\mathbf{x}}$ using its graph Laplacian $\mathbf{L}$ and its associated degree matrix $\mathbf{D}$.
-This is achieved by first removing the mean and then computing the Laplacian score using:
-$L_{s}(\tilde{\psi}_{\mathbf{x}})_{l}=\frac{(\tilde{\psi}_{\mathbf{x}})_{l}^{T}\mathbf{L}(\tilde{\psi}_{\mathbf{x}})_{l}}{(\tilde{\psi}_{\mathbf{x}})_{l}^{T}\mathbf{D}(\tilde{\psi}_{\mathbf{x}})_{l}}$ (11)
-Smaller scores imply that the feature $(\tilde{\psi}_{\mathbf{x}})_{l}$ has a larger projection on the subspace of the eigenvectors associated with the smallest eigenvalues, indicating higher importance with respect to the graph structure.
-Each coordinate $l$ in the embedding space $\tilde{\psi}_{\mathbf{x}}$, namely $(\tilde{\psi}_{\mathbf{x}})_{l}$, was constructed using a single coordinate of the original feature $f_{r}$ space.
-Hence, the feature importance of $(\tilde{\psi}_{\mathbf{x}})_{l}$ can be explicitly interpreted as the importance of the associated original feature.
-8.1 Application to Feature importance
-Figure 2: AWA Dataset Laplacian Score plotline showing in order from high to low, the importance of the original features with respect to the embeddings.
-We demonstrate the effectiveness of our approach using a computer vision example where providing interpretation, including feature importance, is of significant interest.
-Here we show in Figure (2) the estimated Laplacian score of the 85 semantic attributes for the AWA datasets sorted from high to low.
-Because the Laplacian score is a function of explained variance, one can argue that this score is a measure of feature information and we observe that the number of features capturing the highest scores is small (on the order of 10).
-This result highlights that most of the information in the data-set which correlates with the embedding is captured by a small group of variables which suggests their relatively higher importance.
-The Laplacian score ranking allows us to select a small number of variables for further analysis in order to better understand the output of statistical studies performed using the embedded data.
-8.2 Runtime and Computational Complexity
-The execution time of our method with Python code implementation with 32 cores Intel Xeon 8259CL running at 2.50Ghz and 256GB of RAM on the Cancer QC data-set of 48,969 samples and 306 features took 12.44 mins using Method 1.
-The computational complexity of Multi-Scale IMAP is of $O(NDlog(N))$.
+We chose a mix of datasets from varied fields: the Census dataset [14] is a financial dataset containing information about individuals extracted from the 1994 US Census; the Zilionis dataset [40] is a biological dataset containing single-cell sequencing data from different types of cells; the Animals with Attributes (AWA) [37] dataset is an image dataset containing images of animals; The Human Activity Recognition (HAR) [2] dataset which consists of sensor data collected from subjects performing six different activities: walking, walking upstairs, walking downstairs, sitting, standing, and lying.
+Each activity is represented as a feature vector of 561 attributes and 7352 samples; and the Cora dataset [30, 24] is a graph dataset of a citation network consisting of scientific publications.
+Note that we only use the features without the natural graph associated with Cora.
+More information about each dataset can be found in the supplementary material.
+As shown in Table 1, MS-IMAP is competitive and performs the best overall, while also providing an interpretable embedding through its one-to-one correspondence with the original features—unlike the other competing graph embedding methods.
+Runtime and Computational Complexity The computational complexity of Multi-Scale IMAP is of $O(NDlog(N))$.
+The execution time of our method with Python code implementation with 32 cores Intel Xeon 8259CL running at 2.50Ghz and 256GB of RAM on the Cancer QC data-set of 48,969 samples and 306 features took 27 mins, on Cora 5 min, on the AWA dataset 1.35 mins and on Census 7 min.
+The increase in running times compared to methods such as UMAP is primarily due to our extension of manifold learning to include interpretability.
+Our approach generally requires the embedding dimensionality to be at least equal to the input dimensionality of the features, which adds computational complexity but ensures meaningful feature representation and interpretation.
+Our method is practical for handling datasets with millions of points and dozens to hundreds of features within a few hours.
+In fields such as finance and genomics/biology, where typical dataset sizes range from under 100,000 to several million data points, analysis is often conducted offline once the graph embedding is computed.
+Runtimes of up to a few hours are generally acceptable and do not present a significant challenge, making the method well-suited for real-world applications.
 Additional details are provided in the Appendix.
-8.3 Conclusions and Limitations
-Identifying the key drivers of high-dimensional datasets is a necessary ingredient to take advantage of unlabelled data in many practical applications in fields such as finance and biology.
+Data HAR Census Zilionis AWA Cora
+Method / Accuracy ARI AMI ARI AMI ARI AMI ARI AMI ARI AMI
+Features 0.45 0.58 0.11 0.15 0.48 0.63 0.07 0.16 0.01 0.07
+ISOMAP 0.55 0.65 0.18 0.09 0.44 0.56 0.52 0.60 0.12 0.17
+Diffusion Maps N/A N/A 0.06 0.16 0.01 0.02 0.22 0.42 0.03 0.07
+UMAP 0.61 0.73 0.23 0.15 0.52 0.74 0.74 0.82 0.33 0.36
+t-SNE 0.71 0.77 0.17 0.17 0.41 0.70 0.73 0.80 0.21 0.27
+PHATE 0.55 0.66 0.20 0.09 0.65 0.72 0.66 0.73 0.12 0.22
+TriMap 0.59 0.72 0.2 0.14 0.62 0.77 0.74 0.81 0.15 0.2
+PaCMAP 0.62 0.75 0.24 0.15 0.55 0.76 0.75 0.81 0.21 0.26
+HeatGeo 0.6 0.69 0.15 0.10 N/A N/A 0.65 0.74 0.17 0.20
+MS-IMAP 0.67 0.8 0.23 0.15 0.70 0.76 0.74 0.81 0.33 0.36
+Table 1: Clustering results comparison using ARI and AMI on the Census, Zilionis, AWA, and Cora datasets. Note HeatGeo had code execution issues on the Zilionis dataset.
+6 Feature Importance using Explicit Correspondence with Embedding Features
+The correspondence established between the features and the embedding space offers a valuable tool for assessing feature importance.
+We explore two alternative methods for measuring feature importance, leveraging the constructed embedding space, where each dimension can be traced back to its corresponding feature in the original space.
+Specifically, we test these methods in an unsupervised or self-supervised setting where the subset of features are selected without using label information.
+6.1 Assessing Embedding Feature Importance using the Laplacian Score
+The first method utilizes the Laplacian Score (LS) [19], which assesses the importance of each feature by examining its correlation with the eigenvalues of the graph Laplacian.
+The significance of the variable corresponding to the $l$-th coordinate, represented by the feature $f_{l}\in\mathbb{R}^{N}$, is inferred from the relevance of its corresponding coordinate $l$ in the embedding space, $\tilde{\psi}_{\mathbf{x}_{l}}$.
+The Laplacian Score is calculated with respect to the embedding features $\tilde{\psi}_{\mathbf{x}}$ using the graph Laplacian $\mathbf{L}$ and the degree matrix $\mathbf{D}$.
+To compute the Laplacian Score, we first subtract the mean and then calculate it as follows: $L_{s}(\tilde{\psi}_{\mathbf{x}})_{l}=\frac{(\tilde{\psi}_{\mathbf{x}})_{l}^{T}\mathbf{L}(\tilde{\psi}_{\mathbf{x}})_{l}}{(\tilde{\psi}_{\mathbf{x}})_{l}^{T}\mathbf{D}(\tilde{\psi}_{\mathbf{x}})_{l}}$
+Smaller scores indicate that the embedding feature $(\tilde{\psi}_{\mathbf{x}})_{l}$ has a greater projection onto the subspace of eigenvectors corresponding to the smallest eigenvalues, signifying higher importance concerning the global graph structure.
+Therefore, the feature importance of $(\tilde{\psi}_{\mathbf{x}})_{l}$ can be directly interpreted as the importance of the corresponding original feature.
+6.2 Assessing Importance of Embedding Features using Mutual Information
+In the second method, we propose a self-supervised approach that uses Mutual Information (MI) to estimate the importance of each dimension in the embedding space, which can then be explicitly linked to the original features.
+First, we apply $k$-means to cluster the MS-IMAP embedding space and use the resulting clusters as pseudo-labels.
+We then compute MI to quantify the shared information between features and these pseudo-labels, similar to how MI would be computed with ground truth labels.
+This step allows MI to measure how much each feature contributes to the cluster assignments, highlighting those most informative in distinguishing clusters.
+The MI-Based Feature Importance identify features that are informative about cluster assignments.
+Thus, masking these features would especially degrade the clustering performance.
+We summarize the MI-based unsupervised feature selection in Algorithm 2.
+In comparison to the MI based feature importance method, using the Laplacian score [19] we have that the embedding may have combined original features in a way that makes individual Laplacian Scores less indicative of feature importance.
+The Laplacian score prioritize neighborhood preservation, which might not directly translate to cluster discrimination efficacy.
+On the other hand, while the proposed MI based feature importance identifies features that are directly informative for distinguishing clusters, one of its possible limitations may be relaying on the quality of pseudo-labels, which may not align with true clusters.
+  Input: MS-IMAP embeddings $\psi_{\mathbf{x}}$, number of clusters $k$
+1  Output: MI scores for each feature
+  Step 1: Apply k-means clustering on the embedding space $\psi_{\mathbf{x}}$ to obtain cluster labels
+[…]
+  Step 2: Use the resulting cluster labels as pseudo-labels
+[…]
+  Step 3: Initialize an empty list to store mutual information (MI) values for each feature
+[…]
+  Step 4: For each feature $i$ in the original data or embedding space:
+            Step 4.1: Compute mutual information between feature $\psi_{\mathbf{x}}(i)$ and the pseudo-labels
+4            Step 4.2: $MI\_score=\text{MI}(\psi_{\mathbf{x}}(i),c({\psi_{\mathbf{x}}})$)
+5            Step 4.3: Append $MI\_score$ to mi_scores
+  Step 5: Output the list of MI scores
+            Return mi_scores
+Algorithm 2 MI-based Unsupervised Feature Selection
+Figure 2: Clustering results on Cora using a subset of features based on feature importance, comparing MS-IMAP embedding space and the input features.
+6.3 Assessing Feature Importance via Embedding Correspondence
+We demonstrate the application of MS-IMAP for feature importance by evaluating clustering performance using $k$-means as a function of the number of features, based on the most significant features identified by two methods: the Laplacian score and Mutual Information-based feature selection.
+We compared the clustering results (using ARI and AMI) when $k$-means is applied to different subsets of embedding features selected by the Laplacian score and MI-based methods, with results obtained by applying the same feature selection methods directly to the original feature space.
+Unlike most manifold learning and graph embedding methods, MS-IMAP produces embeddings where the dimensions can be directly mapped to the original feature space, enabling the assessment of individual feature contributions.
+On the Cora dataset, Figure 2 shows that for both feature selection methods, the clustering accuracy obtained using the MS-IMAP embedding is significantly higher than when using the raw features across different subsets of features.
+This improvement is due to the robustness of the MS-IMAP embedding, while still offering interpretability by linking the embedding dimensions to the original features.
+7 Discussion
+Identifying the key factors of high-dimensional datasets is a necessary ingredient to take advantage of unlabelled data in many practical applications.
 In this article, we introduce a novel contrastive learning framework for manifold learning via graph embeddings, capitalizing on both low and high-frequency information.
@@
 We then use an SGD-based optimization scheme together with innovative strategies such as 3D Tensor encoding to derive the embeddings.
-We study the theoretical properties of the spectral graph wavelet (SGW) representation by considering functions in Paley-Wiener spaces on combinatorial graphs and prove that the SGW operator provides a more effective representation using the concept of the $\Lambda$set.
-Finally, we show that the embeddings are interpretable using a simple derivation of feature importance of the embedding and original feature spaces.
-The construction of our methodology for generating the embeddings implicitly gives us a way to tie the original and transformed feature spaces together which is lacking in current non-linear manifold learning techniques to the best of our knowledge.
-Limitations: While our work has demonstrated several significant properties with strong performance on challenging datasets, it is not without limitations.
+To motivate the use of the spectral graph wavelet (SGW) for representation, we examine the theoretical properties of the SGW by studying functions in Paley-Wiener spaces on combinatorial graphs.
+We prove that the SGW operator provides a more effective representation through the concept of the $\Lambda$-set.
+Finally, we demonstrate that the embeddings are interpretable using two alternative methods for unsupervised and self-supervised feature importance, applied to both the embedding and original feature spaces.
+The construction of our methodology inherently links the original and transformed feature spaces, a capability that is notably absent in most existing non-linear manifold learning techniques, to the best of our knowledge.
+Limitations and future work include the following.
 Firstly, we rely on the assumption that the features, and thus the similarity measure, adequately capture information to reveal geodesic distances between manifold points.
@@
 Appendix
+8 Encoding Method 1
+[…]
+$\mathbf{c}(\psi_{{f}_{i}}(s_{k},:),\psi_{{f}_{j}}(s_{k},:))=\psi_{{f}_{i}}(s_{k},:)\left|\right|\psi_{{f}_{j}}(s_{k},:)$
+denoting the concatenation of the vectors corresponding to the multi-scale representation $\psi_{{f}_{i}}(s_{k},:)$ and $\psi_{{f}_{j}}(s_{k},:)$.
+For method 1 all features and all scales are concatenated together, the resulting matrix $\psi_{\mathbf{x}}$ can be represented as
+$\psi_{\mathbf{x}}=\psi_{{f}_{1}}(s_{1},:)\|\psi_{{f}_{2}}(s_{1},:)\|....\|\psi_{{f}_{D-1}}(s_{K},:)\|\psi_{{f}_{D}}(s_{K},:)$
+where $\psi_{\mathbf{x}}\in\mathbb{R}^{KD}\times\mathbb{R}^{N}$.
+1. Optimize Embedding Method 1:
+Given the encoded multi-scale representation $\psi_{\mathbf{x}}\in\mathbb{R}^{KD}\times\mathbb{R}^{N}$ perform optimization in the SGW domain, using the following fuzzy cross entropy loss function:
+$\mathcal{L}(\tilde{\psi}_{\mathbf{x}}|\mathbf{W})=\sum_{i,j}\left(w_{ij}\mbox{log}\frac{w_{ij}}{v_{ij}^{\psi_{\mathbf{x}}}}+(1-w_{ij})\mbox{log}\frac{1-w_{ij}}{1-v_{ij}^{\psi_{\mathbf{x}}}}\right)$ (7)
+where $v_{ij}^{\psi_{\mathbf{x}}}=\frac{1}{1+||\psi_{\mathbf{x}_{i}}-\psi_{\mathbf{x}_{j}}||^{2}}$.
+Dropping terms that do not depends on $\psi_{\mathbf{x}_{i}}$, the gradient of the loss is approximated by:
+$\frac{\partial\mathcal{L}(\tilde{\psi}_{\mathbf{x}}|\mathbf{W})}{\partial\psi_{\mathbf{x}_{i}}}\sim\sum_{j}w_{ij}v_{ij}^{\psi_{\mathbf{x}}}(r_{ij})-\sum_{j}\frac{1}{||r_{ij}||^{2}}v_{ij}^{\psi_{\mathbf{x}}}(\psi_{\mathbf{x}_{i}}-\psi_{\mathbf{x}_{j}})$ (8)
+where $r_{ij}=\psi_{\mathbf{x}_{i}}-\psi_{\mathbf{x}_{j}}$, and $\psi_{\mathbf{x}_{i}}$ corresponds to the vector associated with node $i$ multi-scale representation concatenated for all the graph signals and all spectral bands.
+since for method 1 all features and all scales are concatenated together,
+Figure 3: The encoding step of our framework is illustrated as a mapping between each coordinate of the input features and a corresponding dimension in the embedding space.
+This mapping is facilitated through our proposed approach, which aligns the input dimensions of the features with the latent embedding space.
+By establishing this correspondence, we acquire the capability to interpret each dimension in the embedding space.
+Such interpretation allows for various analyses, including providing feature importance relative to the latent space.
+9 Additional experiments: Embedding Feature Importance using Mutual Information and Laplacian Score
+We include additional experiments that demonstrate the application of MS-IMAP for feature importance by evaluating clustering performance using $k$-means as a function of the number of features (in addtion to the one found in the Section 6), based on the most significant features identified by the Laplacian score and Mutual Information-based feature selection.
+We compared the clustering results (using ARI and AMI) when $k$-means is applied to different subsets of embedding features selected by the Laplacian score and MI-based methods, with results obtained by applying the same feature selection methods directly to the original feature space.
+On the Zillions dataset, Figures 5 and 6 shows that for both feature selection methods, the clustering accuracy obtained using the MS-IMAP embedding is significantly higher than when using the raw features across different subsets of features.
+This improvement is due to the robustness of the MS-IMAP embedding, while still offering interpretability by linking the embedding dimensions to the original features.
+We also show in Figure 4 the clustering accuracy obtained using the MS-IMAP copared to using the raw features using AMI, demonstrating similar observations regarding the advantages of MS-IMAP.
+Figure 4: Clustering results (AMI) on Cora using a subset of features based on feature importance, comparing MS-IMAP embedding space and the input features.
+Figure 5: Clustering results (ARI) on Zillions using a subset of features based on feature importance, comparing MS-IMAP embedding space and the input features.
+Figure 6: Clustering results (AMI) on Zillions using a subset of features based on feature importance, comparing MS-IMAP embedding space and the input features.
 10 Dataset Details
@@
 More specifically, to produce these points we use Sci-kit Learn’s sklearn.datasets.make_moons function with $n\_samples=600$ and $noise=0.12$.
-In Figure 3, we show example clusterings for each of the methods mentioned in Section 7.
+In Figure 7, we show example clusterings for each of the methods mentioned in the Experimental Results section of the main paper.
 Census.
@@
 More information can be found in Section 4.1 of [37].
+Cora.
+The Cora dataset has 2,708 scientific publications categorized into seven classes.
+The network has 5,429 links and each publication is described as a binary word vector indicating the presence or absence of a word.
+The dictionary consists of 1,433 words.
+In our experiments, we do not use the given graph, and only use the features.
+[30, 24].
+HAR.
+The Human Activity Recognition (HAR) dataset consists of sensor data collected from 30 subjects performing six different activities: walking, walking upstairs, walking downstairs, sitting, standing, and lying.
+The data was recorded using accelerometer and gyroscope sensors embedded in a smartphone worn on the waist.
+Each activity is represented as a feature vector of 561 attributes and 7352 samples, extracted from the raw sensor signals through signal processing techniques.
+In our experiments, we do not use the temporal dependencies and only utilize the extracted features.
 Figure 7: Clustering results using two moons
-10 Experimental Results on More Datasets
-Here we show further experiments on more datasets.
-Dense cluster inside a sparse circle: In this example, we sampled $500$ points from a dense cluster (using a uniform distribution) situated in the interior of a sparse cluster (using 100 points), maintaining a ratio of 5:1.
-UMAP relies on a well-initialized kNN graph, however the disparate densities in the two manifolds lead to erroneous edge weights, resulting in errors in segmentation.
-Despite utilizing the same initial graph construction as UMAP, our method, employing a multi-scale representation, is more successful in segmenting the two manifolds accurately in the presence of differing densities.
-HeatGeo, a competitive state-of-the-art method, does the best here, and is able to separate the dense-sparse circles from each other.
+11 Experimental Results on two moons Synthetic Dataset
+11.1 Synthetic dataset
+Here we show further experiments on more data, where we assess the robustness of our method using the two moons dataset, which is a 2D manifold depicting two interleaving half-circles.
+We sampled $N=600$ points, and set the Gaussian noise standard deviation to $0.12$.
+While spectral based methods such as UMAP are effective under relatively “modest” noise levels, their performance deteriorates in the presence of larger amounts of noise.
+As shown in Table 1 under the Moons column, our approach is robust and correctly clusters most points despite the large noise level, and competitive with the competing methods.
 Data Dense-sparse clusters
 Method/Accuracy ARI AMI
-UMAP 0.75 0.72
-t-SNE 0.30 0.33
-ISOMAP 0.74 0.65
-Diffusion Maps 0.81 0.76
-HeatGeo 1.0 1.0
-Ms-IMAP Method 1 0.82 0.8
-Ms-IMAP Tensor-encoding Method 2 0.77 0.73
-Table 3: Comparison of clustering performance on Two Moons and Dense-sparse clusters datasets.
-11 HeatGeo Hyperparameter Tuning Details
-We tune HeatGeo on each dataset in Section 7, by tuning over the parameter space:
+Features 0.24 0.18
+UMAP 0.54 0.51
+t-SNE 0.42 0.35
+ISOMAP 0.36 0.3
+Diffusion Maps 0.25 0.19
+PHATE 0.42 0.35
+HeatGeo 0.54 0.52
+MS-IMAP 0.89 0.87
+Table 2: Comparison of clustering performance on the two moons dataset.
+12 Hyperparameter tuning details
+We tune UMAP on each dataset in the Experimental Results section, by tuning over the parameter space in Table 3.
+Hyperparameter Set of values
+n_neighbors 2, 10, 15, 20, 30, 50, 100
+min_dist 0, 0.1, 0.5, 0.99
+n_components 2, 3
+Table 3: Space of parameters in which we tuned UMAP.
+12.1 TSNE Hyperparameter Tuning Details
+We tune TSNE on each dataset in the Experimental Results section, by tuning over the parameter space in Table 4.
+Hyperparameter Set of values
+perplexity 15, 30, 60
+early_exaggeration 12, 24
+n_components 2, 3
+Table 4: Space of parameters in which we tuned TSNE.
+12.2 PHATE Hyperparameter Tuning Details
+We tune PHATE on each dataset in the Experimental Results section, by tuning over the parameter space in Table 5.
+Hyperparameter Set of values
+knn 3, 5, 10
+decay (i.e. $\alpha$) 6, 10, 14, 18, 40
+t auto, 50, 100, 150, 200
+Table 5: Space of parameters in which we tuned PHATE.
+12.3 HeatGeo Hyperparameter Tuning Details
+We tune HeatGeo on each dataset in the Experimental Results section, by tuning over the parameter space in Table 6,
 Hyperparameter Set of values
@@
 20 0.22 0.15 0.70 0.77 0.71 0.79
-Table 5: Ablation study on MS-IMAP with method 2, varying the number of neighbors. The number of filters is kept at 5.
+Table 7: Ablation study on MS-IMAP with encoding method 2, varying the number of neighbors. The number of filters is kept at 5.
 We also study the effect of varying the number filters. In table 8, we also see similar results when using different filters, showing the stability of MS-IMAP.
@@
 7 0.22 0.15 0.72 0.78 0.74 0.81
-Table 6: Ablation study on MS-IMAP with method 2, varying the number of filters. The number of neighbors is kept at 15.
+Table 8: Ablation study on MS-IMAP with encoding method 2, varying the number of filters. The number of neighbors is kept at 15.
 14 Ablation Studies of Hyperparameters for t-SNE, Isomaps, and Diffusion Maps
@@
 Experiments were performed on Virtual Server containers with 32 cores Intel Xeon 8259CL running at 2.50Ghz and 256GB of RAM.
-On the AWA dataset of $5,685$ samples and $2,048$ features, Method 1 and the 3D based Tensor Method 2 both run in 1.35 mins.
+On the AWA dataset of $5,685$ samples and $2,048$ features, the 3D based Tensor and the simple concatenation Method both run in 1.35 mins.
 Finally on the Cancer QC data-set of 48,969 samples and 306 features, Method 1 took 12.44 mins.
@@
 The Paley-Wiener spaces were introduced on combinatorial graphs in [27] and a corresponding sampling theory was developed which resembles the classical one.
-Pesenson proved in [22] that Paley-Wiener functions of low type are uniquely determined by their values on certain subgraphs (which are composed from set of nodes known as the uniqueness sets) and can be reconstructed from such sets in a stable way.
+Pesenson proved in [27] that Paley-Wiener functions of low type are uniquely determined by their values on certain subgraphs (which are composed from a set of nodes known as the uniqueness sets) and can be reconstructed from such sets in a stable way.
 We demonstrate that the SGW operator can represent functions $f$ that reside in the Paley-Wiener space on combinatorial graphs more efficiently than the graph Laplacian operator $\mathcal{L}$.
@@
 In one way, by the ability of the SGW operator to accurately represent functions with larger bandwidth, i.e., $f\in PW_{\omega^{\prime}}(G)$ where $\omega<\omega^{\prime}$.
+We first summarize the main notions and definitions.
+The space $L_{2}(G)$ is the Hilbert space of all complex-valued functions $f:V\rightarrow\mathbb{C}$ with the following inner product $\left\langle f,g\right\rangle=\sum_{v\in G}f(v)\overline{g(v)}$ and the norm
+$||f||=\left(\sum_{v\in V}{|f(v)|^{2}}\right)^{1/2}$ (10)
+The Laplace (normalized) operator $\cal L$ is defined by the formula [27]:
+$\mathcal{L}f(v)=\frac{1}{\sqrt{d(v)}}\sum_{u\sim v}\left(\frac{f(v)}{\sqrt{d(v)}}-\frac{f(u)}{\sqrt{d(u)}}\right),f\in L_{2}(G)$ (11)
 In order to prove these results, we will first need the following definitions:
 Definition 1
-The Paley-Wiener space of $\omega$bandlimited signals is defined as follows:
+The Paley-Wiener space of $\omega$bandlimited signals $f\in L_{2}(G)$ is defined as follows:
 $PW_{\omega}(G)=\left\{f|\hat{f}(\lambda)=0\,\,\forall\,\,\lambda>\omega)\right\}$ (12)
-We summarize the main notions and definitions.
-We consider simple undirected unweighted and connected graph $G=(V,\mathbf{W})$, where $V$ is its set of $N$ vertices and $\mathbf{W}$ is its set of edges.
+We consider a simple, undirected, unweighted, and connected graph $G=(V,\mathbf{W})$, where $V$ is its set of $N$ vertices and $\mathbf{W}$ is its set of edges.
 The degree of $v$ is number of vertices adjacent to a vertex $v$ is and is denoted by $d(v)$.
@@
 Definition 2
-A set of vertices $U\subset V$ is a uniqueness set for a space $PW_{\omega}(G)$ if we have two functions from $PW_{\omega}(G)$ that coincide on $U$ then they coincide on $V$.
-Next the following definition of $L_{2}(S)$ is provided as the set of finite energy signals whose support is contained in $S$.
-The space $L_{2}(G)$ is the Hilbert space of all complex-valued functions $f:V\rightarrow\mathbb{C}$ with the following inner product $\left\langle f,g\right\rangle=\sum_{v\in G}f(v)\overline{g(v)}$ and the norm
-$||f||=\left(\sum_{v\in V}{|f(v)|^{2}}\right)^{1/2}$ (15)
-The Laplace (normalized) operator $\cal L$ is defined by the formula [22]:
-$\mathcal{L}f(v)=\frac{1}{\sqrt{d(v)}}\sum_{u\sim v}\left(\frac{f(v)}{\sqrt{d(v)}}-\frac{f(u)}{\sqrt{d(u)}}\right),f\in L_{2}(G)$ (16)
+A set of vertices $U\subset V$ is a uniqueness set for a space $PW_{\omega}(G)$ if for every two functions from $PW_{\omega}(G)$ that coincide on $U$, then they coincide on $V$.
 Definition 3
@@
 Definition 4
-[22] We say that a set of vertices $S\subset V$ is a $\Lambda$set if for any $\varphi\in L_{2}(S)$ it admits a Poincare inequality with a constant $\Lambda(S)>0$
+[27] We say that a set of vertices $S\subset V$ is a $\Lambda$set if for any $\varphi\in L_{2}(S)$ it admits a Poincare inequality with a constant $\Lambda=\Lambda(S)>0$
 $||\varphi||\leq\Lambda||\mathbf{\cal L}\varphi||,\,\varphi\in L_{2}(S)$ (14)
 The infimum of all $\Lambda(S)>0$ for which $S$ is a $\Lambda$set will be called the Poincare constant of the set $S$ and denoted by $\Lambda$.
-Definition above provides a tool to determine when bandlimited signals in Paley-Wiener spaces $PW_{\omega}(G)$ can be uniquely represented from their samples on a given set.
-The role of $\Lambda$-sets was explained and proved in the following Theorem by Pesenson [22], that shows that if $S\subset V$ then any signal $f\in PW_{\omega}(G)$ can be uniquely represented by its samples in the complement set $U=V(G)\setminus S$:
+The definition above provides a tool to determine when bandlimited signals in Paley-Wiener spaces $PW_{\omega}(G)$ can be uniquely represented from their samples on a given set.
+The role of $\Lambda$-sets was explained and proved in the following Theorem by Pesenson [27], that shows that if $S\subset V$ then any signal $f\in PW_{\omega}(G)$ can be uniquely represented by its function values in the complement set $U=V(G)\setminus S$:
 Theorem 2
 [27] If $S\subset V$ is a $\Lambda$- set, then the set $U=V(G)\setminus S$ is a uniqueness set for any space $PW_{\omega}(G)$ with $\omega<\frac{1}{\Lambda}$.
-Remark: Note that non-trivial uniqueness sets can not exist for functions from any Paley-Wiener subspace $PW_{\omega}(G)$ with any $\lambda_{0}\leq\omega\leq\lambda_{N}$, but can they can exist for some range $\lambda_{0}\leq\omega<\Omega$, as was shown in [22].
+Remark: Note that non-trivial uniqueness sets can not exist for functions from any Paley-Wiener subspace $PW_{\omega}(G)$ with any $\lambda_{0}\leq\omega\leq\lambda_{N}$, but they can exist for some range $\lambda_{0}\leq\omega<\Omega$, as was shown in [27].
 We state one of our main results, in which we employ the SGW operator $\psi$ to characterize the uniqueness set using the $\Lambda_{\psi}$-set, therefore extending the $\Lambda$-set concerning the graph Laplacian operator $\cal L$.
@@
 Let $G=(V,\mathbf{W})$ be a connected graph with $N$ vertices.
-Assume that there exist a set of vertices $S\subset V$ for which the conditions (1)-(2) in Lemma 1 hold true.
+Assume that there exist a set of vertices $S\subset V$ for which the conditions (1)-(2) in Lemma 1 below hold true.
 Let $\psi$ be the SGW operator using a polynomial $p(\mathbf{\cal L})$ with the coefficients $\left\{a_{k}\right\}_{k=0}^{K}$ such that $\psi_{f}=\sum_{k=0}^{K}a_{k}\mathbf{\cal L}^{k}f$.
@@
 Our result Theorem 3 holds similar assumptions as in [27].
+Lemma 1 (Lemma 3.6 of [27])
+Given a connected graph $G=(V,\mathbf{W})$, a set of vertices $S\subset V$, its complement $U=V\setminus S$, for which the following conditions hold true
+  1. 1.
+    For every $s\in S$ there exists $u\in U$ that is a neighbor of $s$, i.e., $w(u,s)>0$.
+  2. 2.
+    For every $s\in S$ there exists at least one neighbor node $u\in U$ whose adjacency set intersects $S$ only over $s$.
+Then there exist a set of vertices $S\subset V$ which is a $\Lambda$-set, with $\Lambda=d(G)$.
+We require one more property of the Laplacian operation, Lemma 3 below. To this end, we use a lemma from [27],
+Lemma 2 (Lemma 3.9 of [27])
+If $S$ is a $\Lambda$-set, then for any $\varphi\in L_{2}(S)$ and all $t\geq 0$, $k=2^{l}$, $l=0,1,2,...$
+$\frac{1}{\Lambda^{k}}\|\mathcal{L}^{t}\varphi\|\leq\|\mathcal{L}^{k+t}\varphi\|,\quad\varphi\in L_{2}(S)$
+Now we are ready to prove a modification of Lemma 3.9 from [27]:
 Lemma 3
-[22] Given a connected graph $G=(V,\mathbf{W})$ graph with $N$ vertices for which the following conditions hold true: (1) For every $s\in S$ there exists $u\in U$ that is a neighbor of $s$, i.e., $w(u,s)>0$.
-(2) for every $s\in S$ there exists at least one neighbor node $u\in U$ whose adjacency set intersects $S$ only over $s$.
-Then there exist a set of vertices $S\subset V$ which is a $\Lambda$-set, with $\Lambda=d(G)$.
+Let $S$ be a $\Lambda$-set, $\varphi\in L_{2}(S)$, and let $m\geq 1$ be an integer. Then
+$\frac{1}{\Lambda^{m}}\|\varphi\|\leq\|\mathcal{L}^{m}\varphi\|.$
+Proof: Because every positive integer has binary representation, then we have for some $j_{i}\in\{0,1,2,...\},i=1,2,...,N$, that
+$m=\sum_{i=1}^{N}2^{j_{i}}.$
+Using Lemma 2, let $k=2^{j_{1}}$ and $t=2^{j_{2}}+\cdots+2^{j_{N}}$, so $m=k+t$. Then we have
+$\|\mathcal{L}^{m}\varphi\|$ $=\|\mathcal{L}^{k+t}\varphi\|$
+$\geq\frac{1}{\Lambda^{k}}\|\mathcal{L}^{t}\varphi\|$
+$=\frac{1}{\Lambda^{2^{j_{1}}}}\|\mathcal{L}^{2^{j_{2}}+\cdots+2^{j_{N}}}\varphi\|$
+We can now let $k_{2}=2^{j_{2}}$ and $t_{2}=2^{j_{3}}+\cdots+2^{j_{N}}$ so that,
+$\|\mathcal{L}^{m}\varphi\|\geq\frac{1}{\Lambda^{2^{j_{1}}}}\|\mathcal{L}^{2^{j_{2}}+\cdots+2^{j_{N}}}\varphi\|$ $=\frac{1}{\Lambda^{2^{j_{1}}}}\|\mathcal{L}^{k_{2}+t_{2}}\varphi\|$
+$\geq\frac{1}{\Lambda^{2^{j_{1}}}}\frac{1}{\Lambda^{k_{2}}}\|\mathcal{L}^{t_{2}}\varphi\|$
+$=\frac{1}{\Lambda^{2^{j_{1}}+2^{j_{2}}}}\|\mathcal{L}^{2^{j_{3}}+\cdots+2^{j_{N}}}\varphi\|$
+Continuing in this way, we eventually get,
+$\|\mathcal{L}^{m}\varphi\|\geq\frac{1}{\Lambda^{m}}\|\varphi\|.$
+completing the proof. $\square$
 In the next theorem, we expand the characterization of the uniqueness set using the $\Lambda$-set concerning the graph Laplacian operator $\mathcal{L}$ to include cases where we employ the SGW operator $\psi$.
@@
 Proof of Theorem 3:
-Assuming $f\in PW_{\omega}(G)$, $f$ can be efficiently represented using a polynomial $p(\mathcal{L})$ instead of a kernel function $g$, where the wavelet at node $i$ and a fixed scale $s$ is calculated using the polynomial $p(\mathbf{\cal L})$ of the Laplacian:
-$\psi(i)=\sum_{k=0}^{K}a_{k}\mathbf{\cal L}^{k}\delta_{i}$ (19)
-and we can write the SGW coefficients with respect to the function $\varphi$ as:
-$\psi_{{\varphi}}=\sum_{l=0}^{N-1}\sum_{k=0}^{K}a_{k}\mathbf{\cal L}^{k}\varphi$
-Following the assumptions of Theorem 1 implies that there exists a subset $U^{*}\subset U$ such that for every $s\in S$ there exists at least one point $u^{*}\in U^{*}$ whose adjacency set intersects $S$ only over $s$ and from this property we have
-$\mathbf{\cal L}\varphi(u^{*})=-\frac{\varphi(s)}{\sqrt{d(s)d(u^{*})}},u^{*}\in U^{*},s\in S$ (20)
-We will use this property to show that the set $S$ is a $\Lambda_{\psi}$-set with respect to the operator $\psi$.
-Using 20 and taking taking powers of the Laplacian operator considering $u^{*}\in U^{*},s\in S$ we have that:
-$\mathbf{\cal L}^{k}\varphi(u^{*})=(-1)^{k}\frac{\varphi(s)}{{(d(s)d(u^{*}))^{k/2}}},u^{*}\in U^{*},s\in S$ (21)
-and
-$|\mathbf{\cal L}^{k}\varphi(u^{*})|\geq\frac{|\varphi(s)|}{\Lambda^{k}},u^{*}\in U^{*},s\in S$ (22)
-Thus we have that $\forall\varphi\in L_{2}(S)$
-$||\psi_{\varphi}||=\left(\sum_{i,j}\sum_{k=0}^{K}|a_{k}({\cal L}^{k}\varphi)_{ij}|^{2}\right)^{1/2}\geq\left(\sum_{s\in S}\sum_{k=0}^{K}|a_{k}({\cal L}^{k}\varphi(u^{*}))|^{2}\right)^{1/2}$ (23)
-using similar argument as in Lemma 1 we have that
-$\left(\sum_{k=0}^{K}\sum_{s\in S}|a_{k}{\cal L}^{k}\varphi(u^{*})|^{2}\right)^{1/2}=\left(\sum_{k=0}^{K}|a_{k}|^{2}\sum_{s\in S}|{\cal L}^{k}\varphi(u^{*})|^{2}\right)^{1/2}\geq||\varphi||\left(\sum_{k=0}^{K}|a_{k}|^{2}\frac{1}{\Lambda^{2k}}\right)^{1/2}$ (24)
-which proves the claim of the Theorem. $\square$
-Remark 1: An important property which can be observed from Theorem 3 is the following: given the SGW $\psi_{{f}}$ as a spectral representation operator, we can choose coefficients $\left\{a_{k}\right\}_{k=0}^{K}$ such that we obtain a $\Lambda_{\psi}$ set associated with the operator $\psi$, which is smaller than the $\Lambda$ set associated with the Laplacian operator $\cal L$.
+Let $S\subset V$ be a set satisfying the conditions in Lemma 1, so $S$ is $\Lambda$-set. Then by Lemma 3, for any integer $k\geq 1$,
+$\frac{1}{\Lambda^{k}}\|\varphi\|\leq\|\mathcal{L}^{k}\varphi\|.$
+The SGW operator of $\varphi$ has the form,
+$\psi_{\varphi}:=\sum_{i,j}\sum_{k=0}^{K}a_{k}\mathbf{\cal L}^{k}\varphi$
+for some $K\geq 1$, $a_{k}\in\mathbb{R}$ are coefficients, and where the summation over $i$ and $j$ represents the summation over the nodes of the graph. Now we calculate,
+$\|\psi_{\varphi}\|$ $=\left(\sum_{i,j}\sum_{k=0}^{K}\left|a_{k}(\mathcal{L}^{k}\varphi)_{ij}\right|^{2}\right)^{1/2}$
+$=\left(\sum_{i,j}\sum_{k=0}^{K}a_{k}^{2}\left|(\mathcal{L}^{k}\varphi)_{ij}\right|^{2}\right)^{1/2}$
+$=\left(\sum_{k=0}^{K}a_{k}^{2}\sum_{i,j}\left|(\mathcal{L}^{k}\varphi)_{ij}\right|^{2}\right)^{1/2}$
+$=\left(\sum_{k=0}^{K}a_{k}^{2}\|\mathcal{L}^{k}\varphi\|^{2}\right)^{1/2}$
+$\geq\left(\sum_{k=0}^{K}a_{k}^{2}\frac{1}{\Lambda^{2k}}\|\varphi\|^{2}\right)^{1/2}$
+$=\|\varphi\|\left(\sum_{k=0}^{K}a_{k}^{2}\frac{1}{\Lambda^{2k}}\right)^{1/2}$
+Letting
+$\Lambda_{\psi}:=\frac{1}{\sqrt{\sum_{k=0}^{K}\frac{a_{k}^{2}}{\Lambda^{2k}}}}$
+Then we have $\|\varphi\|\leq\Lambda_{\psi}\|\psi_{\varphi}\|$, so $S$ is also a $\Lambda_{\psi}$-set.
+$\square$
+From the assumptions and proof of Theorem 3, we get the immediate result:
+Corollary 1
+Let $S\subset V$ be a $\Lambda$-set. Then $S$ is also a $\Lambda_{\psi}$-set.
+Remark 1: An important property which can be observed from Theorem 3 is the following: given the SGW $\psi_{{f}}$ as a spectral representation operator applied on $f$, we can choose coefficients $\left\{a_{k}\right\}_{k=0}^{K}$ such that we obtain a $\Lambda_{\psi}$ set associated with the operator $\psi$, which is smaller than the $\Lambda$ set associated with the Laplacian operator $\cal L$.
 This implies that the operator $\psi$ provides more flexibility and better control over smoothness properties in comparison to the Laplacian operator.
@@
 Remark 2: We note that the results of Theorem 4 concerning the uniqueness set are independent from the stability properties of the representation.
-In order to achieve stability which is important for reconstruction, it is required to construct a wavelet operator using multiple scales $t_{j},j=1,...K$, as proposed in [15], which will yield a collection of $nK$ wavelets.
-More specifically, we can express the function $\varphi$ using multiple scales $t_{j}$ (here we replace the previous notation of scale $s_{j}$ with $t_{j},j=1,..K$, not to confuse with nodes $s\in S$).
+In order to achieve stability which is important for reconstruction, it is required to construct a wavelet operator using multiple scales $t_{j},j=1,...J$, as proposed in [18].
+More specifically, we can express the function $\varphi$ using multiple scales $t_{j}$ (here we replace the previous notation of scale $s_{j}$ with $t_{j},j=1,..J$, not to confuse with nodes $s\in S$).
 Then for a fixed scale $t_{j}$ we have that the SGW is given by $||\psi_{\varphi}(t_{j})||=\left(\sum_{i=1}^{N}\sum_{k=0}^{K}|a_{t_{j},k}({\cal L}^{k}\varphi(i))|^{2}\right)^{1/2}$ and
@@
 In a similar way to the arguments provided in Theorem 3 we can choose coefficients $a_{t_{j},k}$ associated with the Laplacian polynomial such that the inequality $||\varphi||\leq\Lambda_{\psi}||\psi_{{\varphi}}||$ is satisfied.
-17 Sampling approach from network features
-Our approach revolves around the strategic sampling of edges, emphasizing their importance in the overall network structure.
-We introduce innovative method for this purpose, such as one leveraging edge betweenness centrality (EBC).
-In these proposed methods, our approach to edge sampling is guided by the assessed significance of edges with respect to the topological structure of the graph.
-This methodology extends random sampling techniques commonly employed in contemporary graph embedding, which typically rely on information derived solely from sparse graph connectivity, encompassing nodes within a 1-hop distance on the graph.
-To elaborate on our sampling approach, we extend the definition of the network $G=(V,\mathbf{W})$ to a triplet $G=(V,\mathbf{W},\gamma)$ such that $\gamma:V\rightarrow(0,1]$ is some probability distribution over $V$ (or $\mathbf{W}$) providing the significance of each node/edge.
-We choose to estimate $\gamma$ directly from the graph network using kernel density estimation (KDE).
-The kernel density estimator of the distribution ${\gamma}({{V})}$, is given by
-$\hat{\gamma}(x)=\frac{1}{Nh}\sum_{i\in V}{\frac{\mathbf{K}(\bar{v}(i)-x)}{h}}$ (29)
-Where $\bar{v}$ accumulate statistics measuring the extent of diffusion spread using SGW among nodes, $\mathbf{K}$ is symmetric kernel function (chosen to be a smooth Gaussian symmetric function), $h>0$ is the smoothing bandwidth that controls the amount of smoothing.
-In a similar way, one can also construct a probability measure $\gamma:V\times V\rightarrow(0,1]$ describing the significant of an edge.
-In this case, we choose to calculate the distribution from the edge betweenness centrality measurements (EBC), $\left\{w_{i,j}^{BE}|i,j\in V,i\sim j\right\}$ (see Eq.
-31) where we use kernel density estimation (KDE) to estimate the density distribution.
-Let $\gamma:V\times V\rightarrow\mathbb{R}^{+}$ denote the probability distribution over edges in the graph $G=(V,\mathbf{W})$, using the distribution of EBC, (EBC is defined in Eq.
-30).
-We denote the estimated distribution calculated directly from the EBC measurements as $\gamma({\mathbf{W^{BE}})}$.
-Formally, the kernel density estimator of distribution of EBC, ${\gamma}({\mathbf{W^{BE}})}$ is given by
-$\hat{\gamma}(x)=\frac{1}{N(e)h}\sum_{i\sim j}{\frac{\mathbf{K}(w_{i,j}^{BE}-x)}{h}}$ (30)
-Where $N(e)$ is a fixed number of positive edges to be sampled.
-The advantage of using KDE is that it is a non-parametric density estimator, which does not require assumption that the underlying density function is from a parametric family.
-Note that KDE works by estimating the density at each point (node or edge) as a weighted sum of the densities of neighboring points, with the weights given by a kernel function $\mathbf{K}$.
-In our experiments, we used the Gaussian kernel for simplicity, but one can use other kernel choices.
-Figure 4: Sampling from Betweenness Centrality distribution
-17.1 Sample from Network Features using Edge Betweenness Centrality
-Motivation: We employ Edge Betweenness Centrality (EBC) based sampling to create a strategy that samples edges from the EBC distribution.
-This enables us to optimize embeddings by predominantly using edges with low EBC, which are concentrated in dense clusters.
-In contrast, high-EBC edges typically occur in transitions between clusters and form a sparser set that is less frequently sampled.
-During optimization with SGD and negative sampling, these high-EBC edges play a crucial role in revealing the global structure of the network and connections between different clusters.
-Consider the illustration in Fig.
-4, where edges marked with black arrows correspond to edges with high edge betweenness centrality values, indicating their importance in connecting different dense clusters.
-Sampling edges with low EBC assists in identifying dominant local structures, while infrequently sampling high betweenness centrality edges is critical for understanding large clusters and the network’s global structure.
-  Input: $G=(V,\mathbf{W})$, kernel density function $\mathbf{K}$, bandwidth parameter $h$, $N_{e}$ for the number of edge samples.
-  Step 1: Construct the centrality graph $G=(V,\mathbf{W}^{BE})$ using Eq. (31).
-  Step 2: Utilize $\mathbf{K}$, $\bar{v}$, and $\mathbf{W}^{BE}$ in Kernel Density Estimation (Eq. 30) to estimate $\hat{\gamma}$.
-1  Step 3: Initialize a list of positive edges $C_{e}=\left\{\right\}$. Set $i=1$.
-  Step 4: While $i\leq N_{e}$ do:
-  Step 5: Sample an edge $w_{ij}$ using $\hat{\gamma}$ , $w_{ij}\sim\hat{\gamma}(\mathbf{W}^{BE})$.
-  Step 6: Add $w_{ij}$ to $C_{e}$, and and set $i=i+1$.
-  Step 7: Repeat steps 4-6. STOP when $i=N_{e}$
-  Output: Set of edges $C_{e}$ for positive samples in Algorithm 1.
-Algorithm 2 Edge Betweenness Centrality Sampling
-The Edge Betweenness Centrality:
-Edge Betweenness Centrality (EBC) quantifies edge importance based on the fraction of shortest paths that pass through it.
-For a graph with vertices and edge weights, the EBC of an edge $(i,j)$ is given by:
-$w_{i,j}^{BE}=\sum_{k\neq t}\frac{\sigma_{kt}(w_{ij})}{\sigma_{kt}}$ (31)
-Here, $\sigma_{kt}$ is the number of shortest paths from node $k$ to node $t$, and $\sigma_{kt}(w_{ij})$ is the number of those paths that include the edge $w_{ij}$.
-The centrality graph $G=(V,\mathbf{W}^{BE})$ is constructed using these EBC values, modifying the edges of the original graph $G$.
-Sampling Strategy of EBC using KDE estimator:
-Let $\gamma:V\times V\rightarrow\mathbb{R}^{+}$ denote the probability distribution over edges in the graph $G=(V,\mathbf{W})$, using the distribution of EBC.
-We denote the estimated distribution calculated directly from the EBC measurements as $\gamma({\mathbf{W^{BE}})}$.
-Given the set of edge betweenness centrality $\left\{w_{i,j}^{BE}|i,j\in V,i\sim j\right\}$ as an input to the Kernel Density Estimation (KDE) estimator, the density function is estimated automatically.
-With a fixed number of positive edges $N(e)$ to be sampled, we sample $N(e)$ edges from the estimated distribution.
-Although we used the Gaussian kernel in our experiments, other kernel choices are viable.
-In EBC based-sampling, we sample from the distribution of edge betweenness centrality measurements.
-This approach prioritizes the sampling of edges based on their importance in the graph structure.
-By treating edges as bottlenecks, we sample high betweenness centrality edges less frequently, thereby revealing clusters.
-Edge betweenness centrality serves as a measure of an edge’s bottleneck potential, guiding our SGD sampling and prioritizing edges based on their relevance and significance in capturing underlying relationships and connectivity patterns in the graph.
-Combined with the proposed encoding and initial embedding of multi-scale graph structure, the approach provides a comprehensive representation of the data.
+Remark 3: Optimal Choice of coefficients $a_{k}$ for Minimizing $\Lambda_{\psi}$ for optimal choice of $\Lambda_{\psi}$ in the multi-scale setting.
+To find the coefficients $a_{t_{j},k}$ that minimize $\Lambda_{\psi}$ where:
+$\Lambda_{\psi}=\frac{1}{\sqrt{\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{a_{t_{j},k}^{2}}{\Lambda^{2k}}}}$ (20)
+subject to the constraint: $\sum_{j=1}^{J}\sum_{k=0}^{K}a_{t_{j},k}^{2}=1$, we optimize the coefficients $a_{t_{j},k}$, which control how much of the signal at scale $t_{j}$ and power $k$ is captured by the SGW operator.
+We use the Lagrangian $\mathcal{F}$, incorporating the energy constraint into the optimization problem and Define:
+$\mathcal{F}(a_{t_{1},0},\dots,a_{t_{J},K},\mu)=\frac{1}{\sqrt{\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{a_{t_{j},k}^{2}}{\Lambda^{2k}}}}+\mu\left(\sum_{j=1}^{J}\sum_{k=0}^{K}a_{t_{j},k}^{2}-1\right)$ (21)
+[…]
+Taking partial derivatives, solving for $a_{t_{j},k}$, and differentiating $\mathcal{L}$ with respect to each coefficient $a_{t_{j},k}$:
+$\frac{\partial\mathcal{F}}{\partial a_{t_{j},k}}=\frac{-a_{t_{j},k}/\Lambda^{2k}}{\left(\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{a_{t_{j},k}^{2}}{\Lambda^{2k}}\right)^{3/2}}+2\mu a_{t_{j},k}=0.$ (22)
+rearranging the equation above we obtain:
+$\frac{a_{t_{j},k}}{\Lambda^{2k}}=2\mu a_{t_{j},k}\left(\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{a_{t_{j},k}^{2}}{\Lambda^{2k}}\right)^{3/2}$
+assuming $a_{t_{j},k}\neq 0$, we can divide both sides of the equation by $a_{t_{j},k}$, yielding:
+$\frac{1}{\Lambda^{2k}}=2\mu\left(\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{a_{t_{j},k}^{2}}{\Lambda^{2k}}\right)^{3/2}$ (23)
+thus the sum $\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{a_{t_{j},k}^{2}}{\Lambda^{2k}}$ is proportional to the inverse of $\Lambda^{2k}$, implying that the optimal wavelet coefficients should decay proportionally to the inverse powers of the Laplacian eigenvalues $a_{t_{j},k}\propto\frac{1}{\Lambda^{k}}$.
+Next, applying the energy constraint: $\sum_{j=1}^{J}\sum_{k=0}^{K}a_{t_{j},k}^{2}=1$ and substituting $a_{t_{j},k}=\frac{1}{\Lambda^{k}}$, we obtain:
+$\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{1}{\Lambda^{2k}}=1$ (24)
+Thus, the normalized optimal coefficients $a_{t_{j},k}$ are:
+$a_{t_{j},k}=\frac{\frac{1}{\Lambda^{k}}}{\sqrt{\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{1}{\Lambda^{2k}}}}.$ (25)
+Note that the wavelet energy is distributed across all scales and powers of the Laplacian, with a decay that prioritizes low-frequency components (small $k$).
+Using the form of the coefficients $a_{t_{j},k}$ above, the expression for $\Lambda_{\psi}$ is:
+$\Lambda_{\psi}=\frac{1}{\sqrt{\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{1}{\Lambda^{4k}}}}$ (26)
+which is the minimized value of $\Lambda_{\psi}$, showing that the optimal choice of coefficients reduces the overall constant, improving the localization and efficiency of the SGW operator.
+Thus, the optimal choice of wavelet coefficients $a_{t_{j},k}$ that minimizes $\Lambda_{\psi}$ in the multi-scale SGW operator is proportional to $\frac{1}{\Lambda^{k}}$, prioritizing low-frequency components.

### v5

@@
 We theoretically show that in Paley-Wiener spaces on combinatorial graphs, the spectral graph wavelets operator provides greater flexibility and control over smoothness compared to the Laplacian operator, motivating our approach.
-An additional key advantage of the proposed embedding is its ability to establish a correspondence between the embedding and input feature spaces, enabling the derivation of feature importance.
+A key advantage of the proposed embedding is its ability to establish a correspondence between the embedding and input feature spaces, enabling the derivation of feature importance.
 We validate the effectiveness of our graph embedding framework on multiple public datasets across various downstream tasks, including clustering and unsupervised feature importance.
 1 Introduction
-Graph Embeddings and Manifold Learning [29, 3, 34, 35, 5] are pivotal in analyzing complex data structures prevalent in diverse machine learning applications.
-The emergent representations from these techniques offer insights into the underlying data structure in conjunction with downstream tasks like clustering and visualization and are particularly valuable when labels are unavailable or unreliable.
-However, limitations of these techniques include reliance on low frequencies of the graph Laplacian and $k$-Nearest Neighbors graph connectivity, which provide limited information, including in recent contrastive learning methods such as Uniform Manifold Approximation and Projection (UMAP).
-To address these limitations, this paper proposes three main contributions.
-(1) A framework employing multi-scale graph representation using 3D based tensor and optimization using contrastive learning which enhances the expressiveness of an embedding by optimizing low and high-frequency information, capturing intricate data details.
-(2) An application of manifold interpretation to feature importance allows each dimension of our proposed embedding to be uniquely traced back and associated with the original features.
-(3) A characterization of the theoretical properties of the representation power of the spectral graph wavelets (SGW) operator by considering functions sampled from the Paley-Wiener spaces [27] on combinatorial graphs and a demonstration of the flexibility and better control over smoothness properties of the SGW operator in comparison to the Laplacian operator.
-A key distinction between our approach and many manifold learning methods such as UMAP is that we leverage multi-scale graph representation (using SGW) to increase the dimensionality of the data thus enhancing data representation learning.
-We use stochastic gradient descent (SGD) to optimize the embedding space and employ a novel contrastive learning approach using a 3D-tensor to capture both low and high frequencies.
-Generally, manifold learning methods perform nonlinear dimensionality reduction by embedding the data into lower-dimensional spaces, although some approaches may embed the data into higher dimensions [16, 31].
-A common limitation of these and other graph embedding methods is their lack of direct connections to the input features, which restricts their ability to produce interpretable results.
-This can be crucial in fields like finance, where behavioral analysis must be summarized using key features to guide decisions, or in biology, where interpretable features are essential for understanding complex systems.
-Our proposed approach establishes a mapping between the original features and the constructed optimized embeddings and enables the measurement of the importance of the original features with respect to the embedding space.
-To the best of our knowledge, current manifold learning techniques produce embeddings lacking explicit interpretation.
-In contrast, our approach demonstrates the ability to generate interpretable embedding representations, while also being competitive with the state-of-the-art graph embeddings across applications such as finance, vision, and biology.
+Graph Embeddings and Manifold Learning ([29, 3, 34, 35, 5]) play a pivotal role in analyzing complex data structures encountered in a wide range of machine learning applications.
+The representations learned by these techniques uncover the intrinsic geometry of the data and support downstream tasks such as clustering and visualization—especially in scenarios where labels are unavailable or unreliable.
+Most manifold learning methods typically perform nonlinear dimensionality reduction to embed data into lower-dimensional spaces, while some also explore high-dimensional embeddings ([16, 31]).
+Nonetheless, most such methods lack a direct connection to the input features, preventing the evaluation of individual feature contributions and resulting in embeddings with limited interpretability.
+This limitation is critical in domains like finance, where key behavioral indicators must be traced back to individual features, or in biology, where interpretability is vital for uncovering mechanisms within complex systems.
+To address these limitations, this paper makes three primary contributions:
+1.
+We introduce a framework that leverages multi-scale graph representation using a 3D tensor and contrastive learning.
+This enhances the expressiveness of the embedding by jointly optimizing over both low- and high-frequency components, enabling the capture of fine-grained data structure.
+2. We propose a manifold-based interpretation method for feature importance, allowing each dimension of the learned embedding to be directly linked back to a unique original feature.
+3.
+We characterize the theoretical representational capacity of the Spectral Graph Wavelet (SGW) operator by analyzing functions from Paley-Wiener spaces ([27]) on combinatorial graphs.
+This analysis demonstrates that SGW provides greater flexibility and smoother control compared to traditional Laplacian-based operators.
+Our method, MS-IMAP, is designed to associate each embedding coordinate with a unique input feature, thereby enabling the estimation of feature importance relative to the learned embedding space.
+This interpretability arises from a deliberate design choice: each embedding coordinate corresponds to a single, unique input feature and is optimized through spectral graph wavelets to align with the underlying graph structure.
+Unlike standard manifold learning and graph embedding methods, where each embedding dimension typically blends information from all input features, MS-IMAP preserves the semantic identity of each feature by assigning it to a distinct embedding dimension.
+Another key distinction of our approach—compared to both classical ([3]) and modern techniques ([25]), which typically rely on the low-frequency components of the graph Laplacian and thus capture only coarse structural information—is the use of multi-scale graph representations via Spectral Graph Wavelets (SGW) within a contrastive learning framework.
+This allows our method to retain or even expand the dimensionality of the data, yielding richer and more expressive representations while preserving interpretability.
+We further employ stochastic gradient descent (SGD) with a novel 3D-tensor-based contrastive objective to optimize embeddings that capture both global and local graph structure.
+We empirically demonstrate that applying feature importance techniques to MS-IMAP embeddings yields more informative feature subsets for clustering than applying them to raw input data.
+This provides strong evidence of semantic alignment—often absent in nonlinear methods such as deep networks or conventional manifold learning.
+MS-IMAP achieves interpretable and semantically faithful embeddings, while remaining competitive with state-of-the-art graph embedding techniques across a range of tasks.
+Figure 1: Illustration of the proposed method.
+(Left) Initial encoding using SGW in (b) and regularized embedding in (c).
+(Right) Feature embedding space representation, which has a one-to-one correspondence with the input features, can be used along with clustering for interpretability models, such as feature importance.
 2 Related Work
-Extensive efforts have been dedicated to exploring non-linear dimensionality reduction aimed at achieving a low-dimensional embedding that preserves the underlying manifold structure.
-Early manifold learning techniques include Locally Linear Embedding (LLE) [29], Laplacian Eigenmaps [3], Isomap [34], Diffusion Maps [4], and t-distributed Stochastic Neighbor Embedding (t-SNE) [35].
-Balancing local and global structure in the resulting embeddings is challenging in the presence of noise.
-To address the challenge, researchers have introduced de-noising methods [20, 12, 8, 11, 7, 13, 9, 10].
-These techniques seek to enhance algorithms robustness, thereby broadening the applicability of these techniques in real-world scenarios.
-UMAP [25] has proven effective in providing vivid visualizations and cluster separation, yet its reliance on Laplacian Eigenmaps initialization and negative sampling optimization has limitations.
-Laplacian Eigenmaps initialization focuses on low-frequency signal overlooking higher-frequency patterns.
-It uniformly samples edges without considering topological importance, overemphasizing graph connectivity and introducing further distortion.
-Recent work [6] further highlights that negative sampling has a significant impact on the effective loss and introduces distortions with t-SNE and UMAP which motivates the exploration of alternate sampling schemes.
-Other relevant studies include [38] which investigates competing negative sampling strategies, [32] on generalization error bounds , and [23] on topological and sampling bias due to high degree nodes.
-The lack of explicit mapping linking the original high dimensional dataset to its low dimensional embedding has prevented their use in areas where interpretability is critical.
-Establishing this link clearly would lead to a better understanding of the original feature space and improve feature selection, understanding of the resulting visualizations, and robustness to noise [21, 15, 39].
-SGW serve as an efficient tool for localization in both the spectral and vertex domains, playing a crucial role in our method (see Section 3.1 for details).
-SGW are constructed by applying a set of bandpass filters to the eigenvalues of the graph Laplacian, enabling multi-resolution representation of graph signals.
-Also related to our approach are graph embedding and multi-scale framework using Diffusion Wavelets [28] that embed the data in higher dimensions such as the Geometric Scattering for Graph Data Analysis [17] and the Graph Scattering Transform [16] that extract graph features, which are then employed in graph classification tasks.
+Extensive efforts have been dedicated to exploring manifold learning methods that aim to produce low-dimensional embeddings while preserving the intrinsic structure of high-dimensional data.
+Classical manifold learning techniques include Locally Linear Embedding (LLE) [29], Laplacian Eigenmaps ([3]), Isomap ([34]), Diffusion Maps ([4]), and t-distributed Stochastic Neighbor Embedding (t-SNE) ([35]).
+One of the central challenges in these methods is maintaining a balance between local and global structure, especially in the presence of noise.
+To address this, various denoising strategies have been developed ([20, 12, 8, 11, 7, 13, 9, 10]), aiming to improve robustness and extend the applicability of these techniques to real-world, noisy datasets.
+UMAP ([25]) has shown strong empirical performance for visualization and cluster separation.
+It typically employs Laplacian Eigenmaps initialization, which emphasizes low-frequency graph signals while neglecting higher-frequency patterns that may carry important information.
+Recent work ([6]) has shown that negative and positive sampling in UMAP significantly influences the effective loss function.
+These insights have led to studies on alternative sampling schemes ([38]), bounds on generalization error in embedding methods ([32]), and biases introduced by high-degree nodes during graph construction and sampling ([23]).
+Manifold learning and non-linear dimensionality reduction serve purposes beyond visualization; one of their key goals is to learn representations that capture intrinsic geometric structure and approximate geodesic distances.
+The learned embedding space dimensionality can go beyond two or three dimensions.
+However, increasing dimensionality does not necessarily improve robustness: noise and the curse of dimensionality can introduce instability.
+These methods also suffer from computational limitations.
+Moreover, to compute embedding in higher dimensions many algorithms—such as Laplacian Eigenmaps and UMAP—require computing large portions of the Laplacian spectrum, making them computationally intensive for large datasets.
+Our method addresses these issues by producing interpretable and efficient embeddings while leveraging spectral information across multiple frequency scales.
+A key limitation of many manifold learning and graph embedding methods is the absence of an explicit and interpretable mapping, which prevents evaluating the contribution of each input feature to the resulting low-dimensional embedding.
+This gap restricts their utility in domains where interpretability is crucial.
+Providing such a link can enable improved feature selection, clearer interpretation of embeddings, and greater robustness to noise ([21, 15, 39]).
+Spectral Graph Wavelets (SGW) are central to our method (see Section 3.1), offering localization in both spectral and vertex domains.
+SGW apply a family of bandpass filters to the graph Laplacian’s spectrum, allowing for multi-resolution analysis of graph signals and providing richer representations across frequency bands.
+Our approach is also related to graph embedding methods that exploit multiscale structure, such as Diffusion Wavelets ([28]), Geometric Scattering for Graph Data Analysis ([17]), and the Graph Scattering Transform ([16]), which extract graph features for downstream tasks like classification.
+These methods often embed data in higher-dimensional spaces to improve representation power.
 2.1 Preliminaries
-Consider a set of points $\mathbf{x}=\left\{\mathbf{x}_{i}\right\},\,i=1,...N,\mathbf{x}_{i}\in\mathbb{R}^{D}$, which are sampled from an unknown manifold $M\subset\mathbb{R}^{D}$.
-An undirected, weighted graph $G=(V,\mathbf{W})$, is constructed over $\mathbf{x}$, where $V$ corresponds to the nodes and $\mathbf{W}$ to the set of edges on the graph.
-The adjacency matrix $\mathbf{W}=(w_{ij})$ consists of the weights $w_{ij}$ between node $i$ and node $j$.
-The weights can be chosen using different techniques, such as the Gaussian kernel function between the points $\mathbf{x}_{i},\mathbf{x}_{j}$ using the $k$ nearest neighbors of $\mathbf{x}_{i}$, kNN$(\mathbf{x}_{i})$ (which denotes the $k$ nearest neighbors) or adaptive graph construction as proposed in [25].
-$\mathbf{L}$ denotes the combinatorial graph Laplacian, defined as $\mathbf{L}=\mathbf{D}-\mathbf{W}$, with $\mathbf{D}$ the diagonal degree matrix with entries $d_{ii}=d(i)$.
-The degree $d(i)$ of vertex $i$ is defined as the sum of weights of edges that are connected to $i$.
-The normalized Laplacian is defined as ${\mathbf{L}}_{N}=\mathbf{D}^{-\frac{1}{2}}\mathbf{L}\mathbf{D}^{-\frac{1}{2}}=\mathbf{I}-\mathbf{D}^{-\frac{1}{2}}\mathbf{W}\mathbf{D}^{-\frac{1}{2}}$ and its real eigenvalues are in the interval $[0,2]$.
-The eigenvalues and eigenvectors of $\mathbf{L}$ are $\lambda_{1},\ldots,\lambda_{N}$ and $\mathbf{\phi}_{1},\ldots,\phi_{N}$, respectively.
-In this work we use the normalized Laplacian $\mathbf{L}_{N}$, which from now on we refer to as $\mathbf{L}$ for simplicity.
-We will denote the matrix representations of the eigenvectors and eigenvalues of the graph Lapacian $\mathbf{L}$ as $\Phi,\lambda$, respectively.
-Our approach assumes that all graph signals $f_{r}\in\mathbb{R}^{N}$ used to create the embedding correspond to the input coordinates, $\mathbf{x}=\left(f_{1},f_{2},..,f_{r},..f_{D}\right)$.
-The output embedding space provides a multi-scale representation approximating the intrinsic manifold coordinates.
+Consider a set of points $\mathbf{x}=\left\{\mathbf{x}_{i}\right\},i=1,...N,\mathbf{x}_{i}\in\mathbb{R}^{D}$, where each $\mathbf{x}_{i}\in\mathbb{R}^{D}$, sampled from an unknown smooth manifold $M\subset\mathbb{R}^{D}$.
+To model the geometric structure of the data, we construct an undirected, weighted graph $G=(V,\mathbf{W})$ over the point set $\mathbf{x}$, where $V$ is the set of nodes and $\mathbf{W}=(w_{ij})$ is the weighted adjacency matrix encoding edge weights between nodes.
+The weights $w_{ij}$ can be computed using various techniques.
+A common approach is to apply a Gaussian kernel to pairs of points $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ within the $k$ nearest neighbors of $\mathbf{x}_{i}$, denoted as $\text{kNN}(\mathbf{x}_{i})$.
+In this work, we adopt the adaptive Gaussian kernel construction proposed in [25], though other graph construction methods may also be employed.
+Let $\mathbf{L}$ denote the combinatorial graph Laplacian defined as $\mathbf{L}=\mathbf{D}-\mathbf{W}$, where $\mathbf{D}$ is the diagonal degree matrix with entries $d_{ii}=d(i)$, and the degree $d(i)$ of vertex $i$ is the sum of the weights of all edges incident to $i$.
+We use the normalized graph Laplacian, defined as:
+$\mathbf{L}_{N}=\mathbf{D}^{-1/2}\mathbf{L}\mathbf{D}^{-1/2}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{W}\mathbf{D}^{-1/2}$
+which has real eigenvalues in the interval $[0,2]$.
+For simplicity, we refer to $\mathbf{L}_{N}$ as $\mathbf{L}$ for the remainder of the paper.
+Let $\lambda_{1},\ldots,\lambda_{N}$ and $\boldsymbol{\phi}_{1},\ldots,\boldsymbol{\phi}_{N}$ denote the eigenvalues and corresponding eigenvectors of $\mathbf{L}$.
+We denote the matrix of eigenvectors as $\Phi\in\mathbb{R}^{N\times N}$ and the diagonal matrix of eigenvalues as $\Lambda=\text{diag}(\lambda_{1},\ldots,\lambda_{N})$.
+In our setting, each input coordinate $f_{r}\in\mathbb{R}^{N}$ — viewed as a graph signal—is defined across the $N$ nodes of the graph.
+The original data matrix $\mathbf{x}$ can thus be expressed as a collection of such signals: $\mathbf{x}=\left(f_{1},f_{2},..,f_{r},..f_{D}\right)$.
+Our goal is to derive a multi-scale embedding from these signals that approximates the intrinsic manifold coordinates, using spectral information derived from the graph Laplacian.
 2.2 Multi-scale representations using SGW
-In the last two decades, several multi-scale representations using irregular graphs have been suggested, including SGW [18] and Diffusion Wavelets [28].
-In this work we focus on the multi-scale graph transform utilizing SGW.
-SGW provides a natural way of trading off between spectral and spatial resolution, since the SGW coefficients are localized in both domains.
-These wavelets are constructed using a kernel function $g(\mathbf{L})$ operator, which acts on a function by modulating each of its Fourier modes [18], that helps in capturing a trade off between vertex (spatial) and spectral localization.
-Spatial localization in the graph domain is implicitly controlled by a single wavelet scale parameter defined in the spectral domain, such that the more the vertex is localized in the vertex domain, the spectral band is wider.
-To represent a signal ${f}\in\mathbb{R}^{N}$ in multiple scales $S=[s_{1},s_{2},..s_{K}]$, the SGW transform is constructed as follows: Assume that $g(\lambda)$ is a filter in the spectral domain.
-Let $\delta_{i}\in\mathbb{R}^{N}$ be the delta function centred at the vertex $i\in G:\delta_{i}(j)=1\,\,\mbox{if}\,\,i=j$, and $\delta_{i}(j)=0$ otherwise.
-Given a bandpass filter $g(\lambda)$ and a wavelet centered on node $i$ at scale $s$, the wavelets $\psi(s,i),i=1,..N$ are calculated by applying them to the delta function on a single vertex $i$, given by $\psi(s,i)=\Phi g(s\lambda)\Phi^{T}\delta_{i}$.
-The value of $\psi(s,i)$ with respect to a vertex $m$ can be written as $\psi(s,i)(m)=\sum_{l=1}^{N}g(s\lambda_{l})\phi_{l}^{*}(i)\phi_{l}(m)$.
-Given a graph signal ${f}\in\mathbb{R}^{N}$, the SGW coefficient at node $i$ and scale $s$ can be expressed as follows:
+In the past two decades, several multi-scale representations for data on irregular graphs have been developed, including Spectral Graph Wavelets (SGW) [18] and Diffusion Wavelets [28].
+In this work, we focus on the SGW-based multi-scale graph transform.
+SGW offers a principled way to balance spectral and spatial resolution, as its coefficients are localized in both domains.
+These wavelets are constructed using a kernel function operator $g(\mathbf{L})$, which acts on a signal by modulating each of its spectral components (i.e., Fourier modes) [18].
+This design enables a trade-off between vertex-domain (spatial) and frequency-domain (spectral) localization.
+The spatial localization is implicitly controlled by a single scale parameter defined in the spectral domain: greater localization in the vertex domain corresponds to a broader spectral bandwidth.
+To represent a signal $f\in\mathbb{R}^{N}$ at multiple scales $S=[s_{1},s_{2},\ldots,s_{K}]$, the SGW transform is defined as follows.
+Let $g(\lambda)$ be a bandpass filter in the spectral domain.
+Let $\delta_{i}\in\mathbb{R}^{N}$ denote the delta function centered at vertex $i\in G$, where $\delta_{i}(j)=1$ if $i=j$, and $\delta_{i}(j)=0$ otherwise.
+A wavelet centered at node $i$ and scale $s$ is then given by:
+$\psi(s,i)=\Phi g(s\lambda)\Phi^{T}\delta_{i}$
+.The value of $\psi(s,i)$ at vertex $m$ can be written explicitly as:
+$\psi(s,i)(m)=\sum_{l=1}^{N}g(s\lambda_{l})\phi_{l}^{*}(i)\phi_{l}(m)$
+. Given a graph signal $f\in\mathbb{R}^{N}$, the SGW coefficient at node $i$ and scale $s$ is defined as:
 $\psi_{f}(s,i)=\sum_{l=1}^{N}g(s\lambda_{l})\hat{f}(\lambda_{l})\phi_{l}(i)$ (1)
-Fast computation using Chebyshev polynomials: Directly computing the SGW coefficients is computationally intensive - $O(N^{3})$ for $N$ points.
-Instead, Hammond et al.
-[18] suggested computing the SGW using a fast algorithm based on approximating the scaled generating kernels through low-order polynomials of $\mathbf{L}$ applied to the input data (details in supplementary material).
+Fast computation using Chebyshev polynomials:
+Direct computation of SGW coefficients is computationally expensive, requiring $O(N^{3})$ operations for $N$ nodes.
+To address this, Hammond et al.
+[18] proposed an efficient algorithm based on approximating the scaled generating kernels $g(s\lambda)$ using low-order Chebyshev polynomials of the Laplacian $\mathbf{L}$, which are applied directly to the input signal.
+This significantly reduces computational cost (see Appendix for further details).
 3 Our Proposed framework: Multi-Scale IMAP
@@
 Our method imposes a differentiable structure on the mapping $h:\mathbf{x}\rightarrow\psi_{\mathbf{x}}$, supported on a discrete graph $G=(V,\mathbf{W})$, where $\psi_{\mathbf{x}}$ represents the encoded multi-scale graph transform of $\mathbf{x}$.
-Multi-scale IMAP consists of two main components:
-Step 1: Features representation encoding using multi-scale graph representation.
-In this step, our approach constructs a multi-scale graph representation by incorporating feature signals through the SGW Transform across multiple scales and graph frequencies.
+Multi-scale IMAP consists of two main steps: In Step 1, our approach constructs a multi-scale graph representation by incorporating feature signals through the SGW Transform across multiple scales and graph frequencies.
 We suggest a 3D tensor-based approach (detailed in Algorithm 1), where we construct a 3D tensor with dimensions ${K}\times{N}\times{D}$ to provide an effective representation for encoding multi-scale transform.
 We leverage the encoded 3D tensor structure to align the optimized embedding with all scales simultaneously to enforce differentiable structure of the transformed features.
-Note that an alternative would be to use simple concatenation, all of the features transformed for all scales are represented by concatenating all the SGW coefficients for each point in a vector form.
-For completeness we also describe this alternative in the Appendix.
 The 3D tensor-based approach (detailed in Algorithm 1) provides the most effective representation for the following reasons:
-(i) Joint and simultaneous optimization: using the 3D tensor-based method enables the joint and simultaneous optimization of all scales for each sample by leveraging the SGW coefficients of each spectral band within the cross-entropy loss framework while adhering to the fixed graph structure (Equation (3.2)).
+(i) Joint and simultaneous optimization: using the 3D tensor-based method enables the joint and simultaneous optimization of all scales for each sample by leveraging the SGW coefficients of each spectral band within the cross-entropy loss framework while adhering to the fixed graph structure.
 This approach effectively balances coarse and fine-grain resolutions, with summation performed only after the contrastive learning optimization process is complete.
@@
 This combination of optimization efficiency and enhanced interpretability highlights the clear advantages of the 3D tensor-based representation.
-Step 2: Optimization design using multi-scale network structure.
-We integrate the multi-scale representation within the SGD optimization framework, simultaneously leveraging both low and high-frequency information.
+We note that an alternative to the 3D tensor-based representation approach is to use simple concatenation, where all features are transformed using SGW and each data point is represented by a vector formed by concatenating all corresponding SGW coefficients.
+For completeness we describe this alternative in the Appendix.
+In Step 2, We integrate the multi-scale representation within the SGD optimization framework, simultaneously leveraging both low and high-frequency information.
 This integration leads to fine-grained manifold regularization and improved robustness in downstream tasks.
@@
 Note that each dimension in the embedding space is constructed using a single feature in the original feature set, which is an essential characteristic that can be leveraged for interpreting graph embeddings.
-3D-Tensor Encoding: The 3D-Tensor Encoding method concatenates the multi-scale representation for all features at a fixed scale, generating a matrix representation $\psi_{\mathbf{x}}(s_{j},:,:)\in\mathbb{R}^{D}\times\mathbb{R}^{N}$ for each scale $s_{j}$.
+3D-Tensor Encoding: The 3D-Tensor Encoding method concatenates the multi-scale representation for all features at a fixed scale, generating a matrix representation $\psi_{\mathbf{x}}(s_{j},:,:)\in\mathbb{R}^{D}\times\mathbb{R}^{N}$ for each scale $s_{j}$, encoding SGW coefficients across features and nodes.
 After concatenating all $f_{i}$ at a fixed scale $s_{j}$, we have
 $\psi_{\mathbf{x}}(s_{j},:,:)=\psi_{{f}_{1}}(s_{j},:)\|\psi_{{f}_{2}}(s_{j},:)\|...\|\psi_{{f}_{D}}(s_{j},:)$
-[…]
-$\mathbf{c}(\psi_{{f}_{i}}(s_{k},:,:),\psi_{{f}_{j}}(s_{k},:,:))=\psi_{{f}_{i}}(s_{k},:,:)\left|\right|\psi_{{f}_{j}}(s_{k},:,:)$
-The optimization can be performed separately for each scale $s_{j}$, or jointly for all scales $s_{j}$ using the 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$ We detail the proposed feature representation encoding methods in the pseudo code algorithm 1.
+where we designate the concatenation using $\left|\right|$, with $\mathbf{c}(\psi_{{f}_{i}}(s_{k},:,:),\psi_{{f}_{j}}(s_{k},:,:))=\psi_{{f}_{i}}(s_{k},:,:)\left|\right|\psi_{{f}_{j}}(s_{k},:,:)$.
+The optimization can be performed separately for each scale $s_{j}$, or jointly for all scales $s_{j}$ using the 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$ .
+We detail the proposed feature representation encoding methods in the pseudo code algorithm 1.
 Input: Set of points $\left\{\mathbf{x}_{i}\right\}_{i=1}^{N}$.
-  Output: Initial Node Embeddings ${\psi}_{\mathbf{x}}$
+Output: Initial Node Embeddings ${\psi}_{\mathbf{x}}$ .
 Step 1: Construct $G=(V,\mathbf{W})$ from $\left\{\mathbf{x}_{i}\right\}_{i=1}^{N}$.
@@
 Step 3: Compute $\lambda_{\mbox{max}}(\mathbf{L})$.
-1  Step 4: Initial Embedding construction:
+1Step 4: Initial Embedding construction:
 2For $r=1,..,D$ associated with the feature signals $\left\{f_{r}\right\}$ do:
 3Compute $\psi_{{f}_{r}}(s_{j},:)$ at scales $s_{j},j=1,...K$ using Chebyshev approximation.
-  Step 5: Concatenate $\psi_{{f}_{r}}(s_{j},:)$ for $r=1,...,D$ in a matrix form $\psi_{\mathbf{x}}(s_{j},:,:)$ for each fixed scale $s_{j}$, representing all scales using 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$.
+Step 5: Concatenate $\psi_{{f}_{r}}(s_{j},:)$ for r=1,,Dr=1,...,D in a matrix form $\psi_{\mathbf{x}}(s_{j},:,:)$ for each fixed scale $s_{j}$, representing all scales using 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$.
 Algorithm 1 Encoding multi-scale structure in the optimized embeddings
@@
 We outline the steps of the proposed approach.
-3D-Tensor Based Optimization for graph Embedding
-The encoded manifold representation is a 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$. An SGD update rule at iteration $t$, applied to the 3D tensor, is:
+3D-Tensor Based Optimization for graph Embedding Encoding the representation in a 3D tensor offers two key advantages.
+First, is the computational efficiency: the update rule at each iteration, associated with a pair of nodes, is performed jointly and simultaneously for all scales.
+This avoids the need for independent optimization per scale, which would significantly increase computational complexity.
+The second is optimization alignment: jointly optimizing across scales ensures alignment in the optimization process for all scales with respect to a fixed pair of nodes (positive or negative sample).
+This approach is more effective than optimizing scales independently, as it maintains consistency and improves overall performance.
+While SGW captures global graph structure, each embedding update remains feature-specific, preserving the one-to-one correspondence between input features and embedding dimensions, which is essential for interpretability.
+The encoded manifold representation is a 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$.
+An SGD update rule at iteration $t$, applied to the 3D tensor, is:
 $(\tilde{\psi}_{\mathbf{x}}^{(t+1)})_{i,j,k}=(\tilde{\psi}_{\mathbf{x}}^{(t)})_{i,j,k}-\alpha\frac{\partial\mathcal{L}}{\partial(\psi_{\mathbf{x}})_{i,j,k}}$ (2)
@@
 $\left.+(1-w_{ij})\log\frac{1-w_{ij}}{1-v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}}\right)$ (3)
-where $v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}=\frac{1}{1+||\psi_{\mathbf{x}_{i}}(s_{k},:,:)-\psi_{\mathbf{x}_{j}}(s_{k},:,:)||^{2}}$.
+[…]
 We apply the optimization based on SGD with respect to each scale $s_{k}$, where the gradient of the loss is approximated by:
@@
 where $r_{ij}(s_{k})=\psi_{\mathbf{x}_{i}}(s_{k},:,:)-\psi_{\mathbf{x}_{j}}(s_{k},:,:)$.
+Feature updates: Our implementation employs an optimization process that operates jointly across all scales using the 3D tensor, as described in the SGD update rule for iteration $t$ in Eq.(2).
 The optimization process uses SGD with positive and negative sampling for all scales $s_{k}$ given a pair of positive or negative pair of nodes $i,j$.
+At each iteration $t$, associated with either a positive sample (an edge) or a negative sample (a pair of unconnected nodes), the update rule from Eq.(2) is applied.
+The derivative of the loss function, as detailed in Eq.(4), is computed for all scales $(s_{1},...s_{K})$, with the sign in Eq.(2) depending on whether the sample is positive or negative.
+Once the optimization concludes, the final embedding is derived from the 3D tensor by summing over the scales, as specified in Eq.(6).
+Recall that the term $\psi_{\mathbf{x}}(s_{k},:,:)\in\mathbb{R}^{D\times N}$ is a slice of the 3D SGW tensor at scale $s_{k}$, representing spectral graph wavelet coefficients for all features across all nodes $N$.
+The specific vector $\psi_{\mathbf{x}}(s_{k},j,:)\in\mathbb{R}^{N}$ corresponds to the SGW coefficients at scale $s_{k}$ for input feature $j$ across all nodes.
+Note that in Eq.(2), the $j$-th embedding coordinate—aligned with input feature jj—is updated via SGD using two components: (1) the current embedding value and (2) a structural term involving $v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}$, which integrates the graph structure and multiscale information from the SGW tensor.
+At each iteration, updates for positive/negative node pairs are applied jointly across all scales, but each scale’s contribution is weighted by its energy concentration—ensuring that more informative scales have greater influence.
 We compute the final embedding by summing up all the optimized embeddings with respect to each scale $s_{k}$
 $\tilde{\psi}_{\mathbf{x}}=\sum_{k}\tilde{\psi}_{\mathbf{x}}(s_{k},:,:)$ (6)
-Note that the dimensionality of the final embedding
-$\tilde{\psi}_{\mathbf{x}}\in\mathbb{R}^{D}\times\mathbb{R}^{N}$, maintaining the same dimensionality as the original input, and thus there exists a one-to-one correspondence between the coordinates of the input features $\mathbf{x}$ and $\tilde{\psi}_{\mathbf{x}}$.
+The dimensionality of the final embedding is $\tilde{\psi}_{\mathbf{x}}\in\mathbb{R}^{D}\times\mathbb{R}^{N}$, maintaining the same dimensionality as the original input, and thus there exists a one-to-one correspondence between the coordinates of the input features $\mathbf{x}$ and $\tilde{\psi}_{\mathbf{x}}$.
 This correspondence ensures that each dimension of the feature space is directly mapped to a single dimension in the embedding space $\tilde{\psi}_{\mathbf{x}}$.
-Our optimization process involves using SGD with positive and negative sampling, similar to recent graph embeddings methods [33, 25].
+Our optimization process involves using SGD with positive and negative sampling, similar to recent graph embeddings methods ([33, 25]).
 Positive edge samples are associated with attraction (first term on the right hand side), while negative samples (second term on the right hand side) refer to a pair of nodes that are not connected in the graph, which create repulsion among dissimilar points.
-The optimization can be performed jointly for all scales using the 3D tensor.
-Specifically, our implementation employs an optimization process that operates jointly across all scales, as described in the SGD update rule for iteration $t$ in Eq.
-(2).
-At each iteration $t$, associated with either a positive sample (an edge) or a negative sample (a pair of unconnected nodes), the update rule from Eq.(2).
-is applied.
-The derivative of the loss function, as detailed in Eq.
-(4), is computed for all scales $(s_{1},...s_{K})$, with the sign in Eq.(2) depending on whether the sample is positive or negative.
-Once the optimization concludes, the final embedding is derived from the 3D tensor by summing over the scales, as specified in Eq.(6).
-Encoding the representation in a 3D tensor offers two key advantages.
-First, is the computational efficiency: the update rule at each iteration, associated with a pair of nodes, is performed jointly and simultaneously for all scales.
-This avoids the need for independent optimization per scale, which would significantly increase computational complexity.
-The second is optimization alignment: jointly optimizing across scales ensures alignment in the optimization process for all scales with respect to a fixed pair of nodes (positive or negative sample).
-This approach is more effective than optimizing scales independently, as it maintains consistency and improves overall performance.
-4 Theoretical results: Sampling set for Smooth Manifolds with functions defined over Paley-Wiener Spaces
-In this section, we characterize the theoretical properties of the representation power of the SGW operator by considering functions sampled from the Paley-Wiener spaces [27] on combinatorial graphs.
-Pesenson introduced the Paley-Wiener spaces and demonstrated that Paley-Wiener functions of low type are uniquely determined by their values on certain subsets, known as uniqueness sets $U$.
-We show that the SGW operator can represent functions $f$ within the Paley-Wiener space more efficiently than the graph Laplacian operator $\mathbf{L}$.
-This efficiency is demonstrated by showing that the SGW operator is more effective in representing functions with larger bandwidth $\omega$ in the Paley-Wiener spaces (i.e., with higher frequencies) using subsets of nodes from the uniqueness sets $U$.
-To characterize the representation properties of functions defined over $PW_{\omega}(G)$, we employ the notion of the $\Lambda$-set, introduced by Pesenson which is central to our investigation.
-Formally, the Paley-Wiener space of $\omega$-bandlimited signals is defined as $PW_{\omega}(G)=\left\{f\mid\hat{f}(\lambda)=0\,\,\forall\,\,\lambda>\omega\right\}$.
-The space $L_{2}(G)$ is defined as the Hilbert space of all complex-valued functions, and $L_{2}(S)$ is defined as the space of all functions from $L_{2}(G)$ with support in $S$: $L_{2}(S)=\left\{\varphi\in L_{2}(G)\mid\varphi(v)=0,v\in V(G)\setminus S\right\}$.
-The $\Lambda$-set is defined as follows: a set of vertices $S\subset V$ is a $\Lambda(S)$-set if any $\varphi\in L_{2}(S)$ satisfies the inequality $||\varphi||\leq\Lambda||\mathbf{L}\varphi||$ for some constant $\Lambda(S)>0$.
-The infimum of all $\Lambda>0$ for which $S$ is a $\Lambda$-set is denoted by $\Lambda$.
-The ability of the SGW operator to efficiently represent functions $f\in PW_{\omega}(G)$ can be summarized in the following theorem, which highlights the role of the $\Lambda_{\psi}$-set with respect to the operator $\psi$.
-We show that any signal $f\in PW_{\omega}(G)$, where $\lambda_{1}\leq\omega<\Omega_{G}$ for some $\Omega_{G}<\lambda_{N}$, can be uniquely represented by its samples on the uniqueness set $U$ using the SGW operator.
-Under certain conditions related to the SGW operator, its associated $\Lambda_{\psi}$-set is smaller than the $\Lambda$-set associated with the Laplacian operator.
-Theorem 1
-Let $G=(V,\mathbf{W})$ be a connected graph with $n$ vertices and $f\in PW_{\omega}(G)$ for $\lambda_{1}\leq\omega<\lambda_{\text{max}}$.
-The SGW operator $\psi$ can be constructed such that the set $S$ is a $\Lambda_{\psi}$-set and the set $U=V\setminus S$ is a uniqueness set for any space $PW_{\omega}(G)$ with $\omega<1/\Lambda_{\psi}$ and $\Lambda_{\psi}<\Lambda$ for any $\varphi\in L_{2}(S)$, where $\Lambda$ is the $\Lambda$-set of the Laplacian operator.
-The proof is provided in the Appendix.
-Figure 1: Illustration of the proposed method.
-(Left) Initial encoding using SGW in (b) and regularized embedding in (c).
-(Right) Feature embedding space representation, which has a one-to-one correspondence with the input features, can be used along with clustering for interpretability models, such as feature importance.
-5 Experimental Results
-We demonstrate the advantage of our approach over existing approaches using both synthetic and real datasets.
-We evaluate our method by testing its clustering performance using the output embedding and comparing it to several representative methods, including UMAP, t-SNE, Diffusion Maps, ISOMAP, PHATE [26], PaCMAP [36], TriMap [1] and HeatGeo [22].
-For UMAP, t-SNE, PHATE, PaCMAP, TriMap and HeatGeo, we hyperparameter tune the performance on each dataset, running the method five times for each hyperparameter configuration and taking the average.
-Then we report the best average score.
-Additional details are provided in the supplementary material.
-Evaluation metrics: We test performance using the Adjusted Rand Index (ARI) and Adjusted Mutual Information (AMI). In all experiments we used $k$-means to cluster the data in the embedding space.
-5.1 Real Datasets
-We study the performance of MS-IMAP compared to other methods for real datasets.
-We chose a mix of datasets from varied fields: the Census dataset [14] is a financial dataset containing information about individuals extracted from the 1994 US Census; the Zilionis dataset [40] is a biological dataset containing single-cell sequencing data from different types of cells; the Animals with Attributes (AWA) [37] dataset is an image dataset containing images of animals; The Human Activity Recognition (HAR) [2] dataset which consists of sensor data collected from subjects performing six different activities: walking, walking upstairs, walking downstairs, sitting, standing, and lying.
-Each activity is represented as a feature vector of 561 attributes and 7352 samples; and the Cora dataset [30, 24] is a graph dataset of a citation network consisting of scientific publications.
-Note that we only use the features without the natural graph associated with Cora.
-More information about each dataset can be found in the supplementary material.
-As shown in Table 1, MS-IMAP is competitive and performs the best overall, while also providing an interpretable embedding through its one-to-one correspondence with the original features—unlike the other competing graph embedding methods.
-Runtime and Computational Complexity The computational complexity of Multi-Scale IMAP is of $O(NDlog(N))$.
-The execution time of our method with Python code implementation with 32 cores Intel Xeon 8259CL running at 2.50Ghz and 256GB of RAM on the Cancer QC data-set of 48,969 samples and 306 features took 27 mins, on Cora 5 min, on the AWA dataset 1.35 mins and on Census 7 min.
-The increase in running times compared to methods such as UMAP is primarily due to our extension of manifold learning to include interpretability.
-Our approach generally requires the embedding dimensionality to be at least equal to the input dimensionality of the features, which adds computational complexity but ensures meaningful feature representation and interpretation.
-Our method is practical for handling datasets with millions of points and dozens to hundreds of features within a few hours.
-In fields such as finance and genomics/biology, where typical dataset sizes range from under 100,000 to several million data points, analysis is often conducted offline once the graph embedding is computed.
-Runtimes of up to a few hours are generally acceptable and do not present a significant challenge, making the method well-suited for real-world applications.
-Additional details are provided in the Appendix.
-Data HAR Census Zilionis AWA Cora
-Method / Accuracy ARI AMI ARI AMI ARI AMI ARI AMI ARI AMI
-Features 0.45 0.58 0.11 0.15 0.48 0.63 0.07 0.16 0.01 0.07
-ISOMAP 0.55 0.65 0.18 0.09 0.44 0.56 0.52 0.60 0.12 0.17
-Diffusion Maps N/A N/A 0.06 0.16 0.01 0.02 0.22 0.42 0.03 0.07
-UMAP 0.61 0.73 0.23 0.15 0.52 0.74 0.74 0.82 0.33 0.36
-t-SNE 0.71 0.77 0.17 0.17 0.41 0.70 0.73 0.80 0.21 0.27
-PHATE 0.55 0.66 0.20 0.09 0.65 0.72 0.66 0.73 0.12 0.22
-TriMap 0.59 0.72 0.2 0.14 0.62 0.77 0.74 0.81 0.15 0.2
-PaCMAP 0.62 0.75 0.24 0.15 0.55 0.76 0.75 0.81 0.21 0.26
-HeatGeo 0.6 0.69 0.15 0.10 N/A N/A 0.65 0.74 0.17 0.20
-MS-IMAP 0.67 0.8 0.23 0.15 0.70 0.76 0.74 0.81 0.33 0.36
-Table 1: Clustering results comparison using ARI and AMI on the Census, Zilionis, AWA, and Cora datasets. Note HeatGeo had code execution issues on the Zilionis dataset.
+Remark 1: While we adopt the SGW formulation by Hammond et al. ([18]), our framework is compatible with other spectral graph wavelet and multiscale representations.
+Remark 2: While the manifold hypothesis suggests the data lie near a lower-dimensional manifold, this does not preclude learning meaningful structure in the ambient space.
+Our method incorporates high-frequency graph information via spectral graph wavelets, enhancing stability and local structure capture, without requiring expensive eigenspectrum computations as in Laplacian Eigenmaps or UMAP.
 4 Feature Importance using Explicit Correspondence with Embedding Features
-The correspondence established between the features and the embedding space offers a valuable tool for assessing feature importance.
-We explore two alternative methods for measuring feature importance, leveraging the constructed embedding space, where each dimension can be traced back to its corresponding feature in the original space.
-Specifically, we test these methods in an unsupervised or self-supervised setting where the subset of features are selected without using label information.
-6.1 Assessing Embedding Feature Importance using the Laplacian Score
-The first method utilizes the Laplacian Score (LS) [19], which assesses the importance of each feature by examining its correlation with the eigenvalues of the graph Laplacian.
+Feature importance is one of the simplest yet most fundamental tools for model interpretability.
+It associates input features with model outputs, highlighting which features most strongly influence the learned representations.
+We show that feature subsets selected from the MS-IMAP embedding lead to superior clustering performance compared to those selected from the input space using the same feature importance methods.
+Unlike competing approaches, MS-IMAP retains semantic alignment and supports direct correspondence between embedding dimensions and original input features.
+This makes MS-IMAP an interpretable alternative to black-box models.
+We empirically demonstrate that applying feature importance techniques to the MS-IMAP embedding yields more informative feature subsets for clustering than applying the same techniques directly to the raw input.
+This provides evidence of the semantic consistency preserved by our method—often absent in traditional nonlinear embeddings.
+We explore two alternative approaches for estimating feature importance, leveraging the explicit correspondence in the embedding space where each dimension maps back to a specific input feature.
+These methods are evaluated in an unsupervised or self-supervised setting, where feature selection is performed without access to label information.
+4.1 Assessing Embedding Feature Importance using the Laplacian Score:
+The first method utilizes the Laplacian Score (LS) ([19]), which assesses the importance of each feature by examining its correlation with the eigenvalues of the graph Laplacian.
 The significance of the variable corresponding to the $l$-th coordinate, represented by the feature $f_{l}\in\mathbb{R}^{N}$, is inferred from the relevance of its corresponding coordinate $l$ in the embedding space, $\tilde{\psi}_{\mathbf{x}_{l}}$.
-The Laplacian Score is calculated with respect to the embedding features $\tilde{\psi}_{\mathbf{x}}$ using the graph Laplacian $\mathbf{L}$ and the degree matrix $\mathbf{D}$.
-To compute the Laplacian Score, we first subtract the mean and then calculate it as follows: $L_{s}(\tilde{\psi}_{\mathbf{x}})_{l}=\frac{(\tilde{\psi}_{\mathbf{x}})_{l}^{T}\mathbf{L}(\tilde{\psi}_{\mathbf{x}})_{l}}{(\tilde{\psi}_{\mathbf{x}})_{l}^{T}\mathbf{D}(\tilde{\psi}_{\mathbf{x}})_{l}}$
+Computing the Laplacian score with respect to the embedding features $\tilde{\psi}_{\mathbf{x}}$.
+The Laplacian score is calculated with respect to embedding features $\tilde{\psi}_{\mathbf{x}}$ using the Laplacian graph $\mathbf{L}$ and the degree matrix $\mathbf{D}$.
+To compute the Laplacian Score, we first subtract the mean and then calculate it as follows:
+$L_{s}(\tilde{\psi}_{\mathbf{x}})_{l}=\frac{(\tilde{\psi}_{\mathbf{x}})_{l}^{T}\mathbf{L}(\tilde{\psi}_{\mathbf{x}})_{l}}{(\tilde{\psi}_{\mathbf{x}})_{l}^{T}\mathbf{D}(\tilde{\psi}_{\mathbf{x}})_{l}}$
 Smaller scores indicate that the embedding feature $(\tilde{\psi}_{\mathbf{x}})_{l}$ has a greater projection onto the subspace of eigenvectors corresponding to the smallest eigenvalues, signifying higher importance concerning the global graph structure.
 Therefore, the feature importance of $(\tilde{\psi}_{\mathbf{x}})_{l}$ can be directly interpreted as the importance of the corresponding original feature.
+Although the graph is typically constructed using local neighborhoods (e.g., via k-nearest neighbors), the Laplacian operator reflects global smoothness over the entire data manifold.
 4.2 Assessing Importance of Embedding Features using Mutual Information
-In the second method, we propose a self-supervised approach that uses Mutual Information (MI) to estimate the importance of each dimension in the embedding space, which can then be explicitly linked to the original features.
+We also propose a self-supervised approach that uses Mutual Information (MI) to estimate the importance of each dimension in the embedding space, which can then be explicitly linked to the original features.
 First, we apply $k$-means to cluster the MS-IMAP embedding space and use the resulting clusters as pseudo-labels.
 We then compute MI to quantify the shared information between features and these pseudo-labels, similar to how MI would be computed with ground truth labels.
-This step allows MI to measure how much each feature contributes to the cluster assignments, highlighting those most informative in distinguishing clusters.
+This step allows MI to measure how much each feature contributes to cluster assignments, highlighting those that are the most informative to distinguish clusters.
 The MI-Based Feature Importance identify features that are informative about cluster assignments.
@@
 We summarize the MI-based unsupervised feature selection in Algorithm 2.
-In comparison to the MI based feature importance method, using the Laplacian score [19] we have that the embedding may have combined original features in a way that makes individual Laplacian Scores less indicative of feature importance.
-The Laplacian score prioritize neighborhood preservation, which might not directly translate to cluster discrimination efficacy.
-On the other hand, while the proposed MI based feature importance identifies features that are directly informative for distinguishing clusters, one of its possible limitations may be relaying on the quality of pseudo-labels, which may not align with true clusters.
-  Input: MS-IMAP embeddings $\psi_{\mathbf{x}}$, number of clusters $k$
-1  Output: MI scores for each feature
+4.3 Comparison: Laplacian Score (LS) vs. Feature Importance Based-Mutual Information
+Input: MS-IMAP embeddings $\tilde{\psi}_{\mathbf{x}}$, number of clusters $k$
+1Output: MI scores for each feature
 Step 1: Apply k-means clustering on the embedding space $\psi_{\mathbf{x}}$ to obtain cluster labels
-[…]
+2      $\text{clusters}=\text{k-means}(\tilde{\psi}_{\mathbf{x}},k)$
 Step 2: Use the resulting cluster labels as pseudo-labels
-[…]
+      $\text{pseudo\_labels}=c(\tilde{\psi}_{\mathbf{x}})$
 Step 3: Initialize an empty list to store mutual information (MI) values for each feature
-[…]
+3      $\text{mi\_scores}=[]$
 Step 4: For each feature $i$ in the original data or embedding space:
-            Step 4.1: Compute mutual information between feature $\psi_{\mathbf{x}}(i)$ and the pseudo-labels
-4            Step 4.2: $MI\_score=\text{MI}(\psi_{\mathbf{x}}(i),c({\psi_{\mathbf{x}}})$)
-5            Step 4.3: Append $MI\_score$ to mi_scores
+      Step 4.1: Compute mutual information between feature $\tilde{\psi}_{\mathbf{x}}(i)$ and the pseudo-labels
+4      Step 4.2: $MI\_score=\text{MI}(\tilde{\psi}_{\mathbf{x}}(i),c(\tilde{\psi}_{\mathbf{x}})$)
+5      Step 4.3: Append MI_scoreMI\_score to mi_scores
 Step 5: Output the list of MI scores
@@
 Algorithm 2 MI-based Unsupervised Feature Selection
-Figure 2: Clustering results on Cora using a subset of features based on feature importance, comparing MS-IMAP embedding space and the input features.
+(a) Cora dataset (ARI)
+(b) Zillions dataset (ARI)
+Figure 2: Clustering results using Adjusted Rand Index (ARI) on Cora and Zillions datasets with a subset of features based on feature importance, comparing the MS-IMAP embedding space and the input features.
+(a) Cora dataset (AMI)
+(b) Zillions dataset (AMI)
+Figure 3: Clustering results using Adjusted Mutual Information (AMI) on Cora and Zillions datasets with a subset of features based on feature importance, comparing the MS-IMAP embedding space and the input features.
+The Laplacian Score (LS) and Mutual Information (MI)–based feature importance are two complementary approaches to unsupervised feature selection, with LS focusing on geometric smoothness over a graph and MI targeting feature relevance with respect to clustering structure.
+Laplacian Score (LS) evaluates how smoothly a feature varies over a graph constructed from the data.
+It is a fully unsupervised, geometry-driven criterion that does not require clustering or pseudo-labels.
+A feature receives a low score if it varies slowly and consistently across the graph.
+This makes LS particularly effective for identifying features that align with smooth manifold structure.
+However, LS also has several limitations.
+It is not task-aware: it does not assess whether a feature is relevant for a specific downstream objective such as clustering or classification.
+Moreover, LS can overrate irrelevant but structured noise—for instance, a smoothly varying but non-informative feature may receive a deceptively low score simply because it is smooth.
+In contrast, Mutual Information–based feature importance measures how much information a feature provides about cluster assignments.
+It quantifies the statistical dependency between each feature and the resulting clusters (e.g., from k-means or MS-IMAP embeddings).
+A key strength of MI is that it is task-aware: it is self-supervised (or can be seen as supervised with respect to pseudo-labels), capturing which features are actually useful for distinguishing between clusters.
+MI scores are also directly interpretable—a high MI score indicates that the feature contributes to separating groups—and the method is generally robust to irrelevant noise features, which receive low MI if they do not correlate with clustering structure.
+However, MI is sensitive to the quality of the clustering used to generate pseudo-labels—poor or unstable clustering can diminish its effectiveness.
+5 Experimental Results
+We evaluated our approach in both synthetic and real datasets in two application domains.
+First, we demonstrate the use of MS-IMAP embeddings for feature importance—a novel application that highlights the interpretability of our method.
+Next, we assess the performance of MS-IMAP in clustering tasks using the full embedding space, comparing it against a range of manifold learning-based embedding techniques.
+Datasets We study the performance of MS-IMAP compared to other methods for real datasets.
+We chose a mix of datasets from varied fields: the Census dataset ([14]) is a financial dataset containing information about individuals extracted from the 1994 US Census; the Zilionis dataset ([40]) is a biological dataset containing single-cell sequencing data from different types of cells; the Animals with Attributes (AWA) ([37]) dataset is an image dataset containing images of animals; The Human Activity Recognition (HAR) ([2]) dataset which consists of sensor data collected from subjects performing six different activities:; and the Cora dataset ([30, 24]) is a graph dataset of a citation network consisting of scientific publications.
+Note that we only use the features without the natural graph associated with Cora (see more information about each dataset in the Appendix).
 5.1 Assessing Feature Importance via Embedding Correspondence
-We demonstrate the application of MS-IMAP for feature importance by evaluating clustering performance using $k$-means as a function of the number of features, based on the most significant features identified by two methods: the Laplacian score and Mutual Information-based feature selection.
-We compared the clustering results (using ARI and AMI) when $k$-means is applied to different subsets of embedding features selected by the Laplacian score and MI-based methods, with results obtained by applying the same feature selection methods directly to the original feature space.
-Unlike most manifold learning and graph embedding methods, MS-IMAP produces embeddings where the dimensions can be directly mapped to the original feature space, enabling the assessment of individual feature contributions.
-On the Cora dataset, Figure 2 shows that for both feature selection methods, the clustering accuracy obtained using the MS-IMAP embedding is significantly higher than when using the raw features across different subsets of features.
-This improvement is due to the robustness of the MS-IMAP embedding, while still offering interpretability by linking the embedding dimensions to the original features.
+We demonstrate the application of MS-IMAP for feature importance selection by evaluating clustering performance on subsets of the most significant features.
+Feature selection is performed with respect to both the input space and the MS-IMAP embedding space, using two alternative unsupervised methods: Laplacian Score and Mutual Information-based selection.
+We compare clustering results obtained by applying \$k\$-means to subsets of features selected from the embedding, against those selected directly from the original feature space using the same methods.
+Unlike most manifold learning and graph embedding techniques, MS-IMAP produces embeddings in which each dimension corresponds directly to an input feature.
+This one-to-one mapping enables direct assessment of individual feature contributions, preserving interpretability.
+Evaluation metrics: Clustering performance is measured using Adjusted Rand Index (ARI) and Adjusted Mutual Information (AMI).
+In all experiments, $k$-means is applied in the embedding space.
+On the Cora and Zillions datasets, Figures 2(a), 2(b), 3(a), and 3(b) show that, for both feature selection methods and for both evaluation metrics, clustering accuracy is consistently higher when using the MS-IMAP embedding compared to the original features—across varying feature subset sizes.
+This improvement highlights the robustness of the MS-IMAP embedding, while maintaining interpretability through its feature-wise alignment.
+Additional experiments on the AWA and HAR datasets, presented in the Appendix, further confirm these gains.
+5.2 Clustering Output Embedding Space
+We evaluate our method by testing its clustering performance using the full output embedding and comparing it to several representative methods, including UMAP, t-SNE, Diffusion Maps, ISOMAP, PHATE ([26]), PaCMAP ([36]), TriMap ([1]) and HeatGeo ([22]).
+While most of these methods perform dimensionality reduction, we emphasize that dimensionality reduction is a methodological design choice, not an inherent advantage.
+Many manifold learning techniques default to low-dimensional embeddings (e.g., 2D or 3D) due to computational constraints such as eigendecomposition, yet this often comes at the cost of reduced expressiveness and, particularly for nonlinear methods, diminished interpretability.
+However, since most baseline manifold learning methods we compare rely on nonlinear dimensionality reduction, we include clustering results based on the embedding space of the Graph Scattering Transform ([16]) for completeness.
+The Graph Scattering Transform, which is primarily used in graph classification and semi-supervised learning tasks, can be adapted for unsupervised learning via simple vector-based concatenation.
+We note that the Graph Scattering Transform offers less interpretability than MS-IMAP, which preserves a one-to-one correspondence between input features and embedding dimensions.
+Additionally, the increased dimensionality of the embeddings produced by the Graph Scattering Transform can lead to higher computational costs.
+As shown in Table 1, MS-IMAP consistently delivers stable and competitive performance, ranking best or second-best across all benchmarks.
+In contrast, state-of-the-art methods often underperform or show instability on at least one dataset.
+In addition, MS-IMAP provides an interpretable embedding through its one-to-one correspondence with the original features—unlike the other competing graph embedding methods.
+Runtime and Computational Complexity The computational complexity of Multi-Scale IMAP is of
+$O(NDlog(N))$.
+Experiments were performed on Virtual Server containers with 32 cores Intel Xeon 8259CL running at 2.50Ghz and 256GB of RAM.
+Our method is practical for handling datasets with millions of points and dozens to hundreds of features within a few hours, making the method well-suited for real-world applications.
+For example, on the Cancer QC data-set of 48,969 samples and 306 features took 27 mins.
+Additional details and running times on all datasets are provided in the Appendix.
+Note that the increase in running times compared to methods such as UMAP is primarily due to our extension of manifold learning to include interpretability.
+Data HAR Census Zilionis AWA Cora
+Method ARI AMI ARI AMI ARI AMI ARI AMI ARI AMI
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 1: Clustering results comparison using ARI and AMI on the Census, Zilionis, AWA, and Cora datasets.
+Each entry report the mean clustering accuracy and standard deviation.
+The best performance is bolded.
+6 Theoretical results: Sampling set for Smooth Manifolds with functions defined over Paley-Wiener Spaces
+[…]
+Theorem 3 Let $G=(V,\mathbf{W})$ be a connected graph with $n$ vertices and $f\in PW_{\omega}(G)$ for $\lambda_{1}\leq\omega<\lambda_{\text{max}}$.
+The SGW operator $\psi$ can be constructed such that the set $S$ is a $\Lambda_{\psi}$-set and the set $U=V\setminus S$ is a uniqueness set for any space $PW_{\omega}(G)$ with $\omega<1/\Lambda_{\psi}$ and $\Lambda_{\psi}<\Lambda$ for any $\varphi\in L_{2}(S)$, where $\Lambda$ is the $\Lambda$-set of the Laplacian operator.
+The proof is provided in the Appendix (see Theorem 3 in 8.13).
+Theorem 3 motivates our use of SGW within the proposed framework.
+Specifically, it shows that SGW operators provide more compact and effective representations of band limited functions in Paley-Wiener spaces on graphs than the traditional Laplacian operator.
 7 Discussion
-Identifying the key factors of high-dimensional datasets is a necessary ingredient to take advantage of unlabelled data in many practical applications.
-In this article, we introduce a novel contrastive learning framework for manifold learning via graph embeddings, capitalizing on both low and high-frequency information.
-In particular, we use SGW to construct a multi-scale graph representation of the underlying input feature space.
-We then use an SGD-based optimization scheme together with innovative strategies such as 3D Tensor encoding to derive the embeddings.
-To motivate the use of the spectral graph wavelet (SGW) for representation, we examine the theoretical properties of the SGW by studying functions in Paley-Wiener spaces on combinatorial graphs.
-We prove that the SGW operator provides a more effective representation through the concept of the $\Lambda$-set.
-Finally, we demonstrate that the embeddings are interpretable using two alternative methods for unsupervised and self-supervised feature importance, applied to both the embedding and original feature spaces.
-The construction of our methodology inherently links the original and transformed feature spaces, a capability that is notably absent in most existing non-linear manifold learning techniques, to the best of our knowledge.
-Limitations and future work include the following.
-Firstly, we rely on the assumption that the features, and thus the similarity measure, adequately capture information to reveal geodesic distances between manifold points.
-Secondly, utilizing the current optimization framework with SGD may present difficulties when extending to out-of-sample problems.
-Furthermore, although our embedding feature correspondence proves valuable in assessing feature importance through a global method, there remains potential for enhancement through the incorporation of local methods.
-Appendix
-8 Encoding Method 1
-[…]
-$\mathbf{c}(\psi_{{f}_{i}}(s_{k},:),\psi_{{f}_{j}}(s_{k},:))=\psi_{{f}_{i}}(s_{k},:)\left|\right|\psi_{{f}_{j}}(s_{k},:)$
-denoting the concatenation of the vectors corresponding to the multi-scale representation $\psi_{{f}_{i}}(s_{k},:)$ and $\psi_{{f}_{j}}(s_{k},:)$.
-For method 1 all features and all scales are concatenated together, the resulting matrix $\psi_{\mathbf{x}}$ can be represented as
-$\psi_{\mathbf{x}}=\psi_{{f}_{1}}(s_{1},:)\|\psi_{{f}_{2}}(s_{1},:)\|....\|\psi_{{f}_{D-1}}(s_{K},:)\|\psi_{{f}_{D}}(s_{K},:)$
-where $\psi_{\mathbf{x}}\in\mathbb{R}^{KD}\times\mathbb{R}^{N}$.
-1. Optimize Embedding Method 1:
-Given the encoded multi-scale representation $\psi_{\mathbf{x}}\in\mathbb{R}^{KD}\times\mathbb{R}^{N}$ perform optimization in the SGW domain, using the following fuzzy cross entropy loss function:
-$\mathcal{L}(\tilde{\psi}_{\mathbf{x}}|\mathbf{W})=\sum_{i,j}\left(w_{ij}\mbox{log}\frac{w_{ij}}{v_{ij}^{\psi_{\mathbf{x}}}}+(1-w_{ij})\mbox{log}\frac{1-w_{ij}}{1-v_{ij}^{\psi_{\mathbf{x}}}}\right)$ (7)
-where $v_{ij}^{\psi_{\mathbf{x}}}=\frac{1}{1+||\psi_{\mathbf{x}_{i}}-\psi_{\mathbf{x}_{j}}||^{2}}$.
-Dropping terms that do not depends on $\psi_{\mathbf{x}_{i}}$, the gradient of the loss is approximated by:
-$\frac{\partial\mathcal{L}(\tilde{\psi}_{\mathbf{x}}|\mathbf{W})}{\partial\psi_{\mathbf{x}_{i}}}\sim\sum_{j}w_{ij}v_{ij}^{\psi_{\mathbf{x}}}(r_{ij})-\sum_{j}\frac{1}{||r_{ij}||^{2}}v_{ij}^{\psi_{\mathbf{x}}}(\psi_{\mathbf{x}_{i}}-\psi_{\mathbf{x}_{j}})$ (8)
-where $r_{ij}=\psi_{\mathbf{x}_{i}}-\psi_{\mathbf{x}_{j}}$, and $\psi_{\mathbf{x}_{i}}$ corresponds to the vector associated with node $i$ multi-scale representation concatenated for all the graph signals and all spectral bands.
-since for method 1 all features and all scales are concatenated together,
-Figure 3: The encoding step of our framework is illustrated as a mapping between each coordinate of the input features and a corresponding dimension in the embedding space.
-This mapping is facilitated through our proposed approach, which aligns the input dimensions of the features with the latent embedding space.
-By establishing this correspondence, we acquire the capability to interpret each dimension in the embedding space.
-Such interpretation allows for various analyses, including providing feature importance relative to the latent space.
-9 Additional experiments: Embedding Feature Importance using Mutual Information and Laplacian Score
-We include additional experiments that demonstrate the application of MS-IMAP for feature importance by evaluating clustering performance using $k$-means as a function of the number of features (in addtion to the one found in the Section 6), based on the most significant features identified by the Laplacian score and Mutual Information-based feature selection.
-We compared the clustering results (using ARI and AMI) when $k$-means is applied to different subsets of embedding features selected by the Laplacian score and MI-based methods, with results obtained by applying the same feature selection methods directly to the original feature space.
-On the Zillions dataset, Figures 5 and 6 shows that for both feature selection methods, the clustering accuracy obtained using the MS-IMAP embedding is significantly higher than when using the raw features across different subsets of features.
-This improvement is due to the robustness of the MS-IMAP embedding, while still offering interpretability by linking the embedding dimensions to the original features.
-We also show in Figure 4 the clustering accuracy obtained using the MS-IMAP copared to using the raw features using AMI, demonstrating similar observations regarding the advantages of MS-IMAP.
-Figure 4: Clustering results (AMI) on Cora using a subset of features based on feature importance, comparing MS-IMAP embedding space and the input features.
-Figure 5: Clustering results (ARI) on Zillions using a subset of features based on feature importance, comparing MS-IMAP embedding space and the input features.
-Figure 6: Clustering results (AMI) on Zillions using a subset of features based on feature importance, comparing MS-IMAP embedding space and the input features.
-10 Dataset Details
-Two Moons.
-The two moons dataset depicts two interleaving half-circles.
-We sampled $N=600$ points and used a Gaussian noise level having standard deviation $0.12$.
-An example is show in Figure 7.
-More specifically, to produce these points we use Sci-kit Learn’s sklearn.datasets.make_moons function with $n\_samples=600$ and $noise=0.12$.
-In Figure 7, we show example clusterings for each of the methods mentioned in the Experimental Results section of the main paper.
-Census.
-From the UCI Machine Learning Repository [14], this dataset contains 14 features that are a mix of categorical, numerical, and binary.
-Such features include age, marital status, sex, etc.
-The goal is to predict whether a sample makes less than or equal to \$$\$50,000$, or strictly more.
-We use $32,561$ samples in our dataset.
-Lung Cancer.
-The Zilionis dataset is widely used, and consists of single-cell RNA sequencing data.
-It has $306$ features, and $48,969$ samples.
-The data has 20 classes corresponding to cell type.
-More can be found at [40]
-Animals with Attributes.
-The Animals with Attributes (AWA) dataset, contains 5,000 data points corresponding to 10 unseen classes, where the testing image features are obtained from the pre-trained ResNet architecture, with $D=2,048$ dimensions, and the semantic features are provided with $D=85$ dimensions.
-More information can be found in Section 4.1 of [37].
-Cora.
-The Cora dataset has 2,708 scientific publications categorized into seven classes.
-The network has 5,429 links and each publication is described as a binary word vector indicating the presence or absence of a word.
-The dictionary consists of 1,433 words.
-In our experiments, we do not use the given graph, and only use the features.
-[30, 24].
-HAR.
-The Human Activity Recognition (HAR) dataset consists of sensor data collected from 30 subjects performing six different activities: walking, walking upstairs, walking downstairs, sitting, standing, and lying.
-The data was recorded using accelerometer and gyroscope sensors embedded in a smartphone worn on the waist.
-Each activity is represented as a feature vector of 561 attributes and 7352 samples, extracted from the raw sensor signals through signal processing techniques.
-In our experiments, we do not use the temporal dependencies and only utilize the extracted features.
-Figure 7: Clustering results using two moons
-11 Experimental Results on two moons Synthetic Dataset
-11.1 Synthetic dataset
-Here we show further experiments on more data, where we assess the robustness of our method using the two moons dataset, which is a 2D manifold depicting two interleaving half-circles.
-We sampled $N=600$ points, and set the Gaussian noise standard deviation to $0.12$.
-While spectral based methods such as UMAP are effective under relatively “modest” noise levels, their performance deteriorates in the presence of larger amounts of noise.
-As shown in Table 1 under the Moons column, our approach is robust and correctly clusters most points despite the large noise level, and competitive with the competing methods.
-Data Dense-sparse clusters
-Method/Accuracy ARI AMI
-Features 0.24 0.18
-UMAP 0.54 0.51
-t-SNE 0.42 0.35
-ISOMAP 0.36 0.3
-Diffusion Maps 0.25 0.19
-PHATE 0.42 0.35
-HeatGeo 0.54 0.52
-MS-IMAP 0.89 0.87
-Table 2: Comparison of clustering performance on the two moons dataset.
-12 Hyperparameter tuning details
-We tune UMAP on each dataset in the Experimental Results section, by tuning over the parameter space in Table 3.
-Hyperparameter Set of values
-n_neighbors 2, 10, 15, 20, 30, 50, 100
-min_dist 0, 0.1, 0.5, 0.99
-n_components 2, 3
-Table 3: Space of parameters in which we tuned UMAP.
-12.1 TSNE Hyperparameter Tuning Details
-We tune TSNE on each dataset in the Experimental Results section, by tuning over the parameter space in Table 4.
-Hyperparameter Set of values
-perplexity 15, 30, 60
-early_exaggeration 12, 24
-n_components 2, 3
-Table 4: Space of parameters in which we tuned TSNE.
-12.2 PHATE Hyperparameter Tuning Details
-We tune PHATE on each dataset in the Experimental Results section, by tuning over the parameter space in Table 5.
-Hyperparameter Set of values
-knn 3, 5, 10
-decay (i.e. $\alpha$) 6, 10, 14, 18, 40
-t auto, 50, 100, 150, 200
-Table 5: Space of parameters in which we tuned PHATE.
-12.3 HeatGeo Hyperparameter Tuning Details
-We tune HeatGeo on each dataset in the Experimental Results section, by tuning over the parameter space in Table 6,
-Hyperparameter Set of values
-knn 5, 10, 15
-lap_type normalized, combinatorial
-harnack_regul 0, 0.5, 1
-Table 6: Space of parameters in which we tuned HeatGeo.
-13 Ablation Studies of Hyperparameters for MS-IMAP
-Here we study what effect the hyperparameters – the number of nearest neighbors, and the number of filters – have on the clustering performance of MS-IMAP with method 2.
-In table 7, we demonstrate the effect of varying the number of neighbors, between 10, 15, and 20 neighbors.
-We see that the results are mostly the same, thus showing MS-IMAP is robust on real datasets.
-Dataset Census Zilionis AwA
-Number of Neighbors / Accuracy ARI AMI ARI AMI ARI AMI
-10 0.22 0.15 0.71 0.78 0.73 0.80
-15 0.22 0.15 0.70 0.76 0.74 0.81
-20 0.22 0.15 0.70 0.77 0.71 0.79
-Table 7: Ablation study on MS-IMAP with encoding method 2, varying the number of neighbors. The number of filters is kept at 5.
-We also study the effect of varying the number filters. In table 8, we also see similar results when using different filters, showing the stability of MS-IMAP.
-Dataset Census Zilionis AwA
-Number of Filters / Accuracy ARI AMI ARI AMI ARI AMI
-5 0.22 0.15 0.70 0.76 0.74 0.81
-6 0.22 0.15 0.70 0.76 0.73 0.81
-7 0.22 0.15 0.72 0.78 0.74 0.81
-Table 8: Ablation study on MS-IMAP with encoding method 2, varying the number of filters. The number of neighbors is kept at 15.
-14 Ablation Studies of Hyperparameters for t-SNE, Isomaps, and Diffusion Maps
-We study the effect of varying the number of neighbors for the methods: t-SNE, Isomaps, and Diffusion Maps.
-For t-SNE this would be the perplexity hyperparameter, for Isomaps this would be the number of neighbors, and for diffusion maps, this would be the parameter that affects the width of the Gaussian kernel, i.e.
-$\exp(\cdot/\alpha)$.
-In Table 9, we see choosing a smaller perplexity of 15 does worse than the perplexity of 30 and 60.
-Dataset Census Zilionis AwA
-Perplexity / Accuracy ARI AMI ARI AMI ARI AMI
-15 0.03 0.04 0.37 0.67 0.69 0.76
-30 0.15 0.15 0.38 0.68 0.73 0.80
-60 0.17 0.17 0.38 0.69 0.73 0.80
-Table 9: Ablation study on t-SNE, varying the perplexity.
-In Table 10, we see a similar pattern to t-SNE, where choosing too low causes reduction in performance. But performance stabilizes around choosing the number of neighbors as 5-10+.
-Dataset Census Zilionis AwA
-Neighbors / Accuracy ARI AMI ARI AMI ARI AMI
-2 -0.05 0.02 0.41 0.55 0.43 0.58
-5 0.18 0.09 0.44 0.56 0.52 0.60
-10 0.18 0.09 0.44 0.57 0.51 0.62
-Table 10: Ablation study on Isomap, varying the number of neighbors.
-In Table 11, we find the ablation study for Diffusion Maps. Diffusion Maps perform the worst, but as with t-SNE and Isomap, benefits from considering more neighbors.
-Dataset Census Zilionis AwA
-alpha / Accuracy ARI AMI ARI AMI ARI AMI
-1 0 0 0 0.01 0.23 0.44
-2 0 0 0 0.01 0.22 0.43
-3 0.06 0.16 0 0.02 0.22 0.42
-Table 11: Ablation study on Diffusion Map, varying alpha which affects the width of the kernel.
-15 Fast computation using Chebyshev polynomials
-We provide additional details regarding the fast computation of SGW coefficients [18].
-Directly computing the SGW coefficients above requires calculating the entire eigensystem of the Laplacian, which is computationally intensive - $O(N^{3})$ for $N$ points.
-Instead, Hammond et al.
-[18] suggested computing the SGW using a fast algorithm based on approximating the scaled generating kernels through low-order polynomials.
-The wavelet coefficients at each scale are then computed as a polynomial of $\mathbf{L}$ applied to the input data, using approximating polynomials given by truncated Chebyshev polynomials.
-The Chebyshev polynomials $T_{k}(y)$ are computed using the recursive relations: $T_{k}(y)=2yT_{k-1}(y)-T_{k-2}(y)$ for $k\geq 2$, where $T_{0}=1$ and $T_{1}=y$.
-The SGW coefficients are then approximated using wavelet and scaling function coefficients as follows:
-$\psi_{f}(s_{j},i)\sim\left(\frac{1}{2}c_{j,0}f+\sum_{k=1}^{K}c_{j,k}\bar{T}_{j,k}(\mathbf{L})f\right)_{i}$ (9)
-where $c_{j,k},j>0$ are the Chebyshev coefficients and $\bar{T}_{j,k}$ are the shifted Chebyshev polynomials $\bar{T}_{k}(x)=T_{k}\left(\frac{x-a}{a}\right)$ for $x\in[0,\lambda_{\mbox{max}}]$, where $x=a(y+1)$, a = $\lambda_{\mbox{max}}$.
-The scaling function coefficients, which are corresponding to a low-pass filter operation, are approximated in a similar way using Chebyshev polynomials.
-Note that the scaling kernel function is a low pass filter $h$ satisfying $h(0)>0$ and $h(x)\rightarrow 0$ when $x\rightarrow\infty$.
-If the graph is sparse, we obtain a fast computation of the matrix-vector multiplication $\bar{T}_{j,k}(\mathbf{L})f$, where the computational complexity scales linearly with the number of points, resulting in a complexity of $O(N)$ for an input signal $f\in\mathbb{R}^{N}$.
-The SGWs efficiently map an input graph signal (a vector of dimension $N$) to $NK$ scaling and wavelet coefficients.
-16 Runtime and Computational Complexity
-We have performed experimental runtime studies on empirical datasets.
-Experiments were performed on Virtual Server containers with 32 cores Intel Xeon 8259CL running at 2.50Ghz and 256GB of RAM.
-On the AWA dataset of $5,685$ samples and $2,048$ features, the 3D based Tensor and the simple concatenation Method both run in 1.35 mins.
-Finally on the Cancer QC data-set of 48,969 samples and 306 features, Method 1 took 12.44 mins.
-The computational complexity of Multi-Scale UMAP is of $O(NDlog(N))$ for construction of the multi-scale representations which includes the $k$ nearest neighbor graph using k-d tree, the SGW transform which is $O(N)$ for each dimension of the manifold for sparse graphs.
-The optimization stage has a complexity which scale with the number of edges in the graph, which has a complexity of $O(kDN)$.
-17 Theoretical results: Sampling set for Smooth Manifolds with functions defined over Paley-Wiener Spaces
-In this work, we characterize the theoretical properties of the representation power of the SGW operator by considering functions sampled from the Paley-Wiener spaces [27] on combinatorial graphs.
-The Paley-Wiener spaces were introduced on combinatorial graphs in [27] and a corresponding sampling theory was developed which resembles the classical one.
-Pesenson proved in [27] that Paley-Wiener functions of low type are uniquely determined by their values on certain subgraphs (which are composed from a set of nodes known as the uniqueness sets) and can be reconstructed from such sets in a stable way.
-We demonstrate that the SGW operator can represent functions $f$ that reside in the Paley-Wiener space on combinatorial graphs more efficiently than the graph Laplacian operator $\mathcal{L}$.
-The effectiveness of the SGW operator representation in this case can be understood in several ways.
-In one way, by the ability of the SGW operator to accurately represent functions with larger bandwidth, i.e., $f\in PW_{\omega^{\prime}}(G)$ where $\omega<\omega^{\prime}$.
-We first summarize the main notions and definitions.
-The space $L_{2}(G)$ is the Hilbert space of all complex-valued functions $f:V\rightarrow\mathbb{C}$ with the following inner product $\left\langle f,g\right\rangle=\sum_{v\in G}f(v)\overline{g(v)}$ and the norm
-$||f||=\left(\sum_{v\in V}{|f(v)|^{2}}\right)^{1/2}$ (10)
-The Laplace (normalized) operator $\cal L$ is defined by the formula [27]:
-$\mathcal{L}f(v)=\frac{1}{\sqrt{d(v)}}\sum_{u\sim v}\left(\frac{f(v)}{\sqrt{d(v)}}-\frac{f(u)}{\sqrt{d(u)}}\right),f\in L_{2}(G)$ (11)
-In order to prove these results, we will first need the following definitions:
-Definition 1
-The Paley-Wiener space of $\omega$bandlimited signals $f\in L_{2}(G)$ is defined as follows:
-$PW_{\omega}(G)=\left\{f|\hat{f}(\lambda)=0\,\,\forall\,\,\lambda>\omega)\right\}$ (12)
-We consider a simple, undirected, unweighted, and connected graph $G=(V,\mathbf{W})$, where $V$ is its set of $N$ vertices and $\mathbf{W}$ is its set of edges.
-The degree of $v$ is number of vertices adjacent to a vertex $v$ is and is denoted by $d(v)$.
-We assume that degrees of all vertices are bounded by the maximum degree denoted as
-$d(G)=\mbox{max}_{v\in V}d(v)$ (13)
-The following definition [27] explains the uniqueness set.
-Definition 2
-A set of vertices $U\subset V$ is a uniqueness set for a space $PW_{\omega}(G)$ if for every two functions from $PW_{\omega}(G)$ that coincide on $U$, then they coincide on $V$.
-Definition 3
-For a subset $S\subset V$, denote $L_{2}(S)$ as the space of all functions from $L_{2}(G)$ with support in $S$:
-$L_{2}(S)=\left\{\varphi\in L_{2}(G),\varphi(v)=0,v\in V(G)\setminus S\right\}$
-Definition 4
-[27] We say that a set of vertices $S\subset V$ is a $\Lambda$set if for any $\varphi\in L_{2}(S)$ it admits a Poincare inequality with a constant $\Lambda=\Lambda(S)>0$
-$||\varphi||\leq\Lambda||\mathbf{\cal L}\varphi||,\,\varphi\in L_{2}(S)$ (14)
-The infimum of all $\Lambda(S)>0$ for which $S$ is a $\Lambda$set will be called the Poincare constant of the set $S$ and denoted by $\Lambda$.
-The definition above provides a tool to determine when bandlimited signals in Paley-Wiener spaces $PW_{\omega}(G)$ can be uniquely represented from their samples on a given set.
-The role of $\Lambda$-sets was explained and proved in the following Theorem by Pesenson [27], that shows that if $S\subset V$ then any signal $f\in PW_{\omega}(G)$ can be uniquely represented by its function values in the complement set $U=V(G)\setminus S$:
-Theorem 2
-[27] If $S\subset V$ is a $\Lambda$- set, then the set $U=V(G)\setminus S$ is a uniqueness set for any space $PW_{\omega}(G)$ with $\omega<\frac{1}{\Lambda}$.
-Remark: Note that non-trivial uniqueness sets can not exist for functions from any Paley-Wiener subspace $PW_{\omega}(G)$ with any $\lambda_{0}\leq\omega\leq\lambda_{N}$, but they can exist for some range $\lambda_{0}\leq\omega<\Omega$, as was shown in [27].
-We state one of our main results, in which we employ the SGW operator $\psi$ to characterize the uniqueness set using the $\Lambda_{\psi}$-set, therefore extending the $\Lambda$-set concerning the graph Laplacian operator $\cal L$.
-Theorem 3
-Let $G=(V,\mathbf{W})$ be a connected graph with $N$ vertices.
-Assume that there exist a set of vertices $S\subset V$ for which the conditions (1)-(2) in Lemma 1 below hold true.
-Let $\psi$ be the SGW operator using a polynomial $p(\mathbf{\cal L})$ with the coefficients $\left\{a_{k}\right\}_{k=0}^{K}$ such that $\psi_{f}=\sum_{k=0}^{K}a_{k}\mathbf{\cal L}^{k}f$.
-Then, for any $\varphi\in L_{2}({S})$, we have that the following inequality holds:
-$||\varphi||\leq\Lambda_{\psi}||\psi_{{\varphi}}||$ (15)
-and thus the set $S$ is a $\Lambda_{\psi}$-set for the operator $\psi$ with $\Lambda_{\psi}={\frac{1}{\sqrt{\sum_{k=0}^{K}\frac{a_{k}^{2}}{\Lambda^{2k}}}}}$.
-We recall the following results from [27].
-Note that [27] established the construction of a $\Lambda$-set by imposing specific assumptions on the sets $S$ and $U$.
-Our result Theorem 3 holds similar assumptions as in [27].
-Lemma 1 (Lemma 3.6 of [27])
-Given a connected graph $G=(V,\mathbf{W})$, a set of vertices $S\subset V$, its complement $U=V\setminus S$, for which the following conditions hold true
-  1. 1.
-    For every $s\in S$ there exists $u\in U$ that is a neighbor of $s$, i.e., $w(u,s)>0$.
-  2. 2.
-    For every $s\in S$ there exists at least one neighbor node $u\in U$ whose adjacency set intersects $S$ only over $s$.
-Then there exist a set of vertices $S\subset V$ which is a $\Lambda$-set, with $\Lambda=d(G)$.
-We require one more property of the Laplacian operation, Lemma 3 below. To this end, we use a lemma from [27],
-Lemma 2 (Lemma 3.9 of [27])
-If $S$ is a $\Lambda$-set, then for any $\varphi\in L_{2}(S)$ and all $t\geq 0$, $k=2^{l}$, $l=0,1,2,...$
-$\frac{1}{\Lambda^{k}}\|\mathcal{L}^{t}\varphi\|\leq\|\mathcal{L}^{k+t}\varphi\|,\quad\varphi\in L_{2}(S)$
-Now we are ready to prove a modification of Lemma 3.9 from [27]:
-Lemma 3
-Let $S$ be a $\Lambda$-set, $\varphi\in L_{2}(S)$, and let $m\geq 1$ be an integer. Then
-$\frac{1}{\Lambda^{m}}\|\varphi\|\leq\|\mathcal{L}^{m}\varphi\|.$
-Proof: Because every positive integer has binary representation, then we have for some $j_{i}\in\{0,1,2,...\},i=1,2,...,N$, that
-$m=\sum_{i=1}^{N}2^{j_{i}}.$
-Using Lemma 2, let $k=2^{j_{1}}$ and $t=2^{j_{2}}+\cdots+2^{j_{N}}$, so $m=k+t$. Then we have
-$\|\mathcal{L}^{m}\varphi\|$ $=\|\mathcal{L}^{k+t}\varphi\|$
-$\geq\frac{1}{\Lambda^{k}}\|\mathcal{L}^{t}\varphi\|$
-$=\frac{1}{\Lambda^{2^{j_{1}}}}\|\mathcal{L}^{2^{j_{2}}+\cdots+2^{j_{N}}}\varphi\|$
-We can now let $k_{2}=2^{j_{2}}$ and $t_{2}=2^{j_{3}}+\cdots+2^{j_{N}}$ so that,
-$\|\mathcal{L}^{m}\varphi\|\geq\frac{1}{\Lambda^{2^{j_{1}}}}\|\mathcal{L}^{2^{j_{2}}+\cdots+2^{j_{N}}}\varphi\|$ $=\frac{1}{\Lambda^{2^{j_{1}}}}\|\mathcal{L}^{k_{2}+t_{2}}\varphi\|$
-$\geq\frac{1}{\Lambda^{2^{j_{1}}}}\frac{1}{\Lambda^{k_{2}}}\|\mathcal{L}^{t_{2}}\varphi\|$
-$=\frac{1}{\Lambda^{2^{j_{1}}+2^{j_{2}}}}\|\mathcal{L}^{2^{j_{3}}+\cdots+2^{j_{N}}}\varphi\|$
-Continuing in this way, we eventually get,
-$\|\mathcal{L}^{m}\varphi\|\geq\frac{1}{\Lambda^{m}}\|\varphi\|.$
-completing the proof. $\square$
-In the next theorem, we expand the characterization of the uniqueness set using the $\Lambda$-set concerning the graph Laplacian operator $\mathcal{L}$ to include cases where we employ the SGW operator $\psi$.
-We thereby characterize the uniqueness set using the $\Lambda_{\psi}$-set for the SGW operator.
-We now turn to prove Theorem 3, which was stated earlier.
-Proof of Theorem 3:
-Let $S\subset V$ be a set satisfying the conditions in Lemma 1, so $S$ is $\Lambda$-set. Then by Lemma 3, for any integer $k\geq 1$,
-$\frac{1}{\Lambda^{k}}\|\varphi\|\leq\|\mathcal{L}^{k}\varphi\|.$
-The SGW operator of $\varphi$ has the form,
-$\psi_{\varphi}:=\sum_{i,j}\sum_{k=0}^{K}a_{k}\mathbf{\cal L}^{k}\varphi$
-for some $K\geq 1$, $a_{k}\in\mathbb{R}$ are coefficients, and where the summation over $i$ and $j$ represents the summation over the nodes of the graph. Now we calculate,
-$\|\psi_{\varphi}\|$ $=\left(\sum_{i,j}\sum_{k=0}^{K}\left|a_{k}(\mathcal{L}^{k}\varphi)_{ij}\right|^{2}\right)^{1/2}$
-$=\left(\sum_{i,j}\sum_{k=0}^{K}a_{k}^{2}\left|(\mathcal{L}^{k}\varphi)_{ij}\right|^{2}\right)^{1/2}$
-$=\left(\sum_{k=0}^{K}a_{k}^{2}\sum_{i,j}\left|(\mathcal{L}^{k}\varphi)_{ij}\right|^{2}\right)^{1/2}$
-$=\left(\sum_{k=0}^{K}a_{k}^{2}\|\mathcal{L}^{k}\varphi\|^{2}\right)^{1/2}$
-$\geq\left(\sum_{k=0}^{K}a_{k}^{2}\frac{1}{\Lambda^{2k}}\|\varphi\|^{2}\right)^{1/2}$
-$=\|\varphi\|\left(\sum_{k=0}^{K}a_{k}^{2}\frac{1}{\Lambda^{2k}}\right)^{1/2}$
-Letting
-$\Lambda_{\psi}:=\frac{1}{\sqrt{\sum_{k=0}^{K}\frac{a_{k}^{2}}{\Lambda^{2k}}}}$
-Then we have $\|\varphi\|\leq\Lambda_{\psi}\|\psi_{\varphi}\|$, so $S$ is also a $\Lambda_{\psi}$-set.
-$\square$
-From the assumptions and proof of Theorem 3, we get the immediate result:
-Corollary 1
-Let $S\subset V$ be a $\Lambda$-set. Then $S$ is also a $\Lambda_{\psi}$-set.
-Remark 1: An important property which can be observed from Theorem 3 is the following: given the SGW $\psi_{{f}}$ as a spectral representation operator applied on $f$, we can choose coefficients $\left\{a_{k}\right\}_{k=0}^{K}$ such that we obtain a $\Lambda_{\psi}$ set associated with the operator $\psi$, which is smaller than the $\Lambda$ set associated with the Laplacian operator $\cal L$.
-This implies that the operator $\psi$ provides more flexibility and better control over smoothness properties in comparison to the Laplacian operator.
-Remark 2: Since the $\Lambda_{\psi}$ set can be chosen to be smaller than $\Lambda$ set (for a proper choice of the coefficients $\left\{a_{k}\right\}_{k=0}^{K}$ using the operator $\psi$) then the SGW operator $\psi$ provides a more efficient representation for $f\in PW_{\omega^{\prime}}(G)$ with $\omega<\omega^{\prime}$ using the same subsets of nodes from the uniqueness set $U$ in comparison to the Laplacian operator $\cal L$.
-Remark 3: Note that the characterization of the the uniqueness set does not rely on a reconstruction method of the graph signal values of $f(S)$ from their known values on $U$.
-The next Theorem demonstrates the role of the $\Lambda_{\psi}$set with respect to the operator $\psi$, where we show that any signal $f\in PW_{\omega}(G)$ can be uniquely represented by its samples on the uniqueness set $U$.
-This results resembles the role of $\Lambda$-sets with respect to the graph Laplacian operator $\mathbf{\cal L}$, yet with a different bound then Lemma 2.
-Theorem 4
-Let $G=(V,\mathbf{W})$ be a connected graph with $N$ vertices and $f\in PW_{\omega}(G)$ for $\lambda_{1}<\omega<\lambda_{\mbox{max}}$.
-Given the SGW operator $\psi$, and a set $S$ which is a $\Lambda_{\psi}$ set.
-Then the set $U=V\setminus S$ is a uniqueness set for any space $PW_{\omega}(G)$ with $\omega<1/\Lambda_{\psi}$.
-Proof: Given $f,\tilde{f}\in PW_{\omega}(G)$, then $f-\tilde{f}\in PW_{\omega}(G)$.
-Assume that $f\neq\tilde{f}$.
-If $f,\tilde{f}$ coincide on $U=V\setminus S$, then $f-\tilde{f}\in L_{2}(S)$ and therefore
-$||f-\tilde{f}||\leq\Lambda_{\psi}||\psi_{f-\tilde{f}}||$ (16)
-Since $\psi_{{f}}\in\mathbb{R}^{N}$, we have that by properties of a vector space in $\mathbb{R}^{N}$, using the Cauchy–Schwarz inequality and assuming $|a_{k}|\leq 1\,\forall k$, we have:
-$||\psi_{f-\tilde{f}}||\leq\omega||f-\tilde{f}||$ (17)
-Combining the inequalities above and using the inequality $\Lambda_{\psi}\omega<1$ we have that
-$||f-\tilde{f}||\leq\Lambda_{\psi}||\psi_{f-\tilde{f}}||\leq\Lambda_{\psi}\omega||f-\tilde{f}||<||f-\tilde{f}||$ (18)
-which is a contradiction to the assumption that $f\neq\tilde{f}$. Thus, the set $U=V\setminus S$ is a uniqueness set for any space $PW_{\omega}(G)$ with $\omega<1/\Lambda_{\psi}$. $\square$
-Remark 1: Note that $\Lambda\omega<1$ implies that $\Lambda_{\psi}\omega<1$ given $\Lambda_{\psi}<\Lambda$, then we can increase the size number of nodes in the uniqueness set $U$ for $PW_{\omega}(G)$ (for a proper choice of the coefficients $\left\{a_{k}\right\}_{k=0}^{K}$ using the operator $\psi$).
-In other words, we may increase the size of $S$ (thus reducing the size of $U$) and still obtain a uniqueness set with a smaller size for the graph signals in $PW_{\omega}(G)$.
-Remark 2: We note that the results of Theorem 4 concerning the uniqueness set are independent from the stability properties of the representation.
-In order to achieve stability which is important for reconstruction, it is required to construct a wavelet operator using multiple scales $t_{j},j=1,...J$, as proposed in [18].
-More specifically, we can express the function $\varphi$ using multiple scales $t_{j}$ (here we replace the previous notation of scale $s_{j}$ with $t_{j},j=1,..J$, not to confuse with nodes $s\in S$).
-Then for a fixed scale $t_{j}$ we have that the SGW is given by $||\psi_{\varphi}(t_{j})||=\left(\sum_{i=1}^{N}\sum_{k=0}^{K}|a_{t_{j},k}({\cal L}^{k}\varphi(i))|^{2}\right)^{1/2}$ and
-$||\psi_{\varphi}||_{=}\left(\sum_{j}\sum_{i=1}^{N}\sum_{k=0}^{K}|a_{t_{j},k}({\cal L}^{k}\varphi(i))|^{2}\right)^{1/2}$ (19)
-In a similar way to the arguments provided in Theorem 3 we can choose coefficients $a_{t_{j},k}$ associated with the Laplacian polynomial such that the inequality $||\varphi||\leq\Lambda_{\psi}||\psi_{{\varphi}}||$ is satisfied.
-Remark 3: Optimal Choice of coefficients $a_{k}$ for Minimizing $\Lambda_{\psi}$ for optimal choice of $\Lambda_{\psi}$ in the multi-scale setting.
-To find the coefficients $a_{t_{j},k}$ that minimize $\Lambda_{\psi}$ where:
-$\Lambda_{\psi}=\frac{1}{\sqrt{\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{a_{t_{j},k}^{2}}{\Lambda^{2k}}}}$ (20)
-subject to the constraint: $\sum_{j=1}^{J}\sum_{k=0}^{K}a_{t_{j},k}^{2}=1$, we optimize the coefficients $a_{t_{j},k}$, which control how much of the signal at scale $t_{j}$ and power $k$ is captured by the SGW operator.
-We use the Lagrangian $\mathcal{F}$, incorporating the energy constraint into the optimization problem and Define:
-$\mathcal{F}(a_{t_{1},0},\dots,a_{t_{J},K},\mu)=\frac{1}{\sqrt{\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{a_{t_{j},k}^{2}}{\Lambda^{2k}}}}+\mu\left(\sum_{j=1}^{J}\sum_{k=0}^{K}a_{t_{j},k}^{2}-1\right)$ (21)
-[…]
-Taking partial derivatives, solving for $a_{t_{j},k}$, and differentiating $\mathcal{L}$ with respect to each coefficient $a_{t_{j},k}$:
-$\frac{\partial\mathcal{F}}{\partial a_{t_{j},k}}=\frac{-a_{t_{j},k}/\Lambda^{2k}}{\left(\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{a_{t_{j},k}^{2}}{\Lambda^{2k}}\right)^{3/2}}+2\mu a_{t_{j},k}=0.$ (22)
-rearranging the equation above we obtain:
-$\frac{a_{t_{j},k}}{\Lambda^{2k}}=2\mu a_{t_{j},k}\left(\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{a_{t_{j},k}^{2}}{\Lambda^{2k}}\right)^{3/2}$
-assuming $a_{t_{j},k}\neq 0$, we can divide both sides of the equation by $a_{t_{j},k}$, yielding:
-$\frac{1}{\Lambda^{2k}}=2\mu\left(\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{a_{t_{j},k}^{2}}{\Lambda^{2k}}\right)^{3/2}$ (23)
-thus the sum $\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{a_{t_{j},k}^{2}}{\Lambda^{2k}}$ is proportional to the inverse of $\Lambda^{2k}$, implying that the optimal wavelet coefficients should decay proportionally to the inverse powers of the Laplacian eigenvalues $a_{t_{j},k}\propto\frac{1}{\Lambda^{k}}$.
-Next, applying the energy constraint: $\sum_{j=1}^{J}\sum_{k=0}^{K}a_{t_{j},k}^{2}=1$ and substituting $a_{t_{j},k}=\frac{1}{\Lambda^{k}}$, we obtain:
-$\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{1}{\Lambda^{2k}}=1$ (24)
-Thus, the normalized optimal coefficients $a_{t_{j},k}$ are:
-$a_{t_{j},k}=\frac{\frac{1}{\Lambda^{k}}}{\sqrt{\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{1}{\Lambda^{2k}}}}.$ (25)
-Note that the wavelet energy is distributed across all scales and powers of the Laplacian, with a decay that prioritizes low-frequency components (small $k$).
-Using the form of the coefficients $a_{t_{j},k}$ above, the expression for $\Lambda_{\psi}$ is:
-$\Lambda_{\psi}=\frac{1}{\sqrt{\sum_{j=1}^{J}\sum_{k=0}^{K}\frac{1}{\Lambda^{4k}}}}$ (26)
-which is the minimized value of $\Lambda_{\psi}$, showing that the optimal choice of coefficients reduces the overall constant, improving the localization and efficiency of the SGW operator.
-Thus, the optimal choice of wavelet coefficients $a_{t_{j},k}$ that minimizes $\Lambda_{\psi}$ in the multi-scale SGW operator is proportional to $\frac{1}{\Lambda^{k}}$, prioritizing low-frequency components.
+Identifying the key factors in high-dimensional datasets is essential for leveraging unlabeled data in many practical applications.
+In this work, we introduce a novel contrastive learning framework for interpretable manifold learning via graph embeddings, leveraging both low- and high-frequency information to enhance representation quality.
+Our method employs Spectral Graph Wavelets (SGW) to construct a multi-scale graph representation of the input feature space.
+This representation is optimized using a stochastic gradient descent (SGD)-based scheme, combined with a novel 3D tensor encoding strategy that enriches the embedding process.
+MS-IMAP differs fundamentally from traditional dimensionality reduction or visualization techniques.
+While many manifold learning methods (e.g., Laplacian Eigenmaps, Diffusion Maps, UMAP) aim to reveal intrinsic structure through low-dimensional embeddings, MS-IMAP targets a broader objective: constructing structure-preserving, interpretable embeddings suitable for downstream tasks.
+By design, MS-IMAP preserves input dimensionality and enforces a one-to-one correspondence between input features and embedding dimensions.
+To enhance the embedding with structural information, we introduce a 3D tensor that captures SGW coefficients across nodes, features, and frequency bands.
+This enables the model to encode both global and local graph structure.
+Importantly, although SGW captures information across the graph, each embedding coordinate is updated solely based on the SGW coefficients of its corresponding input feature.
+This preserves the interpretability and separability of features throughout training.
+We demonstrate the utility of these interpretable embeddings for both feature selection and traditional clustering tasks.
+Specifically, we assess clustering performance based on features selected from both the input space and the embedding space, using two unsupervised methods: Laplacian Score and Mutual Information-based selection.
+A key advantage of our framework is its inherent linkage between the original and embedded feature spaces—a property rarely offered by existing nonlinear manifold learning techniques.
+Furthermore, we show that the resulting embeddings are highly competitive with state-of-the-art graph embedding methods.
+To further justify the use of SGW for representation, we analyze its theoretical properties by studying functions in Paley-Wiener spaces on combinatorial graphs.
+We show that the SGW operator enables more expressive representations, characterized through the concept of the $\Lambda$-set.
+Limitations and future directions include several aspects.
+First, our approach assumes that the input features—and the similarity measures used—adequately capture the underlying manifold structure, particularly geodesic distances.
+Second, extending the current SGD-based optimization framework to handle out-of-sample generalization remains a challenge.
+Finally, while the embedding-feature correspondence facilitates global feature importance analysis, further improvements may be achieved by incorporating localized importance estimation techniques.